### v2

# K-theory and localizing invariants of large categories

Abstract.
In this paper we introduce and study the so-called continuous -theory for a certain class of “large” stable -categories, more precisely, for dualizable presentable categories. For compactly generated categories, the continuous -theory is simply the usual (non-connective) -theory of the full subcategory of compact objects. More generally, we show that any localizing invariant of small stable -categories can be uniquely extended to a localizing invariant of dualizable categories.
We compute the continuous -theory for categories of sheaves on locally compact Hausdorff spaces. Using the special case for sheaves on the real line, we give an alternative proof of the theorem of Kasprowski and Winges [KW19] on the commutation of -theory with infinite products for small stable -categories.
We also study the general theory of dualizable categories. In particular, we give an “explicit” proof of Ramzi’s theorem [Ram] on the -presentability of the category of dualizable categories. Among other things, we prove that dualizability is equivalent to “flatness” in the category of presentable stable categories.
Contents
- 0 Introduction
1 General theory of dualizable categories
- 1.1 Infinity-categories, limits and colimits
- 1.2 Presentable and accessible stable -categories
- 1.3 Well generated triangulated categories
- 1.4 Dualizable categories
- 1.5 Examples of compactly assembled categories, stable and non-stable
- 1.6 Continuous stabilization
- 1.7 Duality as a covariant involution on dualizable categories
- 1.8 Dualizability via compact morphisms
- 1.9 The axiom (AB6)
- 1.10 Automatic dualizability and images of functors
- 1.11 Calkin categories
- 1.12 Colimits of dualizable categories
- 1.13 Semi-orthogonal decompositions
- 1.14 Finite limits of dualizable categories
- 1.15 General limits of dualizable categories
- 1.16 Example: fpqc descent
- 1.17 Products of dualizable categories
- 2 Dualizability is equivalent to flatness
- 3 Extensions of compactly generated categories
- 4 Localizing invariants
- 5 Sheaves and cosheaves on continuous posets
- 6 Sheaves on locally compact Hausdorff spaces
- A Monomorphisms and epimorphisms of presentable and dualizable categories
- B Essential image of a homological epimorphism
- C Presentability of the category of dualizable categories
- D Urysohn’s lemma for dualizable categories
- E Phantom maps, pure injectivity and Adams representability
- F Analogy between dualizable categories and compact Hausdorff spaces
- G K-theory of products of exact infinity-categories
0. Introduction
It is well known that the Grothendieck group of an additive category vanishes whenever has countable direct sums. More generally, for any stable -category with countable coproducts its -theory spectrum is contractible. This vanishing also holds for any additive invariant, like (topological) Hochschild homology. Therefore, naturally, additive invariants are normally studied only for small stable infinity categories, and not for “large” categories.
In this paper we introduce and study the notion of continuous non-connective -theory (and more general localizing invariants) for a certain class of “nice” large categories. These are the so-called dualizable (also known as compactly assembled) presentable stable categories. Moreover, this continuous -theory in fact extends the usual non-connective -theory for small stable -categories.
We now explain more precisely the basic idea. Let us denote by the -category of small idempotent-complete stable -categories and exact functors between them. We denote by the -category of spectra. We consider the non-connective -theory as a functor We refer to [BGT13, Bar16] for the definition of (non-connective) algebraic -theory of a small stable -category, based on the Waldhausen’s -construction [Wal85]. If is a usual (discrete) associative unital ring, then the -theory space of the stable -category of perfect complexes is equivalent to the Quillen’s -theory space of as defined in [Qui73].
One of the main properties of -theory is the localization. We recall that a short exact sequence in is a fiber-cofiber sequence that is, a biCartesian square of the form
In terms of the (triangulated) homotopy categories, the square (0.1) is Cartesian if and only if the functor induces an equivalence Assuming this, the square (0.1) is coCartesian if and only if the induced functor is fully faithful and the Karoubi completion of its image is equivalent to In other words, a fiber-cofiber sequence (0.1) induces a short exact sequence of triangulated categories, up to direct summands:
Localization property for non-connective -theory means that the -theory functor sends each fiber-cofiber sequence (0.1) to a fiber (=cofiber) sequence in
Now, consider the -category of presentable stable -categories and colimit-preserving functors. Recall that a presentable stable category is called dualizable if it is a dualizable object in the symmetric monoidal category with the Lurie tensor product. This is equivalent to being a retract in of a compactly generated category, by [Lur18, Proposition D.7.3.1].
We call an exact functor of presentable stable -categories strongly continuous if has a right adjoint which is colimit-preserving. We denote by the (non-full) subcategory of dualizable categories and strongly continuous functors. Further, we denote by the subcategory of compactly generated categories.
It is well-known that the functor is an equivalence, and an inverse equivalence is given by (the latter is the full subcategory of compact objects). Therefore, we have a fully faithful functor We claim that the non-connective -theory can be naturally extended to dualizable presentable categories in the following sense: there is a functor and a natural isomorphism of functors Moreover, is uniquely determined (up to a contractible space of choices) by the localization property: it sends short exact sequences in to fiber sequences in
Uniqueness essentially follows from another characterization of dualizable categories: a category is dualizable if and only if there exists a short exact sequence in with strongly continuous functors, such that and compactly generated, see Proposition 1.18. If we require to be a localizing invariant, then we are forced to have
However, it is not very reasonable to take (0.2) as a definition of for the following reasons:
It is not immediately obvious that does not depend on the choice of a short exact sequence; this can be deduced from [Tam18, Theorem 18].
It is not clear that a strongly continuous functor of dualizable categories induces a map
Instead, we define the continuous -theory using Calkin categories. Recall the notion of Calkin algebra from functional analysis: for a Hilbert space the Calkin algebra is defined as the quotient of the algebra of bounded operators by the ideal of compact operators Since is the closure of the ideal of bounded operators of finite rank, we have a natural discrete analogue. Namely, for a vector space over a field the Calkin algebra is defined as the quotient This naturally generalizes to a Calkin category associated to a small stable -category, which we now recall.
For the Calkin category is defined as the (Karoubi completion of) the quotient:
If we ignore set-theoretic issues for now (they are easily resolved), then the short exact sequence implies an isomorphism since
We claim that Calkin categories naturally generalize to dualizable categories. Namely, one can define a functor so that we have a natural equivalence of functors For this we use one more characterization of dualizable categories: a stable presentable -category is dualizable if and only if the colimit functor has a left adjoint. Denoting it by we see that the quotient is compactly generated. We define
It is easy to see that for a small stable -category we have
Ignoring set-theoretic issues again, we define the continuous -theory by the formula We refer to Subsection 4.2 for a more precise definition.
It is not difficult to check that the functor preserves short exact sequences, hence is a localizing invariant. The above discussion also implies an isomorphism Similarly, we can apply the same construction to any localizing invariant. We have the following result.
Theorem 0.1.
Let be a stable -category. The precomposition functor
induces an equivalence between the full subcategories of localizing invariants. The inverse equivalence is given by
The detailed proof is given in Subsection 4.2 below. Moreover, in Subsection 4.3 we show that is simply the right Kan extension of via the functor
The assignment preserves the important properties of localizing invariants. In particular, preserves filtered colimits if and only if so does
Examples of dualizable (but not necessarily compactly generated) categories naturally arise in various contexts, see Subsection 1.5. One class of examples is provided by almost mathematics (Faltings [Fa02], Gabber-Ramero [GabRam03]). Namely, let be an associative unital ring, and a two-sided ideal such that such that for (for example, is flat as a right or left -module). Then the extension of scalars functor is a localization, and its kernel is a dualizable category. We have an equivalence where is the abelian category of almost modules: Moreover, if is contained in the Jacobson radical of then the category has no non-zero compact objects.
Another class of examples which we are interested in this paper are categories of sheaves on a locally compact Hausdorff space Let be a presheaf on with values in We denote by the category of sheaves with values in By [Lur18], the category is dualizable. We have the following general result, see 6.10.
Theorem 0.2.
With the above notation, let be a presentable stable -category, and a localizing invariant which commutes with filtered colimits. Then we have a natural isomorphism
We note that even when the presheaf takes values in compactly generated categories (and even when it is constant), the category is usually not compactly generated, see Subsection 6.4. For example, when is non-compact and connected, and the presheaf is constant and non-zero, then the category has no non-zero compact objects. This is a straightforward generalization of Neeman’s theorem [Nee01b].
The following is a special case of Theorem 0.2.
Corollary 0.3.
Let be a dualizable category. Then for any we have a natural isomorphism
In particular, we have
Restricting to finite CW complexes, we see that the categories of -valued sheaves in fact “categorify” the maps to in the following sense.
Corollary 0.4.
Let be a finite CW complex (hence a compact Hausdorff space). Let be a dualizable category. Then we have a natural isomorphism
In particular, if is a small stable idempotent complete -category, then we have where the RHS is the abelian group of homotopy classes of maps from to the -theory space of
It turns out that Corollary 0.4 holds for arbitrary localizing invariants, not necessarily commuting with filtered colimits, see Theorem 6.1.
Another interesting class of examples of dualizable categories comes from condensed mathematics: these are categories of (strongly) nuclear modules over analytic rings. We consider the following special case (see Example 1.30).
Let be a noetherian commutative ring, and let be an ideal. Clausen and Scholze [CS20] defined the category of nuclear solid modules over This is a dualizable category, and the full subcategory of compact objects in is equivalent to the usual category of perfect complexes We will study the category in [E1], where we will in particular prove that
The paper is organized as follows.
In Section 1 we study the general theory (not necessarily cocomplete) compactly assembled -categories, mostly focusing on the dualizable presentable stable -categories. In Subsection 1.4 we recall the equivalent definitions of dualizable categories from [Lur18] and study the basic properties. In Subsection 1.5 we give examples of compactly assembled categories, stable and non-stable. In Subsection 1.8 we discuss the criterion of dualizability in terms of homotopy categories, which almost follows from H. Krause’s papers [Kr00, Kr05]. In Subsection 1.9 we explain a criterion of dualizabilifty due to Clausen and Scholze: a presentable stable category is dualizable if and only if it satisfies the axiom (AB6), i.e. the distribution of products over filtered colimits. In Subsection 1.11 we define the continuous Calkin category of a dualizable category; this construction is important for the definition of continuous -theory. In Subsection 1.12 we study the colimits in In particular, we show that the inclusion commutes with colimits. We also show that the category satisfies the weak (AB5) axiom: the class of fully faithful functors is closed under filtered colimits (Proposition 1.68). In Subsections 1.14 and 1.15 we discuss the limits of dualizable categories. The (complicated) general description of limits is given by Theorem 1.92. As an example, we show in Subection 1.16 that the functor is a sheaf for the fpqc topology (here denotes the category of ordinary commutative rings). In Subsection 1.17 we show that the category satisfies the weak (AB4*) axiom: the product of epimorphisms is an epimorphism. We also show that the axiom (AB6) holds in
In Section 2 we prove a surprising result: a presentable stable category is dualizable if and only if it is flat in More precisely, is flat if the functor preserves fully faithful functors.
In Section 3 we give a criterion of dualizability for categories which are extensions of compactly generated categories. If such an extension is dualizable, then it is in fact compactly generated. As an application, we show that in the extensions of by are classified by functors where is the category of Tate objects, see [Hen17]. As another application, for a commutative noetherian ring we classify the localizing subcategories of which are dualizable. They correspond bijectively to convex subsets of see Theorem 3.10.
In Section 4 we define and study localizing invariants of dualizable categories. In Subsection 4.1 we recall for each regular cardinal the universal localizing invariant of small stable categories, commuting with -filtered colimits. In subsection 4.2 we explain that a localizing invariant of small categories extends uniquely to dualizable categories, using the Calkin construction. In Subsection 4.3 we explain that this canonical extension is in fact the right Kan extension. In Subsection 4.4 we give an example: we compute finitary localizing invariants of the category of sheaves on and on In Subsection 4.5 we do a similar computation for not necessarily finitary localizing invariants, deducing the commutation of K-theory with products (Theorem 4.29).
In Section 5 we consider the sheaves and cosheaves on continuous posets (these posets are considered as categories, not as sets), and compute their (finitary) localizing invariants. We use this computation later to compute the localizing invariants for sheaves on locally compact Hausdorff spaces.
In Section 6 we study the sheaves on locally compact Hausdorff space with values in a presheaf of dualizable categories. First, in Subsection 6.1 we give a computation of arbitrary localizing invariants for sheaves on finite CW complexes (in the case of constant coefficients). In Subsection 6.2 we explain how to approximate the category of sheaves on an arbitrary compact Hausdorff space by “simpler” categories (Proposition 6.9). Here it is crucial that we use the notion of a -(pre)sheaf from [Lur09]. In Subsection 6.3 we apply the previous results to compute the finitary localizing invariants of categories of sheaves (Theorem 6.10). In Subsection 6.4 we give a simple description of compact objects in the categories of sheaves.
In Appendix A we give the (expected) description of monomorphisms and epimorphisms of presentable and dualizable categories. In Appendix B we show that an image of a homological epimorphism does not have to be a stable subcategory (Proposition B.1). In Appendix C we show that the category is -presentable, and we describe the -compact objects for each uncountable regular (Theorem C.6). The -presentability of is originally due to Ramzi [Ram]. In Appendix D we prove that the category is generated by colimits by a single object – the category of sheaves of spectra on the real line with singular support in In Appendix E we prove an analogue of Adams representability theorem for -compact dualizable categories (Theorem E.11). In Appendix F we explain an analogy between the category of dualizable categories and the opposite category of compact Hausdorff spaces. Finally, in Appendix G we show that -theory of exact -categories commutes with infinite products, using the ideas of Córdova Fedeli [Cor23] and Klemenc [Kle22].
Acknowledgements. I am grateful to Ko Aoki, Alexander Beilinson, Dustin Clausen, Adriano Córdova Fedeli, Vladimir Drinfeld, Boris Feigin, Dennis Gaitsgory, Marc Hoyois, Dmitry Kaledin, David Kazhdan, Maxim Kontsevich, Akhil Mathew, Thomas Nikolaus, Dmitri Orlov, Maxime Ramzi, Victor Saunier, Tomer Schlank, Peter Scholze, Vivek Shende, Vladimir Sosnilo, Georg Tamme, Bertrand Toën, Yakov Varshavsky and Christoph Winges for useful discussions. Part of this work was done while I was a visitor in the Max Planck Institute for Mathematics in Bonn from December 2022 till February 2024, and I am grateful to the institute for their hospitality and support.
1. General theory of dualizable categories
1.1. Infinity-categories, limits and colimits
We will freely use the theory of -categories, functors, limits and colimits as developed in [Lur09, Lur17]. Except for Section 2, we will deal only with -categories. Given an ordinary category we identify it with its nerve which is an -category. In particular, we will consider a poset (partially ordered set) as an -category. Sometimes we will say “category” instead of “-category”, when the meaning is clear from the context.
We denote by the -category of spaces (or equivalently -groupoids). It is freely generated by one object via colimits.
We will use the following convention: a functor between -categories is cofinal if for any the -category is weakly contractible. Equivalently, for any -category and for any functor we have
assuming that one of the colimits exists. Dually, a functor is final if the functor is cofinal.
Recall that an infinite cardinal is called regular if for any collection of cardinals such that and we have For example, any (infinite) successor cardinal is regular. If is regular, then a set is called -small if
We refer to [Lur09, Definition 5.3.1.7] for the notion of a -filtered -category. The argument of [AR94, Theorem 1.5, Remark 1.21] shows that for any -filtered -category there exists an -directed poset and a cofinal functor Recall that a poset is called -directed if is nonempty and any -small collection of elements of has an upper bound. Recall that -filtered -categories resp. -directed posets are simply called filtered -categories resp. directed posets.
Given a small -category and a regular cardinal we denote by the -category which is freely generated by via -filtered colimits. The category can be described as a full subcategory of the category of presheaves which is formed by -filtered colimits of representable presheaves. We will usually denote the -ind objects of by where is a -filtered -category, and we have a functor For we simply write instead of
If is not necessarily small (but still locally small), we will sometimes still consider the (locally small) category which is the directed union of where runs through small full subcategories of
If is an -category with -filtered colimits, then an object is called -compact if the functor commutes with -filtered colimits. We denote by the full subcategory of -compact objects. The -compact objects are called compact objects.
Recall that an -category is called stable if it is pointed, has finite limits and colimits, and a square of the form
is Cartesian if and only if it is coCartesian.
In a general pointed -category, if a square of the form (1.1) is Cartesian then it is also called a fiber sequence, and is called a fiber of Dually, if a square of the form (1.1) is coCartesian then it is also called a cofiber sequence, and is called a cofiber of
If an -category is stable, then the category has a natural triangulated structure. A functor between stable -categories is exact if it is pointed, and takes fiber sequences in to fiber sequences in If is exact, then the induced functor is an exact functor between triangulated categories. By default, the functors between stable categories will be assumed to be exact.
For an -ring we will denote by the (cocomplete, stable) category of right -modules. We denote by the full subcategory of perfect -modules; it is generated by as a stable idempotent-complete subcategory.
We denote by the category of small -categories. Further, we denote by the (non-full) subcategory of small stable categories and exact functors. We denote by the full subcategory of Karoubi complete (idempotent-complete) stable categories.
Given an -category we denote by its Karoubi completion. In particular, the functor is the left adjoint to the inclusion.
By definition, a short exact sequence in is a fiber-cofiber sequence
This means that the functor is fully faithful, the composition is zero, and the functor is an equivalence. We will simply write
for such short exact sequences. We will also use the same terminology and notation for short exact sequences of “large” stable categories, namely presentable and in particular dualizable.
By [BGT13, Corollary 4.25], the categories and are compactly generated. See also Proposition 1.4 and Remark 1.5 for a different argument.
We recall the notion of a homological epimorphism between idempotent-complete stable -categories. A functor is a homological epimorphism if the functor is a quotient functor, or equivalently its right adjoint is fully faithful. In fact, is a homological epimorphism if and only if it is an epimorphism in
1.2. Presentable and accessible stable -categories
An -category is called -accessible for some regular cardinal if where is a small -category. Note that in this case is stable if and only if the category is stable.
Further, an -category is accessible if it is -accessible for some If and are accessible -categories with -filtered colimits (but and are not necessarily -accessible), we say that a functor is -continuous if commutes with -filtered colimits. Further, is continuous if it is -continuous, i.e. commutes with filtered colimits, assuming that and have filtered colimits.
A -continuous functor is called -accessible if moreover and are -accessible. A functor is called accessible if it is -accessible for some
An -category is -presentable if is -accessible and cocomplete. Equivalently, this means that where the category has -small colimits. Further, is presentable if it is -presentable for some Following [Lur09], we denote by the -category of all presentable categories and colimit-preserving functors. Note that the category is not locally small.
A functor between presentable categories is colimit-preserving if and only if it has a right adjoint. On the other hand, has a left adjoint if and only if it is accessible and commutes with limits. This gives an equivalence where the -morphisms in are accessible functors which commute with limits.
If is presentable and is cocomplete, we will denote by the category of colimit-preserving functors. The category has a natural symmetric monoidal structure, for which the internal from to is given by The tensor product of presentable categories is usually called the Lurie tensor product. It is easy to show that for The unit object is given by – the -category of spaces.
For a regular cardinal we denote by the non-full subcategory of -presentable categories, where -morphisms in are colimit-preserving functors which preserve -compact objects. The category is locally small. Moreover, the assignment defines an equivalence between and the category of idempotent-complete small -categories with -small colimits, where the -morphisms are functors which commute with -small colimits. If then the existence of -small colimits automatically implies idempotent-completeness. The Lurie tensor product induces a well-defined symmetric monoidal structure on
For any regular cardinal the inclusion functor commutes with colimits. For the right adjoint to the functor is given by
We denote by the full subcategory of presentable stable categories, and similarly for Note that the category of spectra is an idempotent -algebra in and the subcategory is simply the category of modules over (and similarly for ). In particular, the inclusions and commute with limits and colimits.
Below we will mostly deal with presentable stable categories. By default, the functors between stable categories will be assumed to be exact. Note that an exact functor between presentable stable -categories is continuous if and only if it is colimit-preserving or equivalently commutes with coproducts. Equivalently, this means that the functor between the homotopy categories commutes with coproducts.
A cocomplete stable category is -presentable if and only if it is compactly generated by a set of objects. The assignment defines an equivalence
Definition 1.1.
1) Given an exact functor between small stable categories, we denote by the semi-orthogonal gluing of and via More precisely, is the category of triples where and Equivalently, is the oplax limit of the diagram in the -category of stable categories. As explained in [CDW23, Theorem 2.4.2] (and references therein), the four categories obtained by taking an (op)lax (co)limit of the diagram are naturally equivalent.
2) If and are cocomplete stable categories and is an arbitrary exact functor, we also denote by the same semi-orthogonal gluing as above.
We make the following observation.
Proposition 1.2.
Let be an exact functor between presentable stable categories. The following are equivalent:
the category is presentable;
the functor is accessible.
Moreover, if is -accessible, then is -presentable.
Proof.
Denote by and the inclusion functors given by Clearly, the functors and are continuous.
. Suppose that is presentable. Then the functor (the right adjoint to ) is accessible. Hence, so is the composition which is isomorphic to
. It is sufficient to prove the “moreover” statement. Suppose that and are -presentable and commutes ith -filtered colimuts. Then an object is -compact in whenever We conclude that is -presentable, and the full subcategory is generated by and (as a stable subcategory). ∎
Remark 1.3.
If we denote by the -category of presentable stable categories and accessible functors, then the category has again four interpretations as an (op)lax (co)limit of the arrow
The statements of the following proposition are well known. We include the proof for completeness.
Proposition 1.4.
Let be a regular cardinal.
1) The category is generated by colimits by a single compact object (the linearly ordered set with elements). In particular, is compactly generated.
2) The category is generated by colimits by a single -compact object – the arrow category of spaces. In particular, is -presentable.
3) The category is generated by colimits by a single -compact object In particular, is -presentable and is compactly generated.
Proof.
For 1) it suffices to show that if is a functor between small -categories such that the map is an equivalence of spaces, then is an equivalence.
It is clear that is essentially surjective and the map is an equivalence of spaces. For the space is the fiber of the map over and similarly for This implies that is fully faithful.
To deduce 2) from 1) it suffices to recall that the functor commutes with -filtered colimits.
3) The -compactness of in follows from the above commutation with -filtered colimits. It remains to show that the functor is conservative.
Let be an exact functor between small stable categories such that the induced functor is an equivalence of spaces. We need to show that is an equivalence. It is clear that is essentially surjective. It remains to observe that for and for each the abelian group is naturally a retract of and similarly for ∎
Remark 1.5.
The assignment defines an equivalence between and the category of pairs where and is a subgroup. This follows from Thomason’s theorem [Th97, Theorem 2.1].
In particular, the category is compactly generated and iff and the group is finitely generated. If is an uncountable regular cardinal, then is a -compact object if and only if and is -small.
We will need an explicit characterizations of -compact objects in for uncountable
Proposition 1.6.
Let be an uncountable regular cardinal and let be a small idempotent-complete stable category. The following are equivalent.
the object is -compact in
the triangulated category is generated (as an idempotent-complete triangulated subcategory) by a -small set of objects, and moreover for any the spectrum is -compact in
Proof.
Let be in and choose a -small ind-system of compact objects of such that Then for each the spectra are -compact, . Hence, the spectra are -compact,
Since each is generated by a single object, we conclude that is generated by a -small set of objects.
First suppose that is generated by a single object and consider the -algebra Since the underlying spectrum of is -compact, it follows from Proposition C.2 that is -compact in the category of -rings. Now the category of -rings is equivalent to the full subcategory of the category formed by pairs such that generates It follows that is -compact.
In general, if is generated by a -small collection of objects then is a directed union of idempotent-complete stable subcategories generated by a finite subset of objects By the above, each category is -compact in hence so is ∎
We recall that the the colimit of a functor is equivalent to the limit of the corresponding functor where the transition functors are given by the right adjoints. We denote this colimit by (to avoid confusion with the colimits of small stable categories).
We call an exact functor between presentable stable -categories strongly continuous if it has a right adjoint which is continuous (in other words, if has a twice right adjoint). It is convenient to introduce a non-full subcategory with the same objects and with -morphisms being strongly continuous functors. Note that is locally small. Given presentable stable categories and we denote by the (small) category of strongly continuous functors.
Similarly, for a regular cardinal we denote by the full subcategory of -presentable categories. In particular, is a subcategory of (for we have an equality ).
Note that if and are -presentable stable categories, then a colimit-preserving functor preserves -compact objects (i.e. is a -morphism in ) if and only if its right adjoint commutes with -filtered colimits. For this reason, we will sometimes say that such a functor is -strongly continuous.
Proposition 1.7.
The category is cocomplete and the functor commutes with colimits. In particular, the categories are also cocomplete and all the inclusion functors and commute with colimits.
Proof.
This follows directly from the description of a colimit in as a limit of the opposite diagram. Namely, consider a small -category and a functor and put (the colimit is taken in ). Identify with the limit Since the transition functors in this inverse system are continuous, the coproducts in the limit category are computed “pointwise”.
Let We need to show that a continuous functor is strongly continuous if and only if the compositions are strongly continuous. But the functor corresponds to the compatible system of functors under the identification of with the inverse limit. Hence, is continuous if and only if all are continuous.∎
Remark 1.8.
It seems plausible that for any regular cardinal the category is presentable. However, it is not clear how to describe the functor right adjoint to the inclusion.
We make an observation on the automatic strong continuity.
Proposition 1.9.
Let be a continuous functor between presentable stable categories. Let be a family of strongly continuous functors, where are presentable stable categories. Suppose that the images of generate and the compositions are strongly continuous. Then is also strongly continuous.
Proof.
Indeed, our assumptions imply that the right adjoints are continuous and they form a conservative family. Further, the compositions are continuous. Hence, the functor is also continuous, as required.∎
Given a presentable stable category we will call a full subcategory localizing if is presentable stable and the inclusion functor commutes with colimits. Equivalently, this means that is generated as a cocomplete stable subcategory by a set of objects (or by a single object).
1.3. Well generated triangulated categories
The notion of presentability for a stable -category can be formulated in terms of its triangulated homotopy category Namely, is presentable if and only if the triangulated category is well generated in the terminology of Neeman [Nee01a]. We briefly recall the relevant notions.
Let be a triangulated category. The category is called cocomplete if it has all small direct sums (coproducts). Given an infinite regular cardinal an object is called -compact if for any family of objects for any morphism there exists a subset with such that factors through In particular, is -compact if and only if it is compact, i.e. the functor commutes with small direct sums.
A cocomplete triangulated category is called -well generated for a regular cardinal if there exists a generating set of -compact objects This means that the objects generate via small direct sums, shifts and cones. Further, is called well generated if it is -well generated for some
If is a well generated triangulated category, we say that a full triangulated subcategory is localizing if it is closed under coproducts, and it is generated by a set of objects.
Let now and be cocomplete stable categories, and let be a regular cardinal.
is -presentable if and only if is -well generated;
An exact functor is continuous if and only if the functor commutes with coproducts;
An object is -compact in the -categorical sense if and only if is -compact in the above sense.
If is presentable, then a full subcategory is localizing if and only if is localizing.
Recall that for a triangulated category an additive functor is called cohomological if for any exact triangle the sequence is exact. By [Nee01a, Theorem 1.17], a well generated triangulated category satisfies Brown representability for contravariant functors: an additive functor is representable if and only if is cohomological and commutes with products. In particular, an exact functor between well generated triangulated categories has a right adjoint iff commutes with coproducts.
A much more difficult question is the Brown representability for covariant functors, and it is not known if it holds for abstract well generated categories. But if where is presentable stable, then it follows from [Nee09, Theorems 1.11, 1.17 and 1.22] and [Lur09, Proposition A.3.7.6] that an additive functor is corepresentable if and only if is cohomological and commutes with products. In particular, for such and for any well generated triangulated category an exact functor has a left adjoint iff commutes with products.
1.4. Dualizable categories
We first recall the notion of a compactly assembled (not necessarily stable and not necessarily cocomplete) category from [Lur18].
Definition 1.10.
Let be an accessible -category with filtered colimits. Then is called compactly assembled if the colimit functor has a left adjoint.
If the -category is compactly assembled, then we denote by (or simply ) the (fully faithful) functor which is the left adjoint to Note that the right adjoint to is always given by the Yoneda embedding Any compactly generated (i.e. -accessible) category is compactly assembled. Indeed, if then the functor is obtained by applying to the Yoneda embedding
Compactly assembled ordinary categories were studied in great detail by Johnstone and Joyal [JJ82] under the name “continuous categories”. This is a generalization of the notion of a continuous poset, see Subsection 5.1 for a more precise definition and [GHKLMS80, BH81] for a detailed account.
We will need the following notion of strong continuity for functors between compactly assembled categories.
Definition 1.11.
1) Let and be compactly assembled accessible categories. A functor is called strongly continuous if is continuous and the natural transformation is an isomorphism.
2) We denote by the -category of compactly assembled accessible categories, where the -morphisms are given by strongly continuous functors.
For stable compactly assembled categories, the two notions of strong continuity of exact functors are equivalent by Proposition 1.15.
A presentable stable -category is called dualizable if it is a dualizable object of the symmetric monoidal category (or equivalently of the symmetric monoidal homotopy category ). That is, there exists an -category and continuous functors and such that the compositions
are isomorphic to the identity.
It turns out that a presentable stable category is compactly assembled if and only if is is dualizable. This is part of the following result due to Lurie.
Proposition 1.12.
[Lur18, Proposition D.7.3.1] Let be a presentable stable -category. The following are equivalent.
The -category is dualizable as an object of
is compactly assembled.
is a retract in of a compactly generated category.
Let be another presentable stable category and be a continuous localization, i.e. the right adjoint is fully faithful. Then has a continuous section.
Remark 1.13.
Given a dualizable category we have a natural functor which can be described (at least) in two ways. The first one is the following: given the object corresponds to the continuous functor
Another description is via the universal problem: we have
In the special case where is an -ring, we have – the category of left -modules. For a right -module we have – the usual dual left module of a right module.
Remark 1.14.
Essentially rephrasing the previous remark: given a dualizable category and an object the ind-object corresponds to the functor
Proposition 1.15.
Let be a continuous functor between dualizable categories. Then the right adjoint is continuous if and only if the natural transformation is an isomorphism.
Proof.
The corresponding natural transformation between the right adjoints is given by of functors Tautologically, is an isomorphism if and only if commutes with filtered colimits, or equivalently with all colimits (since is exact).∎
We now define the category of dualizable categories.
Definition 1.16.
The category is the (non-full) subcategory of formed by dualizable categories and strongly continuous functors. Equivalently, is the full subcategory of formed by dualizable categories.
In particular, the category is locally small (since so is ). Note that is a (non-full) subcategory.
We make an immediate observation about short exact sequences.
Proposition 1.17.
Consider a short exact sequence in If is dualizable, then so are and
Proof.
Indeed, both and are retracts of in the right adjoint functors and are continuous, and we have ∎
We need one more characterization of dualizable categories, which is easy to deduce from Proposition 1.12 and which is implicit in the proof of [Lur18, Proposition D.7.3.1].
Proposition 1.18.
Let be a presentable stable -category. Then is dualizable if and only if there exists a small stable -category and a strongly continuous fully faithful functor Equivalently, there is a short exact sequence in of the form
Proof.
If we have a strongly continuous fully faithful functor then is dualizable by Proposition 1.17.
Conversely, suppose that is dualizable. By the criterion (ii) of Proposition 1.12, we have a fully faithful strongly continuous functor As explained in [Lur18, Proof of Proposition D.7.3.1] the essential image of is contained in for some small stable subcategory We obtain a fully faithful strongly continuous functor as required.∎
We make the following observation about compact objects, which follows easily from theorems of Neeman and Thomason.
Proposition 1.19.
Let be a dualizable category, and let be a full stable subcategory. Then we have an equivalence For any compact object the object can be lifted to a compact object of
Proof.
If is compactly generated, then the result follows from [Nee96, Theorem 2.1].
In general, we choose a fully faithful strongly continuous functor where is compactly generated. Then the equivalence induces an equivalence
The second assertion of the proposition follows from Thomason’s theorem [Th97, Theorem 2.1].∎
Next, we show that any dualizable category is -compactly generated. This follows from the following result which is (to the best of our knowledge) due to H. Krause.
Proposition 1.20.
Let be a functor between small stable -categories. Then the category is -compactly generated, and its -compact objects are exactly of the form
Proof.
This is a special case of [Kr10, Theorem 7.4.1].∎
Remark 1.21.
Note that the statement of [Kr10, Theorem 7.4.1] implies that for any uncountable regular cardinals the (non-full) inclusion commutes with fibers (i.e. with kernels). It is easy to deduce that this inclusion commutes with all -small limits.
Indeed, it is sufficient to prove this for fiber products and for -small products. Given -strongly continuous functors and between -presentable categories, we can first form the lax fiber product The functor is -accessible, hence by Proposition 1.2 the category is -presentable. The natural functor preserves -compact objects. Hence, by [Kr10, Theorem 7.4.1] its kernel is -compactly generated by the -compact objects of Finally, if is a -small collection of -presentable categories, then the category is also -presentable, and
Corollary 1.22.
Any dualizable category is -compactly generated. In particular, is naturally a full subcategory of (and a non-full subcategory of ).
Corollary 1.23.
For a dualizable category the essential image of the functor is contained in the full subcategory
Proof.
Since is -compactly generated, the colimit functor factors through
Here the first functor is the right adjoint to the inclusion. Hence, the functor factors through ∎
Remark 1.24.
In general, let be a dualizable category and is a (strictly) full stable subcategory. Then the essential image of is contained in if and only if the functor is a quotient functor. This condition is much stronger than the condition that generates by colimits.
For example, let be a field and consider the -dimensional -algebra and let be the stable subcategory generated by – -linear dual of Then generates by colimits, but the functor is not a quotient functor.
We give a proof that any compactly assembled category (not necessarily cocomplete) is -compactly generated.
Proposition 1.25.
Let be an accessible compactly assembled category. Then is -accessible. Moreover, the -compact objects of are exactly of the form where for each the morphism factors through
Proof.
Suppose that is a sequential colimit as above. Then
Hence, the object is -compact in Since is fully faithful and commutes with filtered colimits, we see that is -compact in Conversely, suppose that is -compact. Since the right adjoint to commutes with filtered colimits, we see that is -compact in Hence, for some sequence Then hence for each there is some such that the map factors through Passing to a subsequence, we may assume that as required.
It remains to show that any object of is an -directed colimit of -compact objects. Let and where is directed. Let be the poset of order-preserving maps Then is directed and
Let be the subset of maps such that for each the map factors through Arguing as above, we see that the inclusion is cofinal, in particular is directed.
Denote by the poset of equivalence classes where if for The partial order on is the following: given elements with representatives we have if for Then the poset is -directed, and the map is cofinal. Moreover, the functor factors uniquely through Therefore, we have
This shows that is an -directed colimit of -compact objects of ∎
Remark 1.26.
One can show that the category is -presentable. More precisely, it is generated by colimits by a single -compact object, given by the poset with the usual order, see Appendix D, more precisely, Remark D.4. In fact, the category shares a lot of good properties with the category However, in this paper we will mostly restrict our attention to dualizable categories, i.e. compactly assembled stable categories.
1.5. Examples of compactly assembled categories, stable and non-stable
We first give a few examples of dualizable (=compactly assembled stable) categories.
Example 1.27.
Let be a quasi-compact quasi-separated scheme. By [BVdB03, Theorem 3.1.1] the derived category of complexes of -modules with quasi-coherent cohomology (considered as a stable -category) is compactly generated, hence dualizable.
Example 1.28.
(Faltings [Fa02], Gabber-Ramero [GabRam03]) Let be an (associative unital) ring, and let be a two-sided ideal such that (we do not assume that or is flat as a left or right module). Then is a Serre subcategory. Denote by the Serre quotient Then the derived category is dualizable.
Indeed, the subcategory is (presentable and) closed under products and coproducts, hence the inclusion functor has both left and right adjoints. Since we have a short exact sequence in
where the functor is the left adjoint to the inclusion.
It is formal that we have a dg ring such that and is identified with In general (without flatness assumptions) we can only say that is connective (i.e. ), and (the non-derived tensor product). If is flat as a right (or left) -module, then we have More generally, if is flat as a right (or left) -module, then is the dg ring
Example 1.29.
Let be a locally compact Hausdorff space. Then the category of sheaves of spectra is dualizable (note that we do not require hyperdescent for sheaves). More generally, for any dualizable category the category is dualizable. Even more generally, if is a presheaf on with values in then the category is dualizable, see Section 6. We will prove in Section 6 that
where denotes the sheafification.
Example 1.30.
Let be a (for simplicity) noetherian commutative ring, and let be an ideal. Clausen and Scholze [CS20] defined the category of nuclear solid modules. This category is dualizable but not compactly generated (unless is nilpotent). The compact objects are given by the usual category of perfect complexes Moreover, the category is closely related with another category which can be defined as an inverse limit in
We refer to Subsection 1.15 for a discussion on general inverse limits in We will study both categories and in great detail in the forthcoming paper [E1]. In particular, we will prove that there is a natural fully faithful strongly continuous functor which induces equivalences on compact objects, and we have
We now give some examples of compactly assembled non-stable categories.
Example 1.31.
Given an associative unital ring the category of flat right -modules is compactly generated by Lazard’s theorem, hence it is compactly assembled.
Example 1.32.
Let be a poset considered as a category. Then is compactly assembled if and only if is a continuous poset, see Subsection 5.1 for the characterization of such posets. For example, the poset with the usual order is continuous, and we have Another example is the poset of open subsets of a locally compact Hausdorff space. In this case we have
Example 1.33.
The following example might be new (to the best of our knowledge).
Example 1.34.
Denote by the category of prenormed -vector spaces, where the morphisms are given by contractive linear maps (i.e. the maps such that ). Then is presentable and compactly assembled. Indeed, the category has colimits, and it is generated by colimits by finite-dimensional normed vector spaces and by the object (-dimensional vector space with zero prenorm). We have
and for a finite-dimensional normed space we have
1.6. Continuous stabilization
We describe the left adjoint to the (non-full) inclusion If is an accessible -category with filtered colimits, we denote by the (presentable stable) category of cofiltered limit-preserving functors If is -accessible, we can identify the category with the category of functors which commute with -small cofiltered limits. The inclusion commutes with limits and -filtered colimits, hence it has a left adjoint
For we denote by the “representable” presheaf, given by For any presentable stable category we have an equivalence
Proposition 1.35.
Let be a compactly assembled accessible category. Then the category is dualizable. The left adjoint to the inclusion is given by
Proof.
By Proposition 1.25 the category is -compactly generated. If is a regular cardinal, then the left adjoint to the inclusion has an explicit description:
We observe that the (a priori partially defined) functor is defined on objects of the form and is given by
Indeed, if is an ind-system in then applying (1.6) we get
Since the category is generated by the objects we conclude that is dualizable and the functor is strongly continuous.
By the above, if is dualizable, then a continuous functor is strongly continuous iff we have This exactly means that the functor is strongly continuous. This proves the proposition.∎
Below we will be especially interested in the special case of Proposition 1.35 for the poset considered as a compactly assembled category.
Remark 1.36.
Note that if is an accessible -category with filtered colimits, and is a presentable stable category, then the tensor product is identified with the category of cofiltered limit-preserving functors This follows from (1.5) and from the equivalence
1.7. Duality as a covariant involution on dualizable categories
Note that the category has a covariant involution We observe that also has a covariant involution so that for
Namely, given a strongly continuous functor between dualizable categories, we see that the dual functor has a left adjoint More precisely, we have where is the right adjoint to
This gives a well-defined functor which is naturally an involution. In particular, the functor commutes with all limits and colimits in
Remark 1.37.
In fact, the involution can be naturally extended to an involution This follows from the following statement, which we will address elsewhere.
Denote by the non-full subcategory of formed by dualizable objects; the -morphisms in are functors which have a colimit-preserving right adjoint. We claim that there is a natural equivalence This equivalence sends a compactly assembled accessible -category to the category of cofiltered limit-preserving functors The inverse functor sends a dualizable (non-stable) category to the category of left exact colimit-preserving functors
Remark 1.38.
Explicitly, the involution can be described as follows. Given the category is the full subcategory formed by pro-objects where is codirected, such that for any there exists such that the map in factors through
For example, if is a locally compact Hausdorff space, then – the poset of compact subsets of with the reverse inclusion order.
1.8. Dualizability via compact morphisms
We give a criterion for dualizability of a presentable stable category in terms of the homotopy category At least conceptually, most of this subsection is covered by the results of [Kr00, Kr05], applied to the smashing subcategory (assuming that is dualizable). In particular, the ideal of compact maps corresponds to this smashing subcategory as in [Kr00, Theorem A] and [Kr05, Theorem 1]. However, the criterion for dualizability does not follow formally from the results in loc.cit.
Let be a presentable stable category. We give the following definition.
Definition 1.39.
A morphism in is called compact if for any ind-object and for any morphism the composition factors through some
Remark 1.40.
1) Definition 1.39 also makes sense for non-stable presentable categories. However, in this section we consider only stable categories.
2) For an object compactness of the identity map is equivalent to the compactness of
The following observation is essentially the motivation for Definition 1.39.
Proposition 1.41.
Suppose that is dualizable. The following are equivalent for a morphism in
is compact.
The map in factors through
Proof.
. This follows from the definition of compactness, applied to the identity map
. For any ind-object and a map we can identify with the map Composing with the map we get an element of It is represented by some map The composition is homotopic to Hence, is compact. ∎
Corollary 1.42.
In a dualizable category any compact morphism is (homotopic to) a composition of two compact morphisms.
Proof.
Let be a compact morphism, and choose its lift Let where is filtered. Since commutes with colimits, we have Hence, the morphism factors through some We obtain compact morphisms and their composition is ∎
Proposition 1.43.
Let be a morphism in a presentable stable category The following are equivalent.
is compact.
for any ind-object such that the map of abelian groups
is zero.
Proof.
. Suppose that where is directed, and take some map Dnote by the identity map
Since is compact, the composition factors through some object It follows that the composition is zero, as required.
. Consider an arbitrary ind-object where is directed. Take a map Put Then Pick some and consider the composition
By (ii) for some the composition is zero. It follows that the composition factors through Hence, is compact. ∎
Proposition 1.44.
Let be a morphism in a presentable stable category Suppose that is -compact. The following are equivalent.
is compact.
for any sequence of objects in the image of the map
is contained in
Proof.
. This is clear: apply the definition of compactness to the sequence
. Consider a directed system and a map Since is -compact, we may assume that Indeed, -compactness of implies that for some sequences of elements of the map factors through
Assume Recall that we have a cofiber sequence
Consider the composition
By (ii) the map factors through a finite direct sum Replacing the sequence by its tail we may and will assume that is a zero map.
Hence, can be lifted to a map By (ii) the composition factors through a finite direct sum It follows that factors through hence is compact. ∎
We can now give a criterion of dualizability on the level of homotopy categories. Proposition 1.44 justifies the following definition.
Definition 1.45.
Let be a well generated triangulated category, and let A morphism is called compact if for any sequence of objects in and for any morphism the composition factors through a finite direct sum for some
Recall that in a triangulated category with countable coproducts we have sequential homotopy colimits, defined up to a non-canonical isomorphism:
Theorem 1.46.
Let be a presentable stable category. Then is dualizable if and only if the triangulated homotopy category satisfies the following condition:
the category is generated (as a localizing subcategory) by a set of objects such that each is isomorphic to a homotopy colimit of a sequence where each is -compact and each map is compact.
Proof.
Suppose that is dualizable. Take some Then is contained in so for some sequence of objects of Since commutes with colimits, we have
It follows that for each there is some such that the map factors through By Proposition 1.41 this means that the map is compact. Passing to a subsequence, we may assume that all maps are compact. Hence, the category satisfies we can take the collection of all -compact objects.
Now suppose that satisfies To show that is dualizable, it suffices to observe the following.
Claim. For any sequence in such that each map is compact, the ind-object is in the left orthogonal to
Proof of Claim..
Suppose that an ind-system satisfies Then each map of spectra
induces zero maps on the homotopy groups. Hence,
This proves the claim. ∎
It follows that the (a priori partially defined) functor is defined on each Since the objects generate it follows that is defined on the whole category ∎
The strong continuity of functors between dualizable stable categories can also be interpreted in terms of compact morphisms in the homotopy categories.
Proposition 1.47.
A continuous functor between dualizable presentable categories is strongly continuous if and only if the functor preserves -compact objects and compact maps between them.
Proof.
Suppose that is strongly continuous. Then preserves -compact objects and It follows from Proposition 1.41 that preserves all compact maps, in particular, preserves compact maps between -compact objects.
Suppose that preserves -compact objects and compact maps between them. By the proof of Theorem 1.46 we see that for each -compact object the map is an isomorphism. Since is -compactly generated, we conclude that is strongly continuous. ∎
Finally, given a dualizable category and two objects we discuss the relation between the following two abelian groups:
the subgroup of which consists of compact maps;
the abelian group
We recall the general notions of an ideal and a quasi-ideal in a small additive category. Given small additive categories and a left resp. right -module is an additive functor resp. and an -bimodule is a biadditive bifunctor Recall that an ideal of a small additive category is a subbimodule of the diagonal bimodule More generally, a quasi-ideal in is a bimodule together with a map (not necessarily injective) such that for any we have
We have the usual notion of a tensor product of left and right -modules. In particular, given bimodules we have a bimodule given by Note that the tensor product of quasi-ideals is naturally a quasi-ideal. We say that an bimodule is flat on the right if each right -module is flat,
Given a quasi-ideal we have a natural map We say that a quasi-ideal is idempotent if is an isomorphism.
Lemma 1.48.
Let be a small additive category. The assignment gives a bijection
The inverse map sends an ideal to the quasi-ideal where the map is given by the multiplication.
Proof.
Let be an idempotent quasi-ideal in and let It is clear that We need to show that as quasi-ideal. Note that if are elements such that or then It follows that the isomorphism factors through hence
Conversely, let be an ideal in such that and is flat on the right. We need to show that is idempotent. It suffices to show that the map
is an isomorphism. Since the map (1.7) is surjective. On the other hand, since is flat on the right, the map (1.7) is injective. Hence, this map is an isomorphism. ∎
Proposition 1.49.
Let be a dualizable category, and let be its homotopy category. Denote by the ideal of compact maps in Consider the bifunctor
Then is an idempotent quasi-ideal in which is flat on the right. Moreover, corresponds to via the bijection from Lemma 1.48. In particular, we have an isomorphism of quasi-ideals
Proof.
Remark 1.50.
We remark that both Lemma 1.48 and Proposition 1.49 hold in the non-additive resp. non-stable context. Namely, if is any (discrete) small category, then we can consider the (bi)functors with values in the category of sets instead of abelian groups. The tensor product and flatness are defined in the usual way, and the notion of a (quasi-)ideal is the same. The statement of Lemma 1.48 still holds, and the proof is the same.
Remark 1.51.
Another generalization of Proposition 1.49 is the following. Let be a compactly generated triangulated category, and consider a smashing subcategory This means that is a full triangulated subcategory closed under coproducts, and the inclusion functor has a right adjoint which commutes with coproducts. Then we have a quasi-ideal in given by Moreover, is idempotent and flat on the right. The image of in is the ideal of maps which factor through an object of We have and we obtain an isomorphism of quasi-ideals Proposition 1.49 is a special case of this statement with and
1.9. The axiom (AB6)
Definition 1.52.
Let be an -category with infinite products and filtered colimits. We refer to the following condition as the axiom (AB6):
- for any collection of filtered categories and for any functors the map
is an isomorphism.
Originally the axiom (AB6) was formulated by Grothendieck [Gro57] for abelian categories, but it makes sense in any -category with infinite products and filtered colimits. The most important observation is that (AB6) holds in the category of sets, hence also in the -category of spaces. We deduce the following statement.
Proposition 1.53.
Let be an -category with infinite products. Then the category also has infinite products, and the following isomorphism holds:
where are filtered categories, and we have functors
Proof.
Since (AB6) holds in the -category of spaces, for any object we have a chain of isomorphisms:
This proves the proposition.∎
We obtain another criterion of dualizability, which is due to Clausen and Scholze.
Proposition 1.54.
(Clausen, Scholze) Let be a presentable stable category. Then is dualizable if and only if (AB6) holds in
Proof.
Suppose that is dualizable. Then the colimit functor has a left adjoint Hence, the functor commutes with infinite products. By Proposition 1.53, this exactly means that (AB6) holds in
Conversely, suppose that holds in Arguing as above, we see that the functor commutes with infinite products. Now, the category is not presentable, so we need the following additional observation to deduce the existence of the left adjoint to
Let be a regular cardinal such that the category is -compactly generated. Denote by the colimit functor. Denoting by the right adjoint to the inclusion, we see that the natural transformation is an equivalence of functors Since both functors and commute with infinite products, we deduce that the functor also commutes with infinite products. Hence, it has a left adjoint Then the composition is left adjoint to hence is dualizable.∎
Remark 1.55.
The same argument shows that a (not necessarily stable) presentable category is compactly assembled if and only if it satisfies (AB6) and the strong version of (AB5): filtered colimits commute with finite limits. If is a complete lattice, then this statement is well known, see [GHKLMS80, Theorem I.2.3].
1.10. Automatic dualizability and images of functors
We have the following sufficient condition for dualizability.
Proposition 1.56.
Let be a presentable stable category, and consider a family of strongly continuous functors where each is dualizable. Suppose that is generated by the images of (by colimits). Then is dualizable.
Proof.
We first reduce to the case when consists of a single element. Consider the category Then is dualizable: it is sufficient to prove this for compactly generated categories, and in this case we have By Proposition 1.7, the induced functor is strongly continuous and its image generates
So assume that and put We give two proofs.
First proof. Note that the right adjoint functor is conservative (since the image of generates), and it commutes with products and filtered colimits. Since the axiom (AB6) holds in we deduce that it also holds in By Proposition 1.54, is dualizable.
Second proof. Let be a generating object, and Then for an ind-object we have an equivalence
Hence, the (a priori partially defined) functor is defined on Since the objects generates by assumption, we conclude that the functor exists, hence is dualizable.∎
An important corollary is that we have well-behaved images of -morphisms in
Definition 1.57.
Let be a continuous functor between presentable stable categories. We denote by the localizing subcategory generated by the image of Equivalently, is the fiber of the cofiber of in
Corollary 1.58.
Let and be presentable stable categories, and suppose that is dualizable. Let be a strongly continuous functor. Then the category is dualizable, and the inclusion is strongly continuous.
Proof.
Remark 1.59.
In the assumptions of Corollary 1.58, if is dualizable, then is also the fiber of the cofiber of in
Remark 1.60.
Note that we have a (not very explicit) description of the functor right adjoint to the inclusion. Let be a presentable stable category, and denote by the poset of (strictly full) dualizable subcategories such that the inclusion is strongly continuous. Note that is indeed a small set: any such is generated by -compact objects of Moreover, is a complete lattice: the supremum (join) of a collection is given by the image of the functor In particular, has the largest element Then is exactly the image of under the functor
1.11. Calkin categories
Fix an uncountable regular cardinal We will define the Calkin and -Calkin categories as follows.
Given a small stable Karoubi complete category we put
It turns out that this construction has a direct analogue for dualizable categories.
Let be a dualizable category. By Corollary 1.22, is -compactly generated. By Corollary 1.23 the image of the functor is contained in the full subcategory Hence, we have a fully faithful strongly continuous functor
Definition 1.61.
For a dualizable presentable stable category we define its continuous -Calkin category as the category of compact objects in the quotient We have a short exact sequence
in
We have fully faithful functors for We denote by the colimit of
We observe that this Calkin construction is functorial: we have a well-defined functor Indeed, if is a strongly continuous functor between dualizable categories, then we have a commutative square
Hence, we have a well-defined functor
Proposition 1.62.
If is compactly generated, we have a natural equivalence
Proof.
Indeed, in this case the functor is identified with Applying to the short exact sequence
we obtain an equivalence ∎
Remark 1.63.
Note that if is compactly generated, then the essential image of the functor is a stable subcategory: it is equivalent to the Verdier quotient However, in general this essential image does not have to be a stable subcategory, see Proposition B.1 below.
Proposition 1.64.
The functor takes short exact sequences to short exact sequences.
Proof.
The only assertion which requires a proof is the following: if is a strongly continuous fully faithful functor between dualizable categories, then the functor is fully faithful. Passing to ind-completions, it suffices to check that the functor is fully faithful. But this is the same as the functor
The latter functor is fully faithful, because the functor is fully faithful.∎
We conclude this subsection with the description of morphisms in the Calkin category.
Proposition 1.65.
Let be a dualizable category. The category is identified with the essential image of the functor In particular, for we have
Proof.
This follows directly from the definitions. Namely, the left adjoint to the inclusion is given by the left Kan extension of the functor
Note that if for some -ring then for any right -modules and Proposition 1.65 gives an isomorphism
1.12. Colimits of dualizable categories
We first observe the following.
Proposition 1.66.
The full subcategory is closed under colimits. In particular, the category is cocomplete and the functor commutes with colimits.
Proof.
The second statement follows from the first one by Proposition 1.7.
We need the following general facts about the directed colimits in and
Proposition 1.67.
Let be a directed system of presentable stable categories, and We denote by the transition functors, and by the structural functors to the colimit. Also, denote by and the corresponding right adjoint functors.
1) We have
2) Suppose that the functors are strongly continuous, i.e. the functors are continuous. Then for any we have an isomorphism
Proof.
We use once again the natural equivalence where the transition functors in the inverse system are Moreover, the composition is identified with the functor
To prove 1), we simply observe that for we have isomorphisms
Proposition 1.68.
The categories and satisfy the weak (AB5) axiom: the class of fully faithful functors is closed under filtered colimits. In particular, a filtered colimit of short exact sequences in is again a short exact sequence.
Remark 1.69.
We explain in Appendix A that in each of the categories and the monomorphisms are exactly the fully faithful functors.
Proof of Proposition 1.68.
By Proposition 1.66, it is sufficient to prove the statement for
Consider a filtered system of strongly continuous fully faithful functors where the transition functors and are strongly continuous. Put and denote by the colimit of Denote by the structural functors. For we can directly compute the composition Namely, applying Proposition 1.67 together with strong continuity and fully faithfulness of we obtain
Hence, i.e. is fully faithful. ∎
Corollary 1.70.
Let be a presentable stable category, and suppose we have a directed family of presentable stable full subcategories where the inclusion functors are strongly continuous. If is generated by as a localizing subcategory, then the natural functor is an equivalence.
Proof.
By Proposition 1.68, the functor is fully faithful. By assumption, is generated by so we actually have an equivalence.∎
Remark 1.71.
The following example shows that the analogues of Proposition 1.68 and 1.70 fail in Denote by the -th prime number, i.e. Put and where We consider as a full subcategory of via restriction of scalars, and denote by the inclusion functor. Hence, we have
Note that the subcategories generate as a localizing subcategory. Indeed, we have and However, we claim that the functor is not an equivalence.
Indeed, otherwise by Proposition 1.67 1) we would have But this clearly fails:
Hence, is not an equivalence.
We have the following important observation about compact objects.
Proposition 1.72.
The functor commutes with filtered colimits.
Proof.
Consider an ind-system of dualizable categories. By Proposition 1.68, we have a short exact sequence
It follows that
Here we used the fact that in filtered colimits commute with finite limits.∎
The following example shows that Proposition 1.72 does not hold for presentable categories.
Proposition 1.73.
Let be a field. Denote by the rectangle (considered as a topological space). Denote by the full subcategory of sheaves whose pullback to vanishes. Then the pullback functors are strongly continuous, the colimit is compactly generated, but we have for
In particular, the functor does not commute with sequential colimits.
Proof.
It is clear that each functor is strongly continuous. Identifying the colimit with the limit of the inverse sequence, we see that is equivalent to the full subcategory of consisting of sheaves whose support is contained in for all (and whose pullback to vanishes). This is simply the category of sheaves supported at the point which is equivalent to
It remains to show that Since all the categories are equivalent, we may assume Suppose that is a non-zero compact object. For put The pullback functor is strongly continuous, hence is a compact sheaf on Since is contractible and compact, it follows from Proposition 6.14 that is a constant sheaf. Since this holds for arbitrarily small and since we deduce that is a non-zero constant sheaf on . The pullback functor is strongly continuous, hence is also a compact sheaf. But is not compact, a contradiction with Proposition 6.14. ∎
We observe that the Calkin construction commutes with sufficiently filtered colimits.
Proposition 1.74.
The functor commutes with -filtered colimits. In particular, the functor commutes with -filtered colimits.
Proof.
Recall that the functor commutes with -filtered colimits. Hence, so is the composition It follows that the functor commutes with -filtered colimits. Passing to compact objects, we deduce that the functor commutes with -filtered colimits.∎
We conclude this subsection with the following elementary statement about pushouts.
Proposition 1.75.
In each of the categories and the class of fully faithful functors is closed under pushouts.
Proof.
It suffices to prove this for Passing to the right adjoints, we deduce this from the standard statement: a pullback of a quotient functor is a quotient functor.∎
1.13. Semi-orthogonal decompositions
We have the natural notion of a semi-orthogonal decomposition in and in
Definition 1.76.
A (finite) semi-orthogonal decomposition in (resp. ) of a presentable stable (resp. dualizable) category is a collection of presentable stable (resp. dualizable) subcategories such that
i) the inclusion functors are strongly continuous;
ii) For we have
iii) The categories generate as a stable subcategory.
In this case we write
Remark 1.77.
Let and be strongly continuous functors between presentable stable categories. Then and define a semi-orthogonal decomposition in if and only if the following conditions hold:
(resp. ) has a strongly continuous left (resp. right) adjoint (resp. );
we have and
we have (equivalently, );
we have a cofiber sequence in
This is a straightforward generalization of a similar statement for small stable categories. Note that the strong continuity of follows from the cofiber sequence in the last condition.
The following observations are immediate.
Proposition 1.78.
Let be a semi-orthogonal decomposition in Then is dualizable if and only if each is dualizable. In this case we also have a semi-orthogonal decomposition on the level of compact objects:
Proof.
If is dualizable, then each is a retract of in hence is dualizable. Conversely, if each is dualizable, then is dualizable by Proposition 1.56.
For the second statement we may assume that Denote by the inclusion functors. Then the functors and satisfy the version of conditions of Remark 1.77 for small stable categories. Hence, they induce a semi-orthogonal decomposition ∎
The usual construction of a semi-orthogonal gluing via a bimodule works in the context of dualizable categories.
Proposition 1.79.
Let be an accessible functor between dualizable categories, and consider the semi-orthogonal gluing with the inclusion functors The following are equivalent.
the category is dualizable and the functors and are strongly continuous.
the functor is continuous.
Proof.
Note that the functor is automatically strongly continuous.
. It suffices to recall that where is the right adjoint to
. Suppose that is continuous. Note that the functor is given by
Hence, is strongly continuous. The dualizability of follows from Proposition 1.78.∎
We have the following expected statement about the functors to and from a dualizable category with a semi-orthogonal decomposition.
Proposition 1.80.
Let be a presentable stable category with a semi-orthogonal decomposition in Denote by the left adjoint to and the right adjoint to Let be another presentable stable category.
1) A continuous functor is strongly continuous if and only if the compositions and are strongly continuous. We have a semi-orthogonal decomposition
2) A continuous functor is strongly continuous if and only if the compositions and are strongly continuous. We have a semi-orthogonal decomposition
Proof.
We first prove part 2). The criterion of strong continuity follows from Proposition 1.9 applied to the functors The functors and are given respectively by precomposition with and The left adjoint to is given by precomposition with and the right adjoint to is given by precomposition with It follows formally that and are fully faithful and the composition is zero. It remains to observe that the cofiber sequence in implies that we have a cofiber sequence in
We now prove part 1). The proof of a semi-orthogonal decomposition is similar; here the inclusion functors are given by the compositions with and The “only if” direction of the criterion is clear. To prove the “if” direction we note that the functors and are strongly continuous, and the cofiber sequence implies that is also strongly continuous. ∎
For a future reference, we define the notion of a poset-indexed (possibly infinite) semi-orthogonal decomposition.
Definition 1.81.
Let be a partially ordered set, and let be a presentable stable (resp. dualizable) category. An -indexed semi-orthogonal decomposition of in (resp. in ) is a collection of presentable stable (resp. dualizable) subcategories such that
i) the inclusion functors are strongly continuous;
ii) For we have
iii) The categories generate as a localizing subcategory.
In this case we write
Proposition 1.82.
Let be a poset and let be a semi-orthogonal decomposition in Then is dualizable iff each is dualizable. In this case we have a semi-orthogonal decomposition
Proof.
The first statement is proved in the same way as Proposition 1.78.
To prove the second statement, we introduce for each finite subset the subcategory generated by By Propositions 1.56 and 1.9 the category is dualizable and the inclusion is strongly continuous. Moreover, we have a semi-orthogonal decomposition By Proposition 1.78, the category is generated as a stable category by Combining Corollary 1.70 and Proposition 1.72 we deduce that the category is generated as a stable subcategory by Hence, we have a semi-orthogonal decomposition as required.∎
1.14. Finite limits of dualizable categories
The limits in the category of dualizable categories are much more complicated then colimits. We start their study with the case of kernels (i.e. fibers).
Proposition 1.83.
Let be a dualizable stable category. For any there exists a functor such that
for any we have
for any the morphism is compact.
Proof.
It is convenient to replace the poset with the isomorphic poset
Choose a sequence in such that and each map is compact. Define the functor by the formula Since any compact morphism in is a composition of two compact morphisms, we can define inductively the functors such that and for any such that the map is compact. The sequence defines a functor
Finally, we define a functor by the formula Then satisfies the required properties (since is a dense linearly ordered set).∎
The following lemma will be useful.
Lemma 1.84.
Let be a dualizable pesentable stable category, and be a full stable subcategory, closed under coproducts. Then there exists the largest dualizable subcategory such that the inclusion functor is strongly continuous.
Moreover, is generated by the objects of the form where is a functor such that for any rational numbers the map is compact in
Proof.
Denote by the full subcategory generated (as a localizing subcategory) by the objects described in the second assertion. We first prove that is dualizable and the inclusion is strongly continuous.
Consider a functor as above. We may and will assume that for each rational the map is an equivalence. Indeed, as in the proof of Proposition 1.83 we may replace by the functor given by
Now, we observe that our assumptions on imply that each object is contained in (because the linearly ordered set is isomorphic to ). It follows that the functor exists and we have Moreover, the inclusion commutes with hence it is strongly continuous.
Now, let be some full dualizable subcategory, such that the inclusion is strongly continuous. We claim that is contained in To see this, take any countably presented object Choose a functor such that each map is compact in (hence in ) and We see that as required.∎
We can give a description of kernels in
Proposition 1.85.
Let be a strongly continuous functor between dualizable categories. Denote by the usual kernel of (the fiber in ). Then the dualizable kernel of (i.e. the fiber in ) is given by the largest dualizable full subcategory such that the inclusion is strongly continuous. More precisely, can be described as in Lemma 1.84.
Proof.
We denote by the category described in Proposition 1.85.
Example 1.86.
Let be a discrete valuation ring with the maximal ideal and with the field of fractions Then Indeed, the usual kernel is given by but the inclusion functor is strongly continuous. The only localizing subcategory of which is not equal to is zero.
Corollary 1.87.
Sequential colimits in do not commute with finite limits. In particular, the category is not compactly assembled.
Proof.
Take some prime number and denote by the ring of integers of the cyclotomic extension Then On the other hand, the colimit is a non-discrete valuation ring, hence
Hence, sequential colimits in do not commute with kernels. This implies that is not compactly assembled.∎
We have the following general statement about fiber products in
Proposition 1.88.
Consider a pair of strongly continuous functors between dualizable categories
1) The fiber product exists in and the natural functor is fully faithful and strongly continuous. More precisely, the category is the largest dualizable subcategory such that the functors and are strongly continuous.
2) If moreover is a localization, then we have and the functor is a localization.
Proof.
1) By Proposition 1.79 the lax fiber product is dualizable. It follows from Proposition 1.80 that where The usual fiber product (taken in ) is simply The rest follows from Proposition 1.85.
2) It is well-known that in a pullback of a localization is a localization. To show that is also the fiber product in it suffices to show that the functor is a localization (assuming that is a localization). This is standard. Explicitly, the right adjoint is given by and ∎
We will need the following statement about the class of situations when the fiber product of compactly generated categories is automatically compactly generated.
Proposition 1.89.
Let be a homological epimorphism of small stable idempotent-complete categories. Then we have equivalences
Proof.
By Proposition 1.88 we only need to show that the category is compactly generated. It suffices to show that the objects of the form are generated by colimits by the essential image of the diagonal functor
By Proposition 1.20, we may assume that is of the form where and is a zero map in Then we have
where the maps are chosen (non-uniquely) in such a way that they induce the pairs of maps This proves the proposition.∎
1.15. General limits of dualizable categories
To describe the general limits in we will use the following non-standard adjunction statement.
Proposition 1.90.
Let be an uncountable regular cardinal. The (non-full) subcategory inclusion functor has a right adjoint, given by
Proof.
We construct the adjunction unit and counit as follows.
Let be a dualizable stable category. Then the adjunction unit on is given by the functor (we know that this functor is strongly continuous).
Let be an -presentable category. Then the adjunction counit is given by the colimit functor Note that this functor is -strongly continuous.
If is dualizable, then the composition is isomorphic to the identity functor. If is -presentable, then the composition
is also isomorphic to the identity. This proves the adjunction. ∎
Remark 1.91.
Given a dualizable category and an -presentable category an -strongly continuous functor corresponds via the above adjunction to the composition
Theorem 1.92.
Let be an -category, and be a functor, Then the limit of exists, and is given by
Proof.
By Proposition 1.90 we see that the limit exists and is given by Let be a dualizable category. Note that a strongly continuous functor corresponds to a strongly continuous functor if and only if the composition is zero. It follows formally that exists and it is the largest dualizable subcategory
such that the inclusion functor is strongly continuous.
Note that for any functor we have a strongly continuous fully faithful embedding To conclude, we simply observe that is equivalent to
and we have a factorization
where the latter functor is fully faithful.
This proves the theorem. ∎
1.16. Example: fpqc descent
To give a feeling of dualizable inverse limits, we prove the following statement. Here we denote by the category of commutatie rings.
Proposition 1.93.
1) The functor is a sheaf in the descendable topology.
2) The functor is a sheaf in the fpqc topology.
Recall that the descendable topology on is given by the pretopology in which the covers are the maps such that is a descendable -algebra, see [Mat16, Definition 3.18]. Namely, this means that the unit object is contained in the stable idempotent-complete ideal of generated by
We need the following lemma.
Lemma 1.94.
Let be a strongly functor between presentable stable categories, and denote by its right adjoint. Suppose that there exists such that each object of is a direct summand of an object with a finite filtration with at most non-zero subquotients, which are contained in Then for any regular cardinal the functor preserves and reflects -compact objects.
Proof.
Since is continuous, takes to
If is an object such that is -compact, then for any collection of objects we have
The assumption on can be reformulated as follows: for any infinite set the category is generated by the full subcategory as a stable idempotent-complete subcategory. We conclude that is -compact. ∎
Proof of Proposition 1.93.
1) Let be a map such that is descendable over Below all the tensor products over are always derived, including the tensor powers of i.e. they are taken in By [Mat16, Proposition 3.22] we have an equivalence By Lemma 1.94, it induces an equivalence By Theorem 1.92, we obtain
It suffices to prove the following statement:
the functor is fully faithful.
To prove recall that by Proposition 1.65 for the complex of morphisms between and in the Calkin category is given by
Since we know the descent for it suffices to show that the map
is an isomorphism. This is equivalent to proving the following
the map is an isomorphism.
This follows easily from the descendability of over Indeed, by [Mat16, Proposition 3.20] the inverse system is pro-constant, hence we have equivalences
2) Descent for finite disjoint unions is obvious. We need to prove that for a faithfully flat map we have As in the proof of 1), we only need to show that and to prove the fully faithfulness statement above.
By Proposition 1.95 below, is isomorphic to an -directed colimit where are descendable faithfully flat -algebras (and is an -directed poset). We know that the functors and commute with -filtered colimits (by Proposition 1.74). Also, the functor commutes with -filtered colimits, since the category is countable. We deduce from part 1) that and the functor is fully faithful. As in the proof of 1), this implies that ∎
The following result is due to Peter Scholze (private communication). We give a short proof based on Positselski’s results on accessibility [Pos23].
Proposition 1.95 (Scholze).
Let be a commutative ring. The category of faithfully flat -algebras is -accessible. A faithfully flat -algebra is -compact if and only if is a countably presented -module (equivalently, a countably presented -algebra).
In particular, any faithfully flat -algebra is an -directed colimit of descendable faithfully flat -algebras.
Proof.
The second assertion follows from the fact that a faithfully flat countably presented -algebra is descendable by [Mat16, Corollary 3.33].
By Lazard’s theorem, the category of flat -modules is equivalent to the ind-completion where is the category of finite projective -modules. The category of flat commutative -algebras is equivalent to the category of modules over the monad The functor commutes with filtered colimits and preserves -compact objects. Applying Proposition 1.98 below, we see that the category is -accessible, and a flat -algebra is -compact if and only if is a countably presented -module.
It remains to observe that if is a map of flat -algebras such that is faithfully flat over then is also faithfully flat over ∎
Remark 1.96.
It seems plausible that for any descendable map and for any uncountable regular we have an equivalence
However, we only checked the fully faithfulness statement (the proof for is the same). It is not clear how to prove the essential surjectivity even for
Recall the following -categorical notion.
Definition 1.97.
[AR94] Consider a pair of functors between ordinary categories, and let be a pair of natural transformations. The equifier of is the full subcategory formed by objects such that
Proposition 1.98.
Let be an uncountable regular cardinal, and let be a -accessible (ordinary) category which has filtered colimits. Let be a monad such that commutes with filtered colimits and preserves -compact objects. Then the category of -modules is -accessible, and a -module is -compact in if and only if its underlying object is -compact in
Proof.
Let be the category of pairs where Clearly, has filtered colimits and the forgetful functor commutes with filtered colimits. By [Pos23, Theorem 4.1] the category is -accessible, and preserves and reflects -compactness.
Consider as a natural transformation Then the category is equivalent to the joint equifier of the pairs and Applying [Pos23, Theorem 3.1], we conclude that the category is -accessible, and the forgetful functor to preserves and reflects -compactness. ∎
1.17. Products of dualizable categories
In this subsection we discuss the products in It turns out that they are very well-behaved: the functor commutes with products, and the category satisfies the weak (AB4*) axiom and the (AB6) axiom.
First, we make an observation about infinite products of ind-objects. Let be a family of small stable categories. Given ind-objects we can consider each as an object of We denote by the product of these objects in We denote by the resulting functor. This is simply the right adjoint to the natural functor We give two more equivalent descriptions of the functor
Proposition 1.99.
1) For a small stable category let us identify the category with the category of exact functors to Then for a collection of ind-objects the ind-object is identified with the functor
2) If are filtered categories such that then the ind-object is given by the formula
Proof.
Part 1) follows directly from the definition, and part 2) follows from Proposition 1.53.∎
Corollary 1.100.
Let be a family of functors between small stable categories. Then we have a commutative square of categories
Proof.
This follows immediately from Proposition 1.99, part 2). ∎
Given a small stable category we denote by the evaluation functor. It is identified with the functor
Proposition 1.101.
Let be a collection of small stable categories, and take some ind-objects Then the natural map
is an equivalence.
Proof.
Let us write where and are filtered categories and . Then applying Proposition 1.99 and (AB6) for we get a chain of equivalences
This proves the proposition. ∎
We now can deduce the statement about infinite products of homological epimorphisms.
Proposition 1.102.
Let be a family of homological epimorphisms between small stable categories.
1) The functor is a homological epimorphism.
2) For a family of objects the object is contained in if and only if each is contained in Moreover, the category is generated by the objects of this form.
Proof.
Put
1) Take some families of objects Since are homological epimorphisms, using Proposition 1.101, we obtain equivalences
Hence, is a homological epimorphism.
2) The first statement holds even without the assumption that are homological epimorphisms, it follows from Corollary 1.100. For the “moreover” statement, note that the category is generated by the objects of the form
But we have
and
This proves part 2). ∎
Now we can describe the products in the category
Proposition 1.103.
Let be a collection of dualizable categories.
1) We have a short exact sequence
2) If the categories are compactly generated, then we have an equivalence
In other words, the functor commutes with products.
Proof.
By Theorem 1.92, we have an equivalence
Now part 1) follows from Proposition 1.102, applied to the homological epimorphisms
Part 2) follows from the fact that an infinite product of short exact sequences in is a short exact sequence, applied to the sequences
Now we want to prove that an infinite product of short exact sequences in is again a short exact sequence. Here the non-trivial part is the weak version of (AB4*) for an infinite product of epimorphisms (quotient functors) in is an epimorphism. We need the following observation.
Lemma 1.104.
1) Let be a collection of dualizable categories. We have a fully faithful strongly continuous functor
The image of contains all objects of the form where and is the evaluation functor for Moreover, the image of is generated by the objects of this form.
2) Let be a family of strongly continuous functors between dualizable categories. Then the following square commutes:
Here the functors and are as in 1). The functor (resp. ) is (resp. ).
Proof.
1) The functor is obtained by applying to the inclusion in the short exact sequence (1.9). Note that the (fully faithful) functor takes to the functor The assertion now follows from Proposition 1.102, part 2), applied to the homologicl epimorphisms
2) We need to show that the image of is contained in the image of By part 1), it suffices to show that for any the object is contained in the image of We have
By part 1), the latter object is in the image of This proves the proposition.∎
We can now deduce that the weak (AB4*) holds in the category
Proposition 1.105.
Consider a collection of short exact sequences in Then the sequence
is also short exact.
Proof.
The only statement that requires a proof is that the functor is a quotient functor. Equivalently, it suffices to show that the functor is a quotient functor, i.e. its right adjoint is fully faithful. By Lemma 1.104, the latter functor is the upper horizontal arrow in a commutative square
Here both vertical arrows are fully faithful, and so is the lower horizontal arrow. Hence, the upper horizontal arrow is also fully faithful, as required.∎
It turns out that the (AB6) axiom holds in the category itself.
Proposition 1.106.
The axiom (AB6) holds in That is, for any filtered categories for any functors the natural functor
is an equivalence.
Proof.
We first observe that (AB6) holds in the category of small stable Karoubi-complete categories, since the category is compactly generated (one can also check (AB6) for directly, by showing that the comparison functor is fully faithful and essentially surjective). Since the functor commutes both with products and with filtered colimits, we deduce that (1.10) is an equivalence when all are compactly generated.
By Propositions 1.68 and 1.105, filtered colimits and products of short exact sequences in are short exact. Consider the map of short exact sequences
By the above discussion, the middle vertical and the right vertical arrows are equivalences. Hence, so is the left vertical arrow. This proves (AB6) for ∎
2. Dualizability is equivalent to flatness
In this section we give a non-trivial criterion of dualizability of a presentable stable category via a “flatness” condition. It can be skipped on the first reading.
Let be a presentable stable category. We can ask the following naive question: is it true that for any fully faithful colimit-preserving functor between presentable stable categories, the functor is also fully faithful?
Definition 2.1.
A category is called flat if the answer to the above question is affirmative, i.e. the tensor product preserves fully faithful functors (=monomorphisms).
We have the following surprising result.
Theorem 2.2.
Let be a presentable stable category. Then is flat if and only if it is dualizable.
The “if” direction is easy: dualizability of implies an equivalence To prove the “only if” direction, we need a certain “extra -functoriality” statement.
Denote by the -category of presentable stable categories and exact accessible functors. We have the following result.
Proposition 2.3.
Let be a presentable stable category. Then we have a natural oplax -functor which induces the usual Lurie tensor product when we restrict to continuous functors. Moreover, if is flat, then we get an actual -functor
Proof.
Givem presentable stable categories and let us denote by the -category of triples where and are fully faithful continuous functors, such that we have a semi-orthogonal decomposition
Then is in fact an -category, and it is naturally equivalent to the category of exact accessible functors Namely, given a triple the corresponding functor is given by where is the right adjoint to Conversely, to an exact accessible functor we associate the category as in Definition 1.1. Recall that is the category of triples where and The inclusions and are given respectively by and
Given correspondences and we can define their composition as follows. We form the pushout (in ), and take the full subcategory generated by the images of the compositions and Then is naturally an object of Moreover, this composition operation is compatible with the composition of the corresponding accessible functors.
Now, we observe that for any presentable and we have a well-defined functor Indeed, given we have Here the fully faithfulness of (resp. ) follows from the existence of a continuous right adjoint to (resp. a left adjoint to ).
Now, given and as above, we observe that we have a natural equivalence Thus, we have a natural transformation
in Moreover, if is flat, then (2.1) is an equivalence.
Identifying the correspondences with accessible functors, we obtain an oplax -functor If is flat, we get an actual -functor.
Finally, if the triple corresponds to a continuous functor then the functor is continuous. Hence, the functor is the right adjoint to and the object corresponds to the functor – the usual Lurie tensor product.∎
Proof of Theorem 2.2..
We already observed that dualizability implies flatness by identifying with
Now, suppose that is flat. Using Proposition 2.3 (extra -functoriality), we will deduce that (AB6) holds in By Proposition 1.54 this will imply that is dualizable.
Consider a family of filtered categories We define the following presentable (in fact, compactly generated) stable categories:
We have the following natural exact accessible functors:
- the product of the colimit functors
- the product functor;
- taking to the functor
- the colimit functor.
Since the axiom (AB6) holds in the category the natural transformation is an equivalence. So we have a commutative square
Applying the extra 2-functoriality of we obtain a commutative square
But it is straightforward to check that the latter square is of the form
Moreover, the functors in this commutative square are analogous to the functors with “” replaced by “”. Therefore, commutativity of (2.2) exactly means that (AB6) holds in ∎
3. Extensions of compactly generated categories
Consider a short exact sequence of presentable stable categories of the form
where and are small idempotent-complete stable categories, and the functors and are strongly continuous.
Question 3.1.
Is it true that the category is automatically dualizable?
It turns out that in general the answer is “no”. To give a precise criterion of dualizability of we introduce some notation.
Denote by the right adjoint to and by the right adjoint to Note that we have a semi-orthogonal decomposition Putting we obtain an equivalence
Note that we have equivalences of categories of functors:
We denote by the functor corresponding to
Recall the following full subcategories of formed by (elementary) Tate objects.
Definition 3.2.
[Hen17, Definition 1] We denote by the full subcategory which is generated by and as a stable subcategory. We denote by the Karoubi completion of
The following result gives a criterion when is dualizable.
Proposition 3.3.
Within the above notation, the following are equivalent.
is dualizable.
is compactly generated.
the image of is contained in
Assuming Proposition 3.3, we deduce the following result.
Theorem 3.4.
Let and be small stable Karoubi complete categories. Denote by the (large, locally small) space of short exact sequences in of the form
Then we have an equivalence
Proof.
Remark 3.5.
The special case of Theorem 3.4 for split extensions is well-known. Namely, short exact sequences of the form (3.2) where both functors have right (resp. left) adjoints correspond to semi-orthogonal decompositions of the form (resp. ), which in turn correspond to functors (resp. ), i.e. -bimodules (resp. -bimodules).
Proof of Proposition 3.3.
Clearly, .
If is dualizable, then for any object there exists an object such that . Such objects together with the objects of form a generating collection of compact objects of hence is compactly generated.
. We use the above observation: is compactly generated if and only if for any object there exists a compact object such that Via the identification the object must be of the form
We need to describe the necessary and sufficient conditions for to be a compact object of Note that commutes with colimits since is a compact object of We have
We deduce that the object is compact if and only if the right-hand side in (3.3) commutes with coproducts.
Note that the functor is exactly the object where corresponds to as above. We conclude that the existence of such that is equivalent to the existence of objects and and an exact triangle
By [Hen17, Theorem 3], this is equivalent to the inclusion By Thomason’s theorem the latter condition is equivalent to the inclusion This proves the equivalence of (ii) and (iii). ∎
Remark 3.6.
It follows from the proof of Proposition 3.3 that the universal extension of by is the following short exact sequence:
Here the functor is given by The functor is given by The functor is given by The category can be thought of as the category of elementary Tate objects of with a choice of a lattice. The functor then simply forgets the lattice.
Example 3.7.
Let be a prime number, and consider the short exact sequence
The associated -linear functor sends to where is considered as a Tate object in the usual way.
We give some natural examples of non-dualizable categories which are extensions of compactly generated categories. We use the notation for the direct sum of copies of an object indexed by the elements of a set
Proposition 3.8.
Let be a field. Consider the category where Equivalently, is the category of triples where and Then is not dualizable.
Proof.
Consider the functor corresponding to By Proposition 3.3, we need to show that
Identifying with the opposite of the category of -linear accessible functors we see that corresponds to the functor Using the fact that (AB6) holds in we obtain isomorphisms
Putting we obtain an isomorphism
To show that we need to show that the image of in is not a pro-constant object. Assume the contrary. Since the maps are split epimorphisms for we deduce that the images must form an eventually constant pro-system in But this is clearly false: for any the map is not an isomorphism in since ∎
Another application of Proposition 3.3 is the classification of dualizable localizing subcategories of where is a noetherian commutative ring (we do not require the inclusion functor to be strongly continuous). For a subset denote by the full localizing subcategory generated by the residue fields By [Nee92, Theorem 2.8] the assignment gives a bijection
Moreover, by [Nee92, Theorem 3.3] the inclusion is strongly continuous if and only if is closed under specialization. In this case the category is compactly generated by the Koszul complexes where are elements such that
Definition 3.9.
Given a poset and two elements we put We call a subset convex if for any we have
Theorem 3.10.
Let be a noetherian commutative ring and a subset. Then is dualizable if and only if is convex with respect to the specialization order. In this case the category is compactly generated.
Example 3.11.
Let is a field. Consider the following full subcategory of the derived category of quasi-coherent sheaves on the affine plane:
Then Theorem 3.10 implies that is not dualizable. Indeed, the corresponding subset contains the ideals and but does not contain the ideal
We need the following auxiliary statement.
Lemma 3.12.
Let be a noetherian local commutative ring, and a complex of -modules. Denote by the inclusion functor. Suppose that the functor commutes with coproducts. Then is a perfect complex of -modules.
Proof.
Denote by the injective hull of the -module Then is a locally -torsion -module, hence Moreover, is an injective cogenerator of the abelian category of -modules.
We first observe that has bounded cohomology. Indeed, suppose that for infinitely many integers For each choose a non-zero map (which exists since is a cogenerator). Since is injective, we obtain the map This map does not factor through a finite direct sum, which contradicts the commutation of with coproducts.
Next, we show that each -module is finitely generated. We may and will assume that Suppose that is not finitely generated. Then we can construct a strictly increasing sequence of submodules
such that each is finitely generated. Put For choose a non-zero map such that (where ). Choose some which extends for each The sequence gives a well-defined map which does not factor through a finite direct sum. Since is noetherian, the module is injective. Choosing an extension of we obtain a morphism which does not factor through a finite direct sum. This contradicts the commutation of with coproducts.
So we showed that has bounded finitely generated cohomology. To prove that is perfect, it suffices to observe that is perfect over by adjunction, the functor commutes with coproducts. ∎
Lemma 3.13.
Let be a local noetherian commutative ring without zero-divisors. Let be a proper subset such that and is closed under specialization. Then the category is not dualizable.
Proof.
Put and denote by the field of fractions of We have a short exact sequence
where and are strongly continuous, and the categories on the left and on the right are compactly generated.
Suppose that is a dualizable category. The proof of Proposition 3.3 shows that there exists an object and a map such that the functor
commutes with coproducts. Since denoting by the inclusion functor, we see that the functor commutes with coproducts. By Lemma 3.12, we deduce that But choosing some we have
a contradiction.∎
Proof of Theorem 3.10.
Suppose that is convex. Denote by the set of all points which are specializations of elements of Put Then and are closed under the specialization, hence the categories and are compactly generated and the inclusion is strongly continuous. Hence, the quotient is also compactly generated.
Now let be a non-convex subset. Suppose that the category is dualizable. Choose some such that and Since is noetherian, we may and will assume that is a minimal element of the set Denote by and respectively the prime ideals corresponding to the points and Put We identify the set with its image in and put
Since the category is dualizable, so is the category Moreover, since is dualizable, the functor
is fully fiathful. Its essential image is clearly the category which is therefore dualizable. But our choice of the ideals and implies that the subset satisfies the conditions of Lemma 3.13. Hence, the category is not dualizable, a contradiction.∎
Remark 3.14.
The “only if” part of Theorem 3.10 fails for non-noetherian rings. Namely, let be a field, and consider the rank valuation ring
Then where is the maximal ideal, and the field of fractions of is
Consider the non-convex subset Then we have
– a compactly generated category.
To see this, it suffices to observe that if is an -torsion -module, then acts by zero on Indeed, if is an element such that then It follows that for we have
Hence, the full subcategories and in are mutually orthogonal.
4. Localizing invariants
4.1. Universal localizing invariants
To avoid set-theoretic issues, we will consider only accessible localizing invariants with values in accessible stable categories.
Let be an accessible stable category with (not necessarily cocomplete). Recall that a functor is a localizing invariant if the following conditions hold:
for any fiber-cofiber sequence of the form
in the square
is a fiber sequence (equivalently, a cofiber sequence);
If is a regular cardinal such that has -filtered colimits, we denote by the category of localizing invariants which commute with -filtered colimits. It follows from [BGT13] that for any regular cardinal there exists a universal localizing invariant such that the category is accessible and has -filtered colimits, and for any other such category we have
Here the right-hand side is the category of exact functors which commute with -filtered colimits.
A clarification is required when is uncountable. In this case the construction of [BGT13, Section 8.2] gives a solution to a different universal problem: the authors construct a presentable stable category and a localizing invariant which commutes with -filtered colimits, such that for any presentable stable category we have
The following proposition describes a relation between these two constructions.
Proposition 4.1.
Let be an uncountable regular cardinal. The category is -accessible, and the category is compactly generated. We have an equivalence
Proof.
It suffices to show that if is an accessible category with -filtered colimits, then the category is equivalent to the category of partially defined localizing invariants (where we consider only short exact sequences of -compact categories). This follows from Lemma 4.2 below. ∎
Lemma 4.2.
Any short exact sequence in is an -filtered colimit of short exact sequenes of -compact categories.
Proof.
Consider a short exact sequence
in Choose an -directed system such that each is in and Consider a poset of pairs where and is a full stable idempotent-complete subcategory generated by at most countable collection of objects. By 1.6, such a category is automatically -compact. The poset is -directed, and we conclude that (4.1) is a -indexed colimit of short exact sequences
of -compact categories. ∎
Localizing invariants which commute with filtered colimits are called finitary. In this case we write
Recall that for each regular the category has a natural (closed) symmetric monoidal structure, and the functor is symmetric monoidal. Below we will often formulate results in terms of the universal localizing invariants. They automatically imply the corresponding results for general localizing invariants.
Remark 4.3.
In the forthcoming papers [E1, E2] we will study the categories and in great detail. In particular, we will prove that the category and its relative versions are rigid as a symmetric monoidal categories in the sense of Gaitsgory and Rozenblyum [GaRo17, Definition 9.1.2]. This in particular will imply that the category is dualizable, which is quite non-obvious from the definition. In this paper however we will use only the universal property of the categories
We recall the following standard notion of a -equivalence.
Definition 4.4.
Let be an exact functor between small stable idempotent-complete categories. We say that is a -equivalence if there exists an exact functor such that we have in and in
It is well-known that a -equivalence induces isomorphisms on all additive invariants, hence on all localizing invariants, see [BGT13].
4.2. Localizing invariants of dualizable categories
The notion of a localizing invariant is defined in the same way as for By Theorem C.6 the category is -presentable, and again we will consider only accessible localizing invariants. It turns out that any localizing invariant defined on small categories has a unique extension to a localizing invariant defined on dualizable categories, so that
Roughly the idea is the following. Proposition 1.62 shows that for a compactly generated presentable category the Calkin category is, up to direct summands, the quotient of by the full subcategory of compact objects For a general dualizable category one can think of as a “virtual quotient of by compact objects”. We will use this point of view to extend localizing invariants from to
Definition 4.5.
Let be a dualizable presentable -category. For any accessible stable category and for any accessible localizing invariant we define the functor by the formula
If is the -theory functor then we refer to the functor as continuous -theory.
Proposition 4.6.
Let and be as in Definition 4.5. The functor is a localizing invariant. In particular, is additive in semi-orthogonal decompositions.
Proof.
This follows immediately from Proposition 1.64.∎
Proposition 4.7.
If and are as in Definition 4.5, and then we have a natural isomorphism
More precisely, we have a natural isomorphism of functors
Proof.
This follows immediately from the short exact sequence (in )
and from the vanishing ∎
We have the following (expected) statement about filtered colimits.
Proposition 4.8.
Let be an accessible stable category with -filtered colimits, and a localizing invariant which commutes with -filtered colimits. Then the functor commutes with -filtered colimits.
Proof.
Corollary 4.9.
Continuous -theory of dualizable categories commutes with filtered colimits.
Summarizing, we can formulate the following theorem.
Theorem 4.10.
Let be a regular cardinal and an accessible stable category with -filtered colimits. The precomposition functor
induces an equivalence between the full subcategories and of localizing invariants which commute with -filtered colimits. The inverse functor is given by
Proof.
By Propositions 4.6 and 4.8, the assignment is indeed a functor By Proposition 4.7, it is a right inverse to the precomposition
On the other hand, for any and any dualizable category we have a short exact sequence
which is functorial in This gives an isomorphism of functors This proves the theorem.∎
Further, we have two straightforward results about commutation of with sufficiently nice pullbacks and pushouts.
Proposition 4.11.
Let be an accessible localizing invariant. Let be a pair of strongly continuous functors between dualizable categories, and assume that is a localization (i.e. a quotient functor). Then the fiber product taken in is also the fiber product taken in and the natural map is an equivalence.
Proof.
The first assertion is the part of Proposition 1.88. Also by loc. cit. the functor is a localization. Clearly, the natural functor is an equivalence. Applying we deduce the commutation of with this fiber product.∎
Proposition 4.12.
Let be an accessible localizing invariant. Let be a pair of strongly continuous functors between dualizable categories, and assume that is fully faithful. Then the natural map is an equivalence.
Proof.
By Proposition 1.75, the functor is fully faithful. Clearly, the functor between the quotients is an equivalence. Applying we obtain the commutation of with the pushout.∎
We observe that there is an alternative equivalent definition of as follows.
Proposition 4.13.
Let be an accessible localizing invariant. Then for a dualizable category we have
Proof.
Finally, we mention the following straightforward result about infinite semi-orthogonal decompositions.
Proposition 4.14.
Let be a dualizable category, a poset, and – an -indexed semi-orthogonal decomposition. Then the natural map is an isomorphism. Hence, the same holds for any finitary localizing invariant.
Proof.
As in the proof of Proposition 1.82, we have an ind-system of dualizable subcategories (with strongly continuous inclusions), where runs through finite subsets of and is generated by We have and has a finite semi-orthogonal decomposition with components We obtain the isomorphisms
Remark 4.15.
Note that the functor is the universal localizing invariant of dualizable categories which commutes with -filtered colimits. It follows formally that is symmetric monoidal: we only need to know that for a dualizable category the functor preserves short exact sequences and filtered colimits. An alternative way to argue is to use Theorem 4.16 below which says that is the right Kan extension of
4.3. Right Kan extension
In this subsection we explain another equivalent way to extend localizing invariants from small stable categories to dualizable categories. We prove the following result.
Theorem 4.16.
Let be an accessible localizing invariant, where is a presentable stable category. Then the functor is the right Kan extension of via
We will again use the fact that the category is -presentable (Theorem C.6).
It is convenient to define the following topologies on the opposite categories and In both cases we define a pretopology for which a family of functors is a cover if consists of one element and the functor is fully faithful. To deal with set-theoretic issues we also consider the induced pretopologies on the small subcategories and (note that these subcategories are closed under finite limits), where is an uncountable regular cardinal. In each case we denote by the corresponding Grothendieck topology.
We make the following observation.
Proposition 4.17.
Let be an accessible localizing invariant. Then and are sheaves for the -topology.
Proof.
It suffices to prove this for Let be a fully faithful strongly continuous functor between dualizable categories. Consider the Čech conerve of i.e. – the cofiber product of copies of It follows from Proposition 4.12 that is the Čech conerve of Since is stable, we conclude that ∎
Next, we observe the following equivalence of categories of sheaves.
Proposition 4.18.
Let be an uncountable regular cardinal, and let be a presentable category. Let be a -sheaf. Then the composition is also a -sheaf. The induced functor
is an equivalence. Its inverse is given by the right Kan extension.
Proof.
The first statement is obvious. The functor sends covers to covers, and it commutes with finite limits. It remains to apply Theorem C.6, part 3), condition (iv): it implies that for any fully faithful strongly continuous functor between -compact dualizable categories we can find a fully faithful strongly continuous functor where is -compact. In particular, the covering of is a refinement of the covering This implies the equivalence of categories of sheaves, as stated.∎
Proof of Theorem 4.16..
Choose an uncountable regular cardinal such that and commute with -filtered colimits. It suffices to show that the restriction of to is the right Kan extension of the restriction of to This follows immediately from Propositions 4.17 and 4.18: and are -sheaves respectively on and and is the composition ∎
4.4. Example: sheaves on the real line
We consider the special case of sheaves on the real line. Let be a dualizable category. Recall that the category is dualizable [Lur18]. In this subsection we give a detailed proof of the following special case of Theorem 0.2.
Proposition 4.19.
For a dualizable category we have
Hence, the same holds for any finitary localizing invariant, for example, for -theory.
Example 4.20.
Let be a field, and consider the dualizable category Combining Propositions 4.19 and 4.13, we see that
Given we can describe its representative in the fiber product as follows. Consider the object The endomorphism algebra of its image in the Calkin category is (the cohomology of the neighborhood of infinity of ). The object represents the class
Recall that if is compactly generated, then we have a notion of a singular support for -valued sheaves on -manifolds [RS18, Definition 4.5]. We will be interested in the following special case.
Denote by the cone of non-positive integers. The -topology on is defined as follows: a subset is -open if is open and In other words, is -open if either or or for some We denote by the set equipped with -topology. Denote by the identity map of considered as a map of topological spaces. Similarly, we denote by the map to equipped with -topology. The following is due to Kashiwara and Schapira.
Proposition 4.21.
[KS18] Let be a presentable stable compactly generated category. The pullback functor is fully faithful and its essential image is exactly the subcategory In particular, we have
Proof.
This is a straightforward generalization of the special case of [KS18, Theorem 2.5].∎
In fact there is a reasonable notion of singular support for sheaves with values in arbitrary presentable stable categories, see Remarks 4.24 and 4.25. To avoid the technicalities, we will use the following definition in the special case
Notation. For a presentable stable category we denote by resp. the category resp. of -valued sheaves on resp.
Note that for we have a tautological equivalence where the latter category was defined in Subsection 1.6. Here the poset is considered as a compactly assembled category. In particular, by Proposition 1.35 the category is dualizable.
Moreover, by Remark 1.36 for any presentable stable category we have
In particular, if is dualizable, then so is the category
Proposition 4.22.
Let be a dualizable category, and choose some dense subset Denote by the set considered as a linearly ordered set with the usual order.
1) We have a short exact sequence in
Here the functor is given by
2) We have
Proof of Proposition 4.22.
1) We first check that is indeed a strongly continuous localization. First, for define the functors
The right adjoint to is given by which is also the left adjoint to It follows that for the right adjoint to the functor sends to the functor given by
Here the latter description of follows from the fact that is a dense linearly ordered set. Now the right adjoint to is given by
Thus, and the maps are zero for In particular, is strongly continuous and the adjunction counit is an equivalence, so is a localization.
Now, we observe that the category is naturally equivalent to the right orthogonal This equivalence sends a sheaf to the functor Note that for we have hence
Using (4.6), we see that the category consists of functors such that for any the map is an isomorphism. This corresponds exactly to the fact that is a sheaf on
Finally, since the functor has a right adjoint, we have equivalences
This proves that we have a short exact sequence (4.4) in This proves part 1).
2) By (4.3), it is sufficient to prove the vanishing for Take and consider the short exact sequence (4.4). We need to show that the map is an equivalence in
Note that the categories and are compactly generated. We have an -indexed semi-orthogonal decomposition where the (fully faithful, strongly continuous) functors are defined in (4.5). Also, we have a tautological -indexed orthogonal decomposition of the category The functor is compatible with these decompositions and it induces the identity functors on the components. Hence, the map is an isomorphism. ∎
Proof of Proposition 4.19.
Again, it suffices to prove the result for We have a short exact sequence in
where the is the inclusion. Hence, it suffices to prove the result for sheaves on
We claim that there is a natural short exact sequence in
Assuming this, the result follows from Proposition 4.22.
The functor in (4.7) is the left adjoint to The functor is given by where is the map (extending ) to the sobrification of Note that as a set is naturally identified with (where is the generic point), and clearly
An explicit description of the functors and is the following. For a convex open subset and for we have
With these descriptions it is straightforward to check that (4.7) is indeed a short exact sequence. ∎
Remark 4.23.
Remark 4.24.
If is a -manifold and is a presentable compactly generated stable category, then the singular support of a -valued sheaf on is defined in [RS18, Definition 4.5], which is a straightforward generalization of [KS90, Definition 4.4.1]. However, if is a dualizable category which is not compactly generated, then already in the -dimensional case the non-characteristic deformation lemma [RS18, Theorem 4.1] fails and the conditions of [KS90, Proposition 5.1.1] are no longer equivalent.
To illustrate this, let Consider a -valued sheaf on given by for Then according to [RS18, Definition 4.5] the singular support would be the zero section But clearly the sheaf is not locally constant, hence the definition of the singular support has to be modified.
We mention several equivalent ways to define the singular support for a sheaf on a -manifold with values in a presentable stable category
1) A point is not contained in if and only if for some (any) choice of local coordinates it satisfies the equivalent conditions and of [KS90, Proposition 5.1.1].
2) The singular support is the smallest conical closed subset such that is contained in the essential image of the functor
This functor is fully faithful since the inclusion has a left adjoint.
3) The singular support of is the support of the conical sheaf (of spectra) on (microlocal Hom), where the bifunctor is defined as in [KS90, Definition 4.4.1].
4) Choose a regular cardinal such that takes values in Then the composition defines a sheaf with values in the compactly generated category (since the Yoneda embedding commutes with limits). We put where the latter singular support is defined as in [RS18, Definition 4.5].
5) If is dualizable, we can define the singular support of to be the singular support of the sheafification of the presheaf where the latter is defined as in [RS18, Definition 4.5].
Remark 4.25.
Essentially the above issue with the definition of the singular support reduces to the fact that Kashiwara’s criterion for a functor to be constant fails when is not compactly generated (even if we require that is dualizable). Namely, if is compactly generated and is a functor, then is constant if and only if we have
see [RS18, Corollary 3.2]. However, if and is given by then the condition (4.8) holds but is not constant.
4.5. Commutation of K-theory with infinite products
Recall that by [KW19] -theory commutes with infinite products of small stable Karoubi-complete categories. Their proof is based on the Grayson’s construction [Gra12] (more precisely, its version for stable -categories) which roughly speaking allows to identify of some category with a direct summand of of another category. We will give an alternative (and in our opinion more conceptual) proof of the commutation of -theory with products, using the category of sheaves on the real line instead of the Grayson’s construction.
Remark 4.26.
We first recall a simple proof of the commutation of with infinite products.
Proposition 4.27 (Heller’s criterion).
[KW19, Lemma 2.12] Let be a small triangulated category, and take two objects The following are equivalent:
we have in
there exist objects and exact triangles
Corollary 4.28.
[KW19, Lemma 5.1] For a collection of small triangulated categories we have
We will prove the following general result.
Theorem 4.29.
1) Let be an accessible stable category with a non-degenerate -structure. Let be accessible localizing invariants and a morphism such that is an isomorphism. Then is an isomorphism.
2) In particular, both -theory and continuous -theory commute with infinite products.
Proof that 1)2)..
By Proposition 1.103 it suffices to prove the statement for continuous -theory. Consider a collection of dualizable categories. We need to show that We may assume that the categories are all equivalent: indeed, each is a retract in of the category
In other words, we need to show that for a set and a dualizable category the map is an isomorphism. But this is the special case of 1) for the standard -structure on Indeed, by Proposition 1.105 the functor is a localizing invariant, and so is the functor When is compactly generated, the map is an isomorphism by Corollary 4.28.∎
To prove part 1), we need the following result on -equivalences.
Proposition 4.30.
Let be a small stable idempotent-complete category with an -indexed semi-orthogonal decomposition Denote by the full stable subcategory generated by and denote by the right adjoint to the inclusion functor. Consider the sequence as an inverse sequence where the transition functors are left adjoint to the inclusions. Then the compositions define a -equivalence
Proof.
Put The functors are compatible with the functors hence we obtain a functor By construction, we have We need to show that in
For each the inclusions define a fully faithful functor Its left adjoint is simply the structural functor Denote by the composition functors, Since the compositions vanish for we see that the direct sum of the functors is well-defined. Moreover, we have
Now, denote by the composition, and put (where ). Equivalently, is the semi-orthogonal projection onto the left orthogonal to (where ). Denote by the natural transformations, Then
Since the composition vanishes for we see that the direct sum is also a well-defined endofunctor of By the above, we obtain an exact triangle
We conclude that
in as required. ∎
We deduce the following statement about commutation of with certain “extremely nice” sequential limits.
Corollary 4.31.
Let be an inverse sequence in such that each functor has a fully faithful right adjoint Then we have an isomorphism of abelian groups
Proof.
We deduce that all accessible localizing invariants vanish for the categories of the form
Proposition 4.32.
Let be a dualizable category. Then for any regular cardinal we have
Proof.
Since we may assume that Applying 4.22 to we obtain a short exact sequence
in Moreover, the source and the target of are compactly generated categories. Denote by the induced functor on the subcategories of compact objects. It suffices to show that is a -equivalence. As in the proof of proposition 4.22, we put and
Note that the functor has an obvious section which corresponds to the collection of “representable” functors where is defined in (4.5). We need to show that in
Choose any bijection Denote by the full stable subcategory generated by the objects Then is equivalent to the category where as usual denotes the poset Clearly, each inclusion has both left and right adjoints. Moreover, we see that the elements form a basis of
It follows from Corollary 4.31 that Since we deduce that
It remains to observe that the functor induces the identity map on it sends each object to itself, ∎
Proof of Theorem 4.29.
. We first prove that is an isomorphism for The argument is basically the same as in [KW19]. We know that is an isomorphism by assumption. Assuming that is an isomorphism for some we see that for any the composition
is also an isomorphism. By induction, we see that is an isomorphism for
Denote by the induced map between the localizing invariants of dualizable categories. Arguing as above, we see that is an isomorphism for It remains to show that if is an isomorphism for some then is also an isomorphism.
Let be a dualizable category. Then the functor is a strongly continuous localization (the right adjoint sends an object of to the corresponding constant sheaf). We denote by the kernel of Applying Proposition 4.32, we see that the composition
is an isomorphism. Therefore, the map is an isomorphism. ∎
5. Sheaves and cosheaves on continuous posets
5.1. Continuous posets
Recall that a partially ordered set is called continuous if has directed supremums (joins), and the functor has a left adjoint Equivalently, this means the following.
Assuming that has directed supremums, one defines another relation on namely, if for any directed system of elements of such that we have for some Then is continuous if and only if for each the poset is directed and we have equality
In this case we have We refer to [GHKLMS80, BH81] for details.
Example 5.1.
Let with the usual linear order. Then is continuous and we have iff
Example 5.2.
Let be a locally compact Hausdorff space, and let – the set of open subsets of ordered by inclusion. Then is continuous and we have iff i.e. and is compact.
From now on we assume that is a continuous poset. We define a Grothendieck topology on (considered as a category) as follows. A sieve on is a covering sieve if and only if it contains all This is indeed a topology, since for any there exists such that Below we will consider sheaves and cosheaves on with respect to this topology.
5.2. Sheaves of spectra on continuous posets
Let be a continuous poset. We denote by the category of sheaves on (where is considered as a category) for the above topology. This is in fact the category from Subsection 1.6.
Namely, since every object of has the smallest covering sieve, a presheaf is a sheaf if and only if
Denote by the sheafification functor. Explicitly, we have
For denote by the “representable” presheaf of spectra,
Its sheafification is given by
Proposition 5.3.
The category is dualizable. For any dualizable category we have an equivalence
Proof.
This is a special case of Proposition 1.35. ∎
5.3. Localizing invariants of categories of sheaves on continuous posets
Consider now a presheaf on with values in for We denote by the restriction functors for and by their right adjoints. Consider the category of -valued presheaves on (i.e. the category of sections of the associated biCartesian fibration over ). For we denote by the functor given by
The right adjoint to is given by
Denote by the category of -valued sheaves on For we have a functor The right adjoint to is given by
Our goal in this section is to prove the following result.
Theorem 5.4.
1) The category is dualizable.
2) For each compact element the functor is strongly continuous and fully faithful.
3) The functors induce an isomorphism
Proof.
For a presheaf we denote by its sheafification. Since any object has the smallest covering sieve we see that
In particular, we have for This shows that for the functor commutes with coproducts. This proves 2).
To prove 1), we first observe that the category is dualizable. Indeed, we have a -indexed semi-othogonal decomposition of the category in
Since are dualizable, so is the category by Proposition 1.82.
Now, it follows from (5.2) that the sheafification functor has a left adjoint We deduce that the category is dualizable.
To prove 3), consider the full subcategory By the above we have a short exact sequence in
where the functor is the left adjoint to the inclusion. We first observe the following.
Claim 1. Let be a presheaf such that for any the map is zero. is Then is contained in
Proof of Claim 1..
This follows directly from the description of the sheafification functor (5.2). ∎
Next, we observe that the category has the following semi-orthogonal decomposition.
Claim 2. For the functor is zero (resp. fully faithful) for (resp. for ). Moreover, we have a semi-orthogonal decomposition
Proof of Claim 2..
Let us describe the functor more explicitly. For we have
Indeed, it follows from (5.2) (by adjunction) that for the object is left orthogonal to On the other hand, it follows from Claim 1 that the object is contained in (since the latter presheaf can take non-zero values only for such that but ). This proves the isomorphism (5.5).
We deduce that
In particular, we see that is zero for If then
Hence, the adjunction unit is an isomorphism. We conclude that the functor is fully faithful for
Similarly, if and then for This implies the semi-orthogonality.
Finally, taking into account the vanishing of for we deduce that the categories generate This proves the claim.∎
5.4. Localizing invariants of categories of cosheaves on continuous posets
Let now be a copresheaf of dualizable categories on for By a -valued copresheaf on we mean a section of the associated biCartesian fibration over We denote by this category of copresheaves. Further, we denote by the full subcategory of cosheaves: it consists of copresheaves such that We denote by the cosheafification functor, i.e. the right adjoint to the inclusion.
Again, we denote by the left Kan extension functor, given by
The right adjoint to is given by We have a composition The following result is basically dual to Theorem 5.4.
Theorem 5.5.
1) The category is dualizable.
2) For each compact element the functor takes values in cosheaves, hence we have a fully faithful strongly continuous functor
3) The functors induce an isomorphism
Proof.
As in the proof of Theorem 5.4, the category is dualizable and we have a -indexed semi-orthogonal decomposition The cosheafification functor is continuous, since we have
It follows that the category is dualizable. This proves 1).
Part 2) is clear from the definitions.
We prove 3). Consider the short exact sequence
where is the kernel of the cosheafification, and is the left adjoint to the inclusion. We claim that the functors are fully faithful for and we have a semi-orthogonal decomposition
Indeed, this follows directly from (5.6): we have
It follows by adjunctions that the functors are fully faithful () and their images are semi-orthogonal as required. Since the functors take values in cosheaves, we obtain the semi-orthogonal decomposition (5.8).
6. Sheaves on locally compact Hausdorff spaces
6.1. Sheaves on finite CW complexes
The following result describes arbitrary (accessible) localizing invariants of categories of sheaves on finite CW complexes, in the case of constant coefficients. Here we consider finite CW complexes both as topological spaces and as -groupoids (i.e. objects of -category of finite spaces).
Theorem 6.1.
Let be a finite CW complex, considered as a topological space. Let be a dualizable category. Then for any accessible localizing invariant we have an isomorphism
Here in the right hand side we consider as an object of
Proof.
Suppose that has -filtered colimits and commutes with -filtered colimits. By Proposition 4.32, we have Arguing as in the proof of Proposition 4.19 we obtain
From the short exact sequence
in we obtain
It follows that for any finite CW complex and for any dualizable category we have
Therefore we have a well defined functor such that for a finite CW complex we have Since it suffices to prove that commutes with pullbacks.
This is straightforward. Namely, consider cellular closed embeddings of finite CW complexes Then both pullback functors and are strongly continuous localizations and we have a pullback square
By Proposition 4.11 we obtain an isomorphism
Hence, commutes with pullbacks. This proves the theorem.∎
6.2. Sheaves and K-sheaves on locally compact Hausdorff spaces
Let now be a locally compact Hausdorff space, and consider a presheaf on with values in In particular, we require that the restriction functors are strongly continuous.
We will consider -valued (pre)sheaves on Recall that is a sheaf if and only if it satisfies the following conditions:
for any open we have a pullback square
for any open the map is an isomorphism.
Recall from Example 5.2 that the poset of open subsets of is continuous. In particular, we have a topology on as in Subsection 5.1. This topology is weaker than the usual one, and we denote by the category of sheaves on with respect to this topology. We call the objects of “continuous presheaves”. These are exactly the presheaves satisfying the condition (iii). We denote by the left adjoint to the inclusion. As in Subsection 5.3 we have
We denote by the sheafification functor. It has a factorization:
We make the following observation.
Proposition 6.2.
Suppose that is a presheaf which satisfies descent for finite covers, i.e. the conditions (i) and (ii). Then is a sheaf. Equivalently, the natural map is an isomorphism.
In particular, the inclusion functor is continuous.
Proof.
Assuming that satisfies finite descent, it suffices to show that satisfies finite descent.
Clearly, so (i) holds. To see that (ii) holds for we first introduce the following notation: for an open we denote by the poset Now take some open subsets and consider the maps
The standard arguments show that both maps (6.1) and (6.2) are cofinal (hence, the maps between the same posets with the opposite order are final). This directly implies that (ii) holds for Thus, is indeed a sheaf.
For the final assertion of the proposition, it suffices to notice that the class of presheaves satisfying finite descent is closed under colimits. ∎
We consider the category as the first approximation of the category Note that if is a presheaf with values in then we know the finitary localizing invariants of the category by Theorem 5.4. To construct the further approximations, it is convenient to use the language of -sheaves, which we know recall.
We first extend the presheaf to locally closed subsets by left Kan extension: for a locally closed we put
where runs over open subsets such that is a closed subset of In particular, we obtain a presheaf of presentable stable categories on
Denote by the poset of compact subsets of
Definition 6.3.
1) Denote by the category of -presheaves, i.e. the category of -valued presheaves on
2) We denote by the full subcategory of continuous -presheaves, i.e. -presheaves such that for any compact we have an isomorphism
Here if contains an open neighborhood of
3) We denote by the full subcategory formed by continuous -presheaves satisfying additional conditions:
i)
ii) For compact subsets the following is a pullback square:
Note that the poset is continuous, and iff Hence, the category is the category of -valued cosheaves for the topology from Subsection 5.1.
The following is an (almost) straightforward generalization of Lurie’s theorem for the case of a constant presheaf
Proposition 6.4.
[Lur09, Theorem 7.3.4.9, Corollary 7.3.4.10] 1) Consider the following functors
They are quasi-inverse equivalences.
Proof.
Part 1) is proved by the same argument as [Lur09, Theorem 7.3.4.9], and in fact we do not need the strong continuity of the restriction functors
We prove 2). Let be a -valued sheaf, and denote by the associated continuous -presheaf. Clearly, Consider compact subsets and put It suffices to show that the square
is the directed colimit of pullback squares
over open neighborhoods
Clearly, we have
Applying Proposition 1.67 and Lemma 6.5 below, we obtain the isomorphisms
(here we used that the restriction functors are strongly continuous). The proof for the remaining two objects of the square (6.5) is similar.
The proof in the other direction is as in [Lur09, Theorem 7.3.4.9] (and here we do not need the strong continuity of the restriction functors). ∎
Lemma 6.5.
Let be a directed poset, and consider a functor Consider a section of the associated biCartesian fibration over Denote by the transition functors, and denote by the functors to the colimit. Then for any we have
Proof.
Fix Applying Proposition 1.67, we obtain the isomorphisms
We introduce some notation. Suppose that is compact. Given a finite collection of closed subsets of such that we put
We consider as a poset with the reverse inclusion order.
Consider the set of all collections of compact (=closed) subsets of of the form as in (6.7). Define the partial order on by setting if Then is a directed poset.
Remark 6.6.
1) If is locally compact, then the poset has a natural Grothendieck pretopology, where the family is a cover if is finite and Then for each compact the poset is a cofinal family of covering sieves for
2) If is compact and is a sieve as above, then we have a final functor
Proposition 6.7.
Let and be as above. Consider the functor which under the equivalence of Proposition 6.4 corresponds to the sheafification functor. Then for we have
Proof.
Consider the full subcategory of presheaves satisfying only the finite descent, i.e. sheaves for the finitary topology on described in Remark 6.6 1). Then the corresponding finitary sheafification functor is given by (6.8). Thus, we need to show the following: for the -presheaf is continuous. This is straightforward. ∎
We observe the following.
Proposition 6.8.
For a closed embedding the functor is a strongly continuous localization.
Proof.
The functor and its right adjoint are given by the formulas
Hence, is continuous and the adjunction counit is an isomorphism.∎
For we define the presentable stable category
Proposition 6.9.
1) For each the functor is a strongly continuous localization. The essential image of the right adjoint consists of continuous presheaves such that for each compact we have
2) Put Then the functor is a localization, and the essential image of its right adjoint is exactly the category of -sheaves In particular, we have an equivalence
Proof.
Observe that the analogue of part 1) for categories holds trivially. To prove the statement 1), we only need to show the following.
Claim. Let be a -presheaf satisfying (6.9). If each of the pullbacks is in then also
Proof of Claim..
We may assume that i.e. Take a compact subset It suffices to show that the map
is an isomorphism, and similarly for and
2) It follows from 1) that the functor is a strongly continuous localization, and the essential image of its right adjoint consists of continuous -presheaves such that for any compact and for any we have
We need to show that this condition holds if and only if is a -sheaf.
The “if” direction is obvious. We show the “only if” direction. Take some pair of compact subsets For any open neighborhoods, denote by the sieve Applying the condition (6.10) to and we see that the following is a pullback square:
Taking the colimit over and (and using Lemma 6.5 again), we obtain the following square:
Hence, this is a pullback square and is a -sheaf. ∎
6.3. Localizing invariants of categories of sheaves
From now on we assume that our presheaf of categories takes values in dualizable categories. Our goal in this section is to prove the following theorem.
Theorem 6.10.
Let be a locally compact Hausdorff space and is a presheaf on with values in Then the category is dualizable and we have the following natural isomorphism in
Hence, this also holds for any finitary localizing invariant, for example, for -theory.
Note that we already know the finitary localizing invariants of the categories of continuous (-)presheaves.
Proposition 6.11.
The category is dualizable and we have
Proof.
This is a special case of Theorem 5.5 since the compact elements of the continuous poset are exactly the open-compact subsets of ∎
We need the following observation about the categories from Proposition 6.9.
Proposition 6.12.
Suppose that is compact. For a sieve the natural functor
is an equivalence. Moreover, we have
Proof.
This is basically an iterated application of Propositions 4.11 and 6.9. We prove the statement by induction on where the case is clear.
Let where We put
Using the induction hypothesis, we see that the functors
are equivalences.
Proposition 6.13.
Let be locally compact Hausdorff, and a presentable stable category. Given a -presheaf we define another -presheaf by the formula
where runs over open-closed subsets of We have a natural map
1) If is continuous, then so is
2) Assuming that is continuous, the map of -sheaves is an isomorphism.
Proof.
1) This follows immediately from the following observation: for any compact the map
is a bijection of sets.
2) Recall that is the sheafification of for the finitary topology on Consider the weaker pretopology on for which a family is a covering if is finite and Denote by the corresponding sheafification functor. It suffices to show that the map is an isomorphism.
Let us observe that the weak sheafification functor is given by
where the colimit is over the decompositions of into a finite disjoint union of compact subsets. Now it is clear that the map is an isomorphism. ∎
Proof of Theorem 6.10..
We first show directly that is dualizable (this is certainly known, but we include a proof for completeness). Given an open subset we denote by the left adjoint to the functor Equivalently, for the sheaf is the sheafification of the presheaf
We observe the following: for any compact inclusion of open subsets, for any objects and for any compact map in the induced map of sheaves is compact in Indeed, this follows from the factorization
and the fact that the functor is strongly continuous.
Now observe that we have a generating collection of sheaves
For any such object since is a dualizable category, we can choose a sequence in such that and all the maps are compact. Choose a sequence of open subsets and compact inclusions such that and each is countable at Then
and by the above discussion each map is compact. This shows that the category is dualizable.
To compute we first reduce to the case when is compact. Denote by the one-point compactification of and denote by the embedding. Let be the presheaf of dualizable categories on obtained by extension by zero of Then hence we may replace the pair with the pair Thus, we may and will assume that is compact.
6.4. Compact objects
Let be as above. The following result describes the category of compact objects of the category It is a generalization of [Nee01b, Theorem 0.1], and the proof is based on the ideas of loc. cit.
Proposition 6.14.
Let be a sheaf. The following are equivalent.
is a compact object in the category
has compact support and can be covered by open subsets such that for some compact object
Proof.
The implication is clear.
. Suppose that is compact.
We first show the following: if for some then there is an open neighborhood such that Indeed, since and is compact, we deduce that is a direct summand of some sheaf of the form where Hence the restriction of to vanishes.
Now let be a point such that Since the functor is strongly continuous, the object is compact in Hence, there is an open neighborhood and a compact object such that Shrinking if necessary, we can lift the identity map of to a map in the category Denote by the associated map of sheaves. Choosing an open such that we see that is a compact object of whose stalk at vanishes. By the above, we deduce that vanishes in a neighborhood of Hence, shrinking further, we may assume that as required.
Finally, to see that has a compact support, note that Since is compact, it is a direct summand of some sheaf which has compact support. ∎
Corollary 6.15.
Let be a non-empty locally compact Hausdorff space, and let be a non-zero presentable stable category. The following are equivalent.
The category is compactly generated.
The category is compactly generated, and is locally profinite.
Proof.
Consider the collection of sheaves
Then each sheaf in is compact and the right orthogonal to consists of sheaves such that for every open-compact Since is locally profinite, this means that Hence, generates the category
Choose a point The pullback to is a strongly continuous quotient functor Since the category is compactly generated, so is
Next, we reduce to the case when is compact. If not, consider the short exact sequence
Since the categories on the left and on the right are compactly generated, so is the category in the middle. Hence, we may and will assume that is compact.
Suppose that is not profinite. Choose a connected component such that is not a single point. Then is closed, hence compact Hausdorff. The functor is a strongly continuous localization. Since the category is compactly generated, so is the category Replacing with we may and will assume that is connected.
Choose a point Choose a non-zero compact object and consider the skyscraper sheaf on Then we can find a directed system of compact -valued sheaves such hat By Proposition 6.14 each is a locally constant sheaf with compact stalks.
Since is compact in we can find and a splitting Using compactness of again, we can find an open neighborhood and a map which lifts Choose a compact such that is a neighborhood of Let be the connected component of in Then hence we may and will assume that
So we have the map Take some Since the sheaf is locally constant, the set
is open-closed in We have Since is connected, we have Choosing a point we see that the map is non-zero. But a contradiction. ∎
Appendix A Monomorphisms and epimorphisms of presentable and dualizable categories
Here we show that the monomorphisms and epimorphism in and are as expected: these are respectively the fully faithful functors and the quotient functors.
Proposition A.1.
Let be a continuous functor between presentable stable categories.
1) is a monomorphism in if and only if is fully faithful.
2) is an epimorphism in if and only if is a quotient functor.
Proof.
1) The “if” part is clear. For the “only if” part, we see that the map of groupoids
is a monomorphism, hence is fully faithful (since and are stable).
2) Again, the “if” part is clear. For the “only if” part, note that where both functors are Hence, the functor is an equivalence, in particular, it is fully faithful. Hence, for any the square
is biCartesian, thus is fully faithful and is a quotient functor.∎
Proposition A.2.
Let be a strongly continuous functor between presentable stable categories.
1) is a monomorphism in if and only if is fully faithful.
2) is an epimorphism in if and only if is a quotient functor.
Proof.
1) The “if” part is clear. We prove the “only if” part.
Suppose that is a monomorphism. Then for a presentable stable category and strongly continuous functors we have By adjunction, we have
Put and consider the semi-orthogonal gluing Since is continuous, the inclusions of the two copies of are strongly continuous. Denote by the semi-orthogonal projections (respectively, the left and the right adjoints to the inclusions). Then and are strongly continuous. Applying (A.1) to we obtain
hence which exactly means that is fully faithful. ‘
2) follows from Proposition A.1 since the functor is conservative and commutes with colimits. ∎
Proposition A.3.
Let be a strongly continuous between dualizable categories.
1) is a monomorphism in if and only if is fully faithful.
2) is an epimorphism in if and only if is a quotient functor.
Proof.
1) The “if” direction is clear. We prove the “only if” direction.
Suppose that a functor is a monomorphism in By Proposition 1.88 the functor is fully faithful. Hence, the functor is fully faithful. Arguing as in the proof of Proposition A.1, we conclude that is fully faithful.
2) follows from Proposition A.1 since the functor is conservative and commutes with colimits. ∎
Remark A.4.
In fact the description of monomorphisms in can be proved similarly to the case of dualizable categories. However, we did not discuss the pullbacks in hence we gave a different proof.
Appendix B Essential image of a homological epimorphism
Consider a homological epimorphism where and are small stable idempotent-complete categories. We know that if the category is compactly generated, then In particular, the essential image of is a stable subcategory of In this case by Thomason’s theorem every object of of the form can be lifted to
It turns out that if is not compactly generated then the image of might not be a stable subcategory of More precisely, the following result shows that in general for a dualizable category there may exist an object such that the object cannot be lifted to
Proposition B.1.
Let be a field and consider the dualizable category Then there exists a non-zero idempotent-complete stable subcategory such that and the intersection of with the image of is zero. In particular, the image of the homological epimorphism is not a stable subcategory.
Proof.
Consider the constant sheaf We have
– the cohomology of the neighborhood of infinity of The first resp. second copy of corresponds to the neighborhood of resp. We denote by the corresponding idempotents in the Calkin category, and denote their images by We obtain fully faithful functors which send to resp. We claim that any non-zero object of the image of is not contained in the image of
Assume the contrary: suppose that for some sheaf the image of in the Calkin category is isomorphic to a non-zero object of Take some real number Since the constant sheaf is a compact object of it follows that the image of in is zero. Hence, the pullback of to is a compact object, i.e. for some perfect complex This holds for any hence for some Then and the image of in is isomorphic to the direct sum which is not contained in the image of This gives a contradiction.∎
Appendix C Presentability of the category of dualizable categories
In this section we give an alternative, “explicit” proof of Ramzi’s theorem [Ram] which states that the category is -presentable. We also describe explicitly the -compact objects in for any uncountable regular We start with the following observation.
Proposition C.1.
Let be a dualizable stable category. Consider the full subcategory which consists of pairs such that the image of generates by colimits. Then the functor is an equivalence of categories.
The inverse functor is given by
Proof.
Indeed, for any the functor is conservative and commutes with all colimits, in particular, with geometric realizations. Hence, the functor is monadic by Lurie-Barr-Beck theorem, and we have an equivalence
It remains to check that for any the category is dualizable and the functor is strongly continuous. The latter assertion is obvious, and the former one follows from Proposition 1.56.∎
We will repeatedly use the following.
Proposition C.2.
Let be a presentable stable -monoidal category. Let be a regular cardinal such that is -compactly generated, the tensor product of -compact objects is -compact, and the unit object is -compact (equivalently, is an -algebra in ).
1) The category of -algebras in is -compactly generated.
2) If is uncountable, then an -algebra is -compact in if and only if the underlying object of is -compact in
Proof.
1) For an object denote by the tensor algebra of i.e. the free -algebra generated by Then is the left adjoint to the forgetful functor. If then our assumptions imply that Since the forgetful functor commutes with -filtered colimits, the objects are -compact in Since the forgetful functor is conservative and is -compactly generated, we deduce that the category is -compactly generated. More precisely, the category is the smallest full subcategory of which contains the objects and is closed under -small colimits.
2) The forgetful functor is monadic, and the corresponding monad on is given by the functor Since preserves -compact objects, the assertion follows from [BSY22, Lemma 6.17].∎
We will apply Proposition C.2 to monoidal categories of the form where is a dualizable category. Since this category of functors is also dualizable, it is -compactly generated.
Proposition C.3.
Let and be as in Proposition C.2, and suppose that is uncountable. Let be a left -module in i.e. is -presentable and the action is -strongly continuous. Let be a -compact -algebra in and an -module. The following are equivalent.
is -compact in
the underlying object of is -compact in
Proof.
Indeed, the forgetful functor is continuous and monadic, and the monad preserves -compact objects. Hence, the equivalence follows from [BSY22, Lemma 6.17].∎
Recall the category of sheaves of spectra on with singular support in For we denote by the sheaf As usual, we denote by the constant sheaf with value
Proposition C.4.
Let be a dualizable stable category. Given a strongly continuous functor we define a functor The assignment defines an equivalence between and the category of functors such that
for any we have
for any the morphism is compact.
Proof.
Since is dense in the category of functors with required properties is equivalent to the category of functors with analogous properties. The assertion now follows as a special case of Proposition 5.3 applied to ∎
Corollary C.5.
For any dualizable category and for any countably presented object there exists a strongly continuous functor such that
We prove the following result on the presentability of the category and on the -compact objects of this category.
Theorem C.6.
Put
1) The category of dualizable categories is -presentable.
2) Let be an uncountable cardinal and a dualizable category. The following are equivalent.
is a -compact object of
there exists a -small set and a -compact monad such that
Moreover, if then can be assumed to be a one-element set.
The functors and are -strongly continuous, i.e. their right adjoints commute with -filtered colimits. Equivalently, is dualizable in
There exists a short exact sequence in where and are -compact in
is -compact in
We need a few auxiliary statements. Let be an uncountable regular cardinal.
Lemma C.7.
Let and be dualizable categories, and suppose that the functors and are -strongly continuous. The following are equivalent for a continuous functor
the object is -compact in
the functor is -strongly continuous.
Proof.
This is essentially a standard statement, we include the proof for completeness.
. Since the functors and are -strongly continuous, so is the composition
This composition is exactly the functor
. Since the functors and are -strongly continuous, so is the composition
This composition sends to ∎
Lemma C.8.
Consider a short exact sequence in
1) If is -strongly continuous, then so is
2) If is -strongly continuous, then so is
3) If and are -strongly continuous, then so is
4) If and are -strongly continuous, then so is
Proof.
1) The functor is the composition of -strongly continuous functors
hence it is -strongly continuous.
2) Similarly, the functor is the composition of -strongly continuous functors
hence it is -strongly continuous.
3) It suffices to prove that is in We have an exact triangle
Since is -compact, it suffices to show that is -compact. The latter object corresponds under the self-duality of to the composition
The latter functor is -strongly continuous by our assumptions. Applying Lemma C.7 (to the categories and ) we conclude that is -compact.
4) This is proved by reversing the proof of 3).∎
Lemma C.9.
Let be a strongly continuous functor between dualizable categories, and denote by its right adjoint. Suppose that the image of generates i.e. is conservative. Suppose that the functors and are -strongly continuous. Then the functors and are also -strongly continuous.
Proof.
We first show that is -strongly continuous. Since it suffices to show that the composition
is -strongly continuous. But it is identified with another composition
The latter composition is -strongly continuous since both functors are -strongly continuous.
Next, we show that the functor is -strongly continuous, i.e. the object is -compact. Denote by the continuous monad on associated with Then is a monad on and we have Since is a -compact monad, the -compactness of is equivalent to the -compactness of in The latter object is simply under the identification Buy our assumptions, is -compact. ∎
Proof of Theorem C.6..
We first show the following.
Claim. For any dualizable category there exists a set and a monad such that
Proof of Claim.
Next, we prove some implications from part 3).
. Recall that is a -compact object in Hence, any dualizable object in is also -compact. This proves the implication.
. Let be a -directed system in and We only need to prove the following: if is some -strongly continuous functor such that the composition is strongly continuous, then there exists some such that the composition is -strongly continuous. To see this, note that the composition
is zero. By Proposition 1.74, we have We deduce that for some the composition
is zero. This exactly means that the composition is strongly continuous.
. Note that it is sufficient to prove that if is -compact in then the functors and for are -strongly continuous. Indeed, assuming this, we can apply Lemma C.8.
So let be in It follows directly from Proposition 1.6 that the functor is -strongly continuous: this exactly means that the spectra are -compact. To show that (the diagonal -bimodule) is -compact, choose a -small ind-system of compact (finitely presented) objects of such that By a straightforward generalization of [TV07, Proposition 2.14], for each the object is compact in Hence, the object is -compact: it is a -small filtered colimit of compact objects, namely
This proves the implication.
. First, we have a short exact sequence
Hence, satisfies and by the above it also satisfies If is a -small set, then also satisfies If is -compact, then by Lemma C.9 we see that the category satisfies
We now prove 1). By the above, it suffices to prove that the category is generated by colimits by the categories of the form where is an -compact monad. Let be an arbitrary dualizable category. Choose a generating collection of -compact objects of such that for each the object is a formal colimit of a sequence of objects of Then is a directed union of countable subsets with the same property. Denoting by the full subcategory generated by by colimits, we see that is dualizable, and the inclusion is strongly continuous. By Corollary 1.70, we have Hence, we may assume that is countable. Then is generated by a single -compact object. Arguing as in 1), we obtain an equivalence for some monad By Proposition C.2, we can find an -directed system of -compact monads such that Then This proves 1).
. Suppose that is -compact in We argue as in the proof of Claim. We see that is generated by a -small collection of -compact objects We obtain a monad such that Since is -compact in we deduce that is -compact in the category of dualizable categories under By Proposition C.1, this implies that is -compact.
We already proved that the conditions (i), (ii), (iii) and (v) are equivalent, and they are implied by (iv). It remains to prove the implication . So let be -compact in By we have a strongly continuous functor such that the image generates. Put Define the category by the pushout square in
Since is compactly generated, so is Since the categories and are -compact in so is It follows that the category is -compact in
Since the functor is fully faithful, so is Consider the quotient Then is compactly generated and is -compact in Therefore, is -compact in Hence, the short exact sequence
has the required properties.∎
Appendix D Urysohn’s lemma for dualizable categories
In this section we prove the following result.
Theorem D.1.
The category is generated by colimits by a single -compact object, namely by the category
Note that the classical Urysohn’s lemma for compact Hausdorff spaces essentially says that the category (the opposite category of compact Hausdorff spaces) is generated by colimits by the unit interval The latter is easily seen to be an -compact object.
We will again use the equivalence
given by Proposition 5.3. Slightly abusing the terminology, we will call a functor strongly continuous if its left Kan extension to is strongly continuous. We denote the category of such functors by
The following proposition essentially implies Theorem D.1.
Proposition D.2.
1) Let be a dualizable category. The functor is a Verdier quotient. Its kernel is the full subcategory of ind-zero functors, i.e. of functors such that for any there exists such that the map is zero.
2) Denote by the right adjoint to Then for a strongly continuous functor we have
Here runs through strictly monotone functions such that and for
The notation simply means that we look only at In other words, the last condition on means that the endofunctor commutes with existing colimits.
We will need the following basic fact about the sequential limits of directed colimits.
Lemma D.3.
Let be a presentable -category such that filtered colimits in commute with finite limits (i.e. strong (AB5) holds in ) and (AB6) for countable products holds in (for example, can be any compactly assembled presentable category, such as or ). Let be a directed poset, and consider as a poset with the usual order. Let be a functor. Then we have a natural isomorphism
where runs through order-preserving maps. Here the set is considered as a subposet of
Proof.
Note that the inclusion
is final. It follows that for any functor we have
where the two maps have components resp.
Note also that the poset of order-preserving maps is cofinal in the poset of all maps Using our assumptions on we obtain the isomorphisms
Proof of Proposition D.2.
Since we already know that the functor is essentially surjective, it suffices to prove part 2). Indeed, (D.1) implies that i.e. is a homological epimorphism. If is a function with required properties such that for all then the cofinality of implies that is an ind-zero functor Hence, the kernel of is generated by colimits by ind-zero functors, as required.
We now prove 2). Denote by the poset of functions with required properties. Then is a cofinal subposet of the poset of all monotone functions such that Since is directed, so is Consider another subposet which consists of monotone functions such that takes values in and is constant on and on each semi-open interval for Then is also cofinal in
Proof of Theorem D.1.
It suffices to prove the following: if a strongly continuous functor between dualizable categories induces an equivalence then is an equivalence.
It is clear that is essentially surjective. The fully faithfulness follows directly from Proposition D.2. Namely, for strongly continuous we have
Remark D.4.
Suppose that we know the cocompleteness of the category and the -compactness of in (both are not difficult to check). It is easy to deduce from the proof of Proposition D.2 that is generated by colimits by
The same applies to the category of compactly assembled presentable categories and strongly continuous right exact functors if we replace the poset with the poset with the usual order (the former poset is not a presentable category since it does not have an initial object).
Appendix E Phantom maps, pure injectivity and Adams representability
Let be a dualizable category, and let be its homotopy category. As in Subsection 1.8, we denote by the ideal of compact maps in Recall that
Definition E.1.
1) We say that a right -module is almost zero if it is annihilated by In other words, for any compact morphism in the map is zero.
2) We denote by the quotient of by the full (Serre) subcategory of almost zero modules.
The subcategory of almost zero modules is indeed a Serre subcategory since We make the following observation.
Proposition E.2.
The quotient functor has an exact left adjoint given by
Proof.
We observe the following interpretation of local coherence of abelian categories.
Proposition E.3.
Let be a Grothendieck abelian category which is compactly generated. The following are equivalent.
is locally coherent.
The functor is exact.
Proof.
. If is locally coherent, then the category is abelian and the inclusion functor is exact. Passing to ind-completions, we get the functor which is therefore exact.
. Let be a morphism in We need to show that the kernel of in is also a compact object. Note that the maps and are isomorphisms. Since both functors and are left exact, we conclude that the map is also an isomorphism. This exactly means that as required.∎
We deduce the following basic properties of the abelian category
Proposition E.4.
1) If is compactly generated, then we have
2) In general, the category is a Grothendieck abelian category which satisfies (AB4*) and (AB6). In particular, it is compactly assembled. The functor is exact.
Proof.
1) If is compactly generated, then is an ideal of maps in which factor through an object of Hence, the kernel of the (exact) restriction functor is the category of almost zero modules. Since its left adjoint is fully faithful, the assertion follows.
2) Recall that the category of almost zero modules is closed under limits and colimits. Since the axioms (AB4*) and (AB6) hold in they also hold in this is a special case of Roos’ theorem [Roo65]. It is clear that is a Grothendieck category. Finally, the functor is the composition of exact functors:
Here the middle functor is exact by Proposition E.3 since the category has weak kernels. ∎
Consider the functor which sends to considered as an almost -module.
Lemma E.5.
is a cohomological functor which commutes with infinite coproducts. Moreover, is conservative.
Proof.
It is clear that is a cohomological functor. Let be a collection of objects of By the definition of compact morphisms, we see that the -module
is almost zero. This shows that commutes with coproducts.
To show that is conservative, it suffices to check that if then But the vanishing of means that any compact morphism is zero. This implies that is zero.∎
Definition E.6.
Let be a morphism in
1) is phantom if for any compact morphism we have
2) is a pure monomorphism if the map is phantom.
3) is a pure epimorphism if the map is phantom.
Note that is a phantom morphism resp. a pure monomorphism resp. a pure epimorphism in if and only if the morphism is zero resp. a monomorphism resp. an epimorphism in
Definition E.7.
We say that an object is pure injective if for any phantom morphism and for any morphism we have Equivalently, for any pure monomorphism and for any map there exists a map such that
We refer to [Kr00, GP04] for the usual notions of phantom maps, pure monomorphisms and epimorphisms, and pure injective objects for compactly generated triangulated categories. We obtain the following description of the category of pure injective objects, which is a straightforward generalization of the corresponding result for compactly generated categories [GP04, Theorem 4.1].
Proposition E.8.
The functor induces an equivalence from the category of pure injective objects of to the category of injective objects of
Proof.
Let be an injective object of Consider the functor given by This is a cohomological functor (since is injective), which commutes with products. Hence, it is represented by some object Moreover, is a pure injective object of since annihilates phantom morphisms. We obtain a functor from the category of injective almost -modules to the category of pure injective objects of Note that we have a functorial isomorphism
Let now be a pure injective object, and choose a monomorphism where is injective. Then the map is a pure monomorphism. Since is pure injective, is a direct summand of Hence, is a direct summand of We conclude that is injective.
It remains to recall that is conservative, hence the map is an isomorphism for any pure injective ∎
Recall that if is an additive category with weak kernels, then a right -module is called fp-injective if for any finitely presented right -module we have
Proposition E.9.
Let be a small triangulated category and an additive functor. The following are equivalent.
is cohomological.
is flat as an -module.
is fp-injective as an -module.
Proof.
For an object we denote by the corresponding representable right -module.
Flatness of means that for any finitely presented module any morphism factors through a representable module. Let for some morphism in Then the map is given by an element of Note that is a weak cokernel of in Hence, the map factors through a representable module iff is in the image of the map We conclude that is flat if and only if is cohomological.
Let be a finitely presented module, for some morphism in Let Then we have an exact sequence of -modules
It follows that We conclude that is fp-injective if and only if is cohomological. ∎
We say that an almost module is almost cohomological if it is isomorphic to an image of a cohomological -module. Equivalently, this means that the -module is flat.
Proposition E.10.
The functor takes values in almost cohomological almost -modules.
Proof.
Indeed, for any object the -module is flat. ∎
We can now prove a generalization of Adams representability theorem for -compact dualizable categories.
Theorem E.11.
Let be a dualizable category which is -compact in Denote by the homotopy category of Then the functor is full and conservative, and its essential image is exactly the full subcategory of almost cohomological almost modules. More precisely, for we have a functorial short exact sequence of abelian groups:
In particular, the abelian group of phantom maps is identified with The composition of any two phantom maps is zero.
Note that if is compactly generated, then is an -compact object of if and only if the triangulated category is countable. Thus, Theorem E.11 is a generalization of Neeman’s theorem [Nee97, Theorem 5.1, Proposition 4.11] if we restrict to triangulated categories with -enhancement.
Proof.
Essentially the theorem follows from the following statement.
Claim. For any the object has injective dimension at most
Assume that Claim holds. Take some objects and choose an injective resolution of of length
Consider the pure injective objects and the map (where the functor is defined in the proof of Proposition E.8). We obtain a (non-unique) map which becomes an isomorphism after applying Since is conservative, we obtain an exact triangle
It follows that we have a short exact sequence
By construction of the functor we see that the sequence E.2 is exactly of the form (E.1).
In particular, we see that any phantom map factors through (and vice versa). If is another phantom map, then since is pure injective.
Proof of Claim..
By Theorem C.6 there exists a fully faithful strongly continuous functor where is -compact in Denote by the homotopy category of By Proposition 1.6, the category is countable, i.e. it has at most countably many isomorphism classes of objects, and the sets of morphisms are at most countable.
Since the functor preserves compact morphisms, we have a well-defined quotient functor
Its left adjoint is fully faithful. Moreover, takes almost cohomological almost modules to flat modules, because the extension of scalars functor preserves flatness. We also observe that is exact: composing with the conservative exact functor we obtain the functor which is isomorphic to the composition of exact functors
Now, applying Proposition E.9, we see that is fp-injective. Since is countable, it follows that has injective dimension at most Since is exact, the functor preserves injective objects. Since is also exact, we conclude that has injective dimension at most ∎
Theorem is proved. ∎
Remark E.12.
One can also prove Theorem E.11 without choosing an embedding One can define the notion of fp-injective objects in any compactly assembled Grothendieck abelian category (equivalently, a Grothendieck abelian category satisfying (AB6)), such that the functor is exact. Namely, an object is fp-injective if for any and for any compact morphism the map
is zero. If moreover is a Serre subcategory of then one can show that any fp-injective object of has injective dimension at most
Now, for any dualizable category an almost -module is almost cohomological if and only if it is fp-injective in If is -compact, it can be shown that is a Serre subcategory in This gives an alternative (intrinsic) way to show that has injective dimension at most for any
Remark E.13.
If is a Grothendieck abelian category with (AB6) and (AB4*), then one can define flat objects of intrinsically. Namely, an object is flat if for any compact morphism and for any object the map is zero.
If is a dualizable category, then a right almost -module is almost cohomological if and only if it is flat in
Appendix F Analogy between dualizable categories and compact Hausdorff spaces
Denote by the category of compact Hausdorff spaces, and denote by the full subcategory of profinite sets. We have a functor By Corollary 6.15, the category is compactly generated if and only if is profinite.
It turns out that a lot of statements about dualizable categories discussed in this paper have very natural analogues for compact Hausdorff spaces. The table below summarizes this analogy. Here for a topological space we denote by its set of quasi-components, with the quotient topology. Recall that quasi-components of are equivalence classes of points, where if any open-closed subset of containing also contains
| The category of dualizable categories | The opposite category of compact Hausdorff spaces |
| The full subcategory of compactly generated categories | The full subcategory of profinite sets |
| The functor | The functor (the space of quasi-components) |
| For we have | For we have i.e. quasi-components are connected |
| The category is generated by colimits by the single compact object | The category is generated by colimits by the single compact object |
| The category is generated by colimits by the single -compact object | The category is generated by colimits by the single -compact object |
| The category satisfies weak (AB4*) and (AB6) | The category satisfies weak (AB4*) and (AB6) |
| Weak (AB5) holds in the class of fully faithful functors is closed under filtered colimits | Weak (AB5*) holds in the class of surjective maps is closed under cofiltered limits |
| The functor commutes with filtered colimits | The functor commutes with cofiltered limits |
| The inclusion commutes with infinite products | The inclusion commutes with infinite coproducts |
| In sequential colimits do not commute with finite limits | In sequential limits do not commute with finite colimits |
| Localizations (=quotient functors=epimorphisms) in | Injective maps (=monomorphisms) in |
| Fully faithful functors (=monomorphisms) in | Surjective maps (=epimorphisms) in |
| Localizations are stable under pullbacks in | Injective maps are stable under pushouts in |
| Fully faithful functors are stable under pushouts in | Surjective maps are stable under pullbacks in |
| The non-full inclusion is conservative and commutes with colimits | The (faithful) forgetful functor is conservative and commutes with limits |
| Pullbacks of localizations are preserved by the functor | Pushouts of injective maps are preserved by the forgetful functor |
Below we prove the statements from the right column of Table LABEL:tab:analogy which are non-standard. Consider the category of Banach vector spaces over where the morphisms are contractive linear maps (i.e. the bounded maps of norm ). We have a functor
Lemma F.1.
The category satisfies weak (AB4*) and (AB6).
Proof.
The product in of a family is given by the subspace of the naive product, formed by elements such that with the norm given by
Epimorphisms in are maps with a dense image. It follows that satisfies weak (AB4*), i.e. the class of epimorphisms is closed under products.
Filtered colimits in are obtained by taking the completion (of the separation of) of the naive filtered colimit. Here the norm on the naive (filtered) colimit is defined by
To show that (AB6) holds in recall the presentable compactly assembled category (of prenormed vector spaces and contractive maps) from Example 1.34. Since is compactly assembled, it satisfies (AB6). It remains to observe that the completion functor commutes with all colimits and with infinite products.∎
Lemma F.2.
The functor is conservative, detects epimorphisms and commutes with products and with filtered colimits.
Proof.
The functor is conservative by Urysohn’s lemma. By Tietze extension theorem, the functor detects epimorphisms (and also detects regular epimorphisms).
Now, the product of in is given by – Stone-Čech compactification of the disjoint union. For any topological space continuous functions on are identified with bounded continuous functions on It follows that the functor commutes with products.
Finally, consider a codirected system where is directed, and put By Stone-Weierstrass theorem, the union of the images is a dense subalgebra. It remains to observe that the map from the naive colimit to is isometric. Indeed, this follows from the fact that for any and for any neighborhood of there is some such that It follows that the functor commutes with filtered colimits. ∎
Corollary F.3.
The category satisfies weak (AB4*) and (AB6).
Proposition F.4.
The functor commutes with coproducts.
Proof.
Consider a family of profinite sets We need to show that the space is totally disconnected. It suffices to show that the locally constant -valued functions on are dense in the space of all continuous functions.
Consider a function corresponding to a family of continuous functions such that for some we have Take some and choose an integer Since is profinite, we can find a locally constant function such that and takes values in the finite set Then the corresponding function is also locally constant and ∎
Proposition F.5.
The category is generated by colimits by the closed unit interval which is an -compact object. In particular, the category is -presentable, and the -compact objects are precisely the metrizable spaces.
Proof.
To see that the object is -compact in it suffices to observe that -filtered colimits in can be computed naively, i.e. the naive colimit is already complete (and separated). By Urysohn’s lemma, the functor is conservative. Hence, the object generates the category by colimits.
Now, the full subcategory of metrizable spaces in contains the final object and is closed under fiber products and countable products. Hence, metrizable spaces in are closed under -small colimits. It remains to show that they are generated by via -small colimits.
Let be a metrizable compact Hausdorff space. Choose an injective map Denote by the pushout Then is also metrizable, and we choose an injective map Composing it with the two maps we obtain two self-maps The equalizer is isomorphic to ∎
Proposition F.6.
The functor commutes with cofiltered limits.
Proof.
Consider a codirected system and put for Passing to Boolean algebras, we need to show that the map is an isomorphism. Given an open-closed consider its characteristic function Then for some we have a continuous function such that Then is an open-closed subset of and its preimage in equals
Suppose that is an open-closed subset such that Then hence for some we have i.e. ∎
Proposition F.7.
Sequential limits in do not commute with finite colimits.
Proof.
Consider the surjective map Put and Then
Put Then the diagram is the limit of but is a one-element set for Hence, the sequential limits do not commute with coequalizers in ∎
Appendix G K-theory of products of exact infinity-categories
In this section we prove that -theory of exact -categories commutes with infinite products.
G.1. Products of additive infinity-categories
Given a small additive -category we denote by its stabilization. By [Bon, Theorem 5.3.1, Proposition 6.2.1], if is Karoubi complete, then so is Given we denote by the full subcategory generated by We put and similarly
The following result is due to A. Córdova Fedeli.
Theorem G.1.
[Cor23, Proposition 2.10] Let be a collection of Karoubi complete additive small -categories. We have an isomorphism in
Since we will need an analogous statement for -structures, we recall the proof from [Cor23] with some simplifications.
We introduce some notation. Put Given functions we put
For any function we denote by the product of the functors
Lemma G.2.
2) Analogous statement holds for the same functors
Proof.
In both cases, we have the following equalities in
This proves the lemma.∎
Lemma G.3.
[Cor23, Proof of Proposition 2.7] The map is an isomorphism in
Proof.
Since (AB6) holds in we have
Hence, it suffices to prove that for any the map
is an isomorphism. This follows from Lemma G.2.∎
Lemma G.4.
Proof.
Essential surjectivity of . Take some object Then for some we have for all Choose exact triangles
Then and the map becomes an isomorphism in This proves that is essentially surjective.
Let now Let be objects of and denote by and their images in
Fullness of Consider a map represented by a roof where and Take such that and Choose and the maps such that Then we can choose factorizations and automatically
Putting we have and Therefore, the roof represents a morphism such that
Faithfulness of Consider a map such that We may assume that can be lifted to a morphism Then there exists a map such that and Arguing as above, we can construct a factorization of such that and Hence ∎
G.2. Products of stable categories with bounded -structures
We will need the following analogue of Theorem G.1. Let be a family of stable categories with bounded -structures. Denote by the full subcategory formed by objects such that the objects have uniformly bounded homological amplitude.
Theorem G.5.
We have an isomorphism
Proof.
The proof is completely analogous to the proof of Theorem G.1 with minor changes, which we indicate.
For a stable category with a bounded -structure and for integers we denote by the full subcategory formed by objects whose homology is concentrated in degrees between and and similarly for
Analogously to Lemma G.3, one shows that we have an isomorphism
Analogously to Lemma G.4, one proves that for any function the functor is an equivalence.
Then one concludes that proving the theorem.∎
G.3. Products of exact -categories
Recall the notion of an exact -category.
Definition G.6.
[Bar15, Definition 3.1] An exact -category is a triple where is an additive -category and are subcategories such that the following conditions hold.
1) Every morphism (resp. ) is in (resp. ).
2) Pushouts of morphisms in exist, and the class of morphisms in is closed under pushouts. Pullbacks of morphisms in exist and the class of morphisms in is closed under pullbacks.
3) For a commutative square
the following are equivalent:
a) the square is coCartesian, and
b) the square is Cartesian, and
By definition, a short exact sequence is a biCartesian square of the form
where and A total cofiber of such a square is by definition a primitive acyclic object. The full subcategory of acyclic objects is the stable subcategory generated by the primitive acyclic objects. The stable hull of is the category we follow the notation of [Kle22].
Proposition G.7.
For an exact -category the category has a bounded -structure, and the functor is -exact.
Proof.
Note that if is a triangulated category with a -structure, and is a weak Serre subcategory of the heart, then the full triangulated subcategory generated by has a bounded -structure with the heart In particular, the functor is -exact.
Theorem G.8.
Let be a family of small idempotent-complete exact -categories. We have an isomorphism
Corollary G.9.
-theory of exact -categories commutes with infinite products.
Proof of Theorem G.8..
We have a map of short exact sequences of stable categories:

### v3

@@
 We also study the general theory of dualizable categories.
-In particular, we give an “explicit” proof of Ramzi’s theorem [Ram] on the -presentability of the category of dualizable categories.
+In particular, we give an “explicit” proof of Ramzi’s theorem [Ram24a] on the -presentability of the category of dualizable categories.
 Among other things, we prove that dualizability is equivalent to “flatness” in the category of presentable stable categories.
@@
 We denote by the (non-full) subcategory of dualizable categories and strongly continuous functors.
-Further, we denote by the subcategory of compactly generated categories.
+Further, we denote by the full subcategory of compactly generated categories.
 It is well-known that the functor is an equivalence, and an inverse equivalence is given by (the latter is the full subcategory of compact objects).
@@
 Another class of examples which we are interested in this paper are categories of sheaves on a locally compact Hausdorff space Let be a presheaf on with values in We denote by the category of sheaves with values in By [Lur18], the category is dualizable.
-We have the following general result, see 6.10.
+We have the following general result, see Theorem 6.11.
 Theorem 0.2.
@@
 In Section 2 we prove a surprising result: a presentable stable category is dualizable if and only if it is flat in More precisely, is flat if the functor preserves fully faithful functors.
+We also give a more general criterion for flatness in for an uncountable regular cardinal
 In Section 3 we give a criterion of dualizability for categories which are extensions of compactly generated categories.
 If such an extension is dualizable, then it is in fact compactly generated.
-As an application, we show that in the extensions of by are classified by functors where is the category of Tate objects, see [Hen17].
+As an application, we recover a result from [CDH+20, Appendix A] stating that in the extensions of by are classified by functors where is the category of Tate objects, see [Hen17].
 As another application, for a commutative noetherian ring we classify the localizing subcategories of which are dualizable.
@@
 In Subsection 4.3 we explain that this canonical extension is in fact the right Kan extension.
-In Subsection 4.4 we give an example: we compute finitary localizing invariants of the category of sheaves on and on In Subsection 4.5 we do a similar computation for not necessarily finitary localizing invariants, deducing the commutation of K-theory with products (Theorem 4.29).
+In Subsection 4.4 we give an example: we compute finitary localizing invariants of the category of sheaves on and on In Subsection 4.5 we do a similar computation for not necessarily finitary localizing invariants, deducing the commutation of K-theory with infinite products (Theorem 4.28).
 In Section 5 we consider the sheaves and cosheaves on continuous posets (these posets are considered as categories, not as sets), and compute their (finitary) localizing invariants.
@@
 In Subsection 6.2 we explain how to approximate the category of sheaves on an arbitrary compact Hausdorff space by “simpler” categories (Proposition 6.10).
-Here it is crucial that we use the notion of a -(pre)sheaf from [Lur09].
+Here it is crucial that we use the notion of a -(pre)sheaf from [Lur09a].
 In Subsection 6.3 we apply the previous results to compute the finitary localizing invariants of categories of sheaves (Theorem 6.11).
@@
 In Appendix B we show that an image of a homological epimorphism does not have to be a stable subcategory (Proposition B.1).
-In Appendix C we show that the category is -presentable, and we describe the -compact objects for each uncountable regular (Theorem C.6).
-The -presentability of is originally due to Ramzi [Ram].
-In Appendix D we prove that the category is generated by colimits by a single object – the category of sheaves of spectra on the real line with singular support in In Appendix E we prove an analogue of Adams representability theorem for -compact dualizable categories (Theorem E.11).
+In Appendix C we show that the category is -presentable, and we describe the -compact objects for each uncountable regular cardinal (Theorem C.6).
+The -presentability of is originally due to Ramzi [Ram24a].
+In Appendix D we prove that the category is generated by colimits by the single object – the category of sheaves of spectra on the real line with singular support in In Appendix E we prove two closely related versions of Adams representability theorem for -compact dualizable categories (Theorems E.2 and E.4).
 In Appendix F we explain an analogy between the category of dualizable categories and the opposite category of compact Hausdorff spaces.
-Finally, in Appendix G we show that -theory of exact -categories commutes with infinite products, using the ideas of Córdova Fedeli [Cor23] and Klemenc [Kle22].
+Finally, in Appendix G we show that -theory of exact -categories commutes with infinite products, using the results of Córdova Fedeli [Cor23] and Klemenc [Kle22].
 Acknowledgements.
-I am grateful to Ko Aoki, Alexander Beilinson, Dustin Clausen, Adriano Córdova Fedeli, Vladimir Drinfeld, Boris Feigin, Dennis Gaitsgory, Marc Hoyois, Dmitry Kaledin, David Kazhdan, Maxim Kontsevich, Akhil Mathew, Thomas Nikolaus, Dmitri Orlov, Maxime Ramzi, Victor Saunier, Tomer Schlank, Peter Scholze, Vivek Shende, Vladimir Sosnilo, Georg Tamme, Bertrand Toën, Yakov Varshavsky and Christoph Winges for useful discussions.
+I am grateful to Ko Aoki, Alexander Beilinson, Dustin Clausen, Adriano Córdova Fedeli, Vladimir Drinfeld, Boris Feigin, Dennis Gaitsgory, Marc Hoyois, Dmitry Kaledin, David Kazhdan, Maxim Kontsevich, Akhil Mathew, Thomas Nikolaus, Dmitri Orlov, Maxime Ramzi, Victor Saunier, Tomer Schlank, Peter Scholze, Vivek Shende, Vladimir Sosnilo, German Stefanich, Georg Tamme, Bertrand Toën, Yakov Varshavsky and Christoph Winges for useful discussions.
 Part of this work was done while I was a visitor in the Max Planck Institute for Mathematics in Bonn from December 2022 till February 2024, and I am grateful to the institute for their hospitality and support.
@@
 1.1. Infinity-categories, limits and colimits
-We will freely use the theory of -categories, functors, limits and colimits as developed in [Lur09, Lur17].
+We will freely use the theory of -categories, functors, limits and colimits as developed in [Lur09a, Lur17].
 Except for Section 2, we will deal only with -categories.
@@
 If is regular, then a set is called -small if
-We refer to [Lur09, Definition 5.3.1.7] for the notion of a -filtered -category.
-The argument of [AR94, Theorem 1.5, Remark 1.21] shows that for any -filtered -category there exists an -directed poset and a cofinal functor Recall that a poset is called -directed if is nonempty and any -small collection of elements of has an upper bound.
+We refer to [Lur09a, Definition 5.3.1.7] for the notion of a -filtered -category.
+By [Lur09a, Proposition 5.3.1.18] for any -filtered -category there exists a -directed poset and a cofinal functor Recall that a poset is called -directed if is nonempty and any -small collection of elements of has an upper bound.
 Recall that -filtered -categories resp.
@@
 Equivalently, this means that where the category has -small colimits.
-Further, is presentable if it is -presentable for some Following [Lur09], we denote by the -category of all presentable categories and colimit-preserving functors.
+Further, is presentable if it is -presentable for some Following [Lur09a], we denote by the -category of all presentable categories and colimit-preserving functors.
 Note that the category is not locally small.
@@
 For any regular cardinal the inclusion functor commutes with colimits. For the right adjoint to the functor is given by
-We denote by the full subcategory of presentable stable categories, and similarly for Note that the category of spectra is an idempotent -algebra in and the subcategory is simply the category of modules over (and similarly for ).
-In particular, the inclusions and commute with limits and colimits.
+We denote by the full subcategory of presentable stable categories, and similarly for The inclusions and commute with limits and colimits.
 Below we will mostly deal with presentable stable categories.
@@
 It is sufficient to prove the “moreover” statement.
-Suppose that and are -presentable and commutes ith -filtered colimuts.
+Suppose that and are -presentable and commutes with -filtered colimuts.
 Then an object is -compact in whenever We conclude that is -presentable, and the full subcategory is generated by and (as a stable subcategory).
@@
 In particular, the category is compactly generated and iff and the group is finitely generated. If is an uncountable regular cardinal, then is a -compact object if and only if and is -small.
-We will need an explicit characterizations of -compact objects in for uncountable
+We will need an explicit characterization of -compact objects in for uncountable
 Proposition 1.6.
@@
 We call an exact functor between presentable stable -categories strongly continuous if it has a right adjoint which is continuous (in other words, if has a twice right adjoint).
-It is convenient to introduce a non-full subcategory with the same objects and with -morphisms being strongly continuous functors.
+It is convenient to introduce the non-full subcategory with the same objects and with -morphisms being strongly continuous functors.
 Note that is locally small.
@@
 Hence, is continuous if and only if all are continuous.∎
-Remark 1.8.
-It seems plausible that for any regular cardinal the category is presentable. However, it is not clear how to describe the functor right adjoint to the inclusion.
+It follows from Proposition 1.72 that the category is not presentable, more precisely, the object is not -compact for any regular cardinal A different example showing this was explained to me by German Stefanich.
 We make an observation on the automatic strong continuity.
@@
 Given a presentable stable category we will call a full subcategory localizing if is presentable stable and the inclusion functor commutes with colimits.
-Equivalently, this means that is generated as a cocomplete stable subcategory by a set of objects (or by a single object).
+Equivalently, this means that is generated as a cocomplete stable subcategory by a set of objects (or equivalently by a single object).
 1.3. Well generated triangulated categories
@@
 If is presentable, then a full subcategory is localizing if and only if is localizing.
-Recall that for a triangulated category an additive functor is called cohomological if for any exact triangle the sequence is exact.
+Recall that for a triangulated category an additive functor is called homological if for any exact triangle the sequence is exact.
+It is convenient to call a functor with the same property cohomological.
 By [Nee01a, Theorem 1.17], a well generated triangulated category satisfies Brown representability for contravariant functors: an additive functor is representable if and only if is cohomological and commutes with products.
@@
 A much more difficult question is the Brown representability for covariant functors, and it is not known if it holds for abstract well generated categories.
-But if where is presentable stable, then it follows from [Nee09, Theorems 1.11, 1.17 and 1.22] and [Lur09, Proposition A.3.7.6] that an additive functor is corepresentable if and only if is cohomological and commutes with products.
+But if where is presentable stable, then it follows from [Nee09, Theorems 1.11, 1.17 and 1.22] and [Lur09a, Proposition A.3.7.6] that an additive functor is corepresentable if and only if is homological and commutes with products.
 In particular, for such and for any well generated triangulated category an exact functor has a left adjoint iff commutes with products.
@@
 is a retract in of a compactly generated category.
-Let be another presentable stable category and be a continuous localization, i.e. the right adjoint is fully faithful. Then has a continuous section.
+Let be another presentable stable category and let be a continuous localization, i.e. the right adjoint is fully faithful. Then has a continuous section.
 Remark 1.12.
 Given a dualizable category we have a natural functor which can be described (at least) in two ways. The first one is the following: given the object corresponds to the continuous functor
-Another description is via the universal problem: we have
+Another description is via the universal property: we have
 In the special case where is an -ring, we have – the category of left -modules. For a right -module we have – the usual dual left module of a right module.
@@
 The compact objects are given by the usual category of perfect complexes Moreover, the category is closely related with another category which can be defined as an inverse limit in
-We refer to Subsection 1.15 for a discussion on general inverse limits in We will study both categories and in great detail in the forthcoming paper [E1].
+We refer to Subsection 1.15 for a discussion of general inverse limits in We will study both categories and in great detail in the forthcoming paper [E1].
 In particular, we will prove that there is a natural fully faithful strongly continuous functor which induces equivalences on compact objects, and we have
@@
 Example 1.33.
-Denote by the category of prenormed -vector spaces, where the morphisms are given by contractive linear maps (i.e.
+Denote by the category of seminormed -vector spaces, where the morphisms are given by contractive linear maps (i.e.
 the maps such that ).
 Then is presentable and compactly assembled.
-Indeed, the category has colimits, and it is generated by colimits by finite-dimensional normed vector spaces and by the object (-dimensional vector space with zero prenorm).
+Indeed, the category has colimits, and it is generated by colimits by finite-dimensional normed vector spaces and by the object (-dimensional vector space with zero seminorm).
 We have
@@
 However, the criterion for dualizability does not follow formally from the results in loc.cit.
-Let be a presentable stable category. We give the following definition.
+Definition 1.38 and a version of Theorem 1.45 below were explained to me by Dustin Clausen.
+Proposition 1.48 seems to be new.
+Let be a presentable stable category.
 Definition 1.38.
@@
 Proof.
-. Suppose that where is directed, and take some map Dnote by the identity map
+. Suppose that where is directed, and take some map Denote by the identity map
 Since is compact, the composition factors through some object It follows that the composition is zero, as required.
@@
 . This is clear: apply the definition of compactness to the sequence
-. Consider a directed system and a map Since is -compact, we may assume that Indeed, -compactness of implies that for some sequences of elements of the map factors through
+. Consider a directed system and a map Since is -compact, we may assume that Indeed, -compactness of implies that for some sequence of elements of the map factors through
 Assume Recall that we have a cofiber sequence
@@
 and an -bimodule is a biadditive bifunctor Recall that an ideal of a small additive category is a subbimodule of the diagonal bimodule More generally, a quasi-ideal in is a bimodule together with a map (not necessarily injective) such that for any we have
-We have the usual notion of a tensor product of left and right -modules.
-In particular, given bimodules we have a bimodule given by Note that the tensor product of quasi-ideals is naturally a quasi-ideal.
-We say that an bimodule is flat on the right if each right -module is flat,
+We have the usual notion of a tensor product of left and right -modules. In particular, given bimodules we have a bimodule given by
+Note that the tensor product of quasi-ideals is naturally a quasi-ideal. We say that an bimodule is flat on the right if each right -module is flat,
 Given a quasi-ideal we have a natural map We say that a quasi-ideal is idempotent if is an isomorphism.
@@
 Let be a small additive category. The assignment gives a bijection
-The inverse map sends an ideal to the quasi-ideal where the map is given by the multiplication.
+The inverse map sends an ideal to the quasi-ideal where the map is given by the multiplication (i.e. composition).
 Proof.
-Let be an idempotent quasi-ideal in and let It is clear that We need to show that as quasi-ideal. Note that if are elements such that or then It follows that the isomorphism factors through hence
-Conversely, let be an ideal in such that and is flat on the right. We need to show that is idempotent. It suffices to show that the map
+Let be an idempotent quasi-ideal in and let It is clear that We need to show that as quasi-ideal.
+Note that if are elements such that or then It follows that the isomorphism factors through hence Conversely, let be an ideal in such that and is flat on the right.
+We need to show that is idempotent.
+It suffices to show that the map
 is an isomorphism. Since the map (1.7) is surjective. On the other hand, since is flat on the right, the map (1.7) is injective. Hence, this map is an isomorphism. ∎
@@
 The statement of Lemma 1.47 still holds, and the proof is the same.
+If is a compactly assembled accessible -category (not necessarily cocomplete), then the analogue of Proposition 1.48 still holds, and the proof is the same.
+If is an ordinary category, then is the quasi-ideal of wavy arrows in the terminology of [JJ82, Section 2].
 Remark 1.50.
@@
 Definition 1.51.
-Let be an -category with infinite products and filtered colimits. We refer to the following condition as the axiom (AB6):
-- for any collection of filtered categories and for any functors the map
+Let be an -category with infinite products and filtered colimits. We refer to the following condition as the axiom (AB6): - for any collection of filtered categories and for any functors the map
 is an isomorphism.
@@
 Proof.
-Suppose that is dualizable. Then the colimit functor has a left adjoint Hence, the functor commutes with infinite products. By Proposition 1.53, this exactly means that (AB6) holds in
-Conversely, suppose that holds in Arguing as above, we see that the functor commutes with infinite products.
-Now, the category is not presentable, so we need the following additional observation to deduce the existence of the left adjoint to
-Let be a regular cardinal such that the category is -compactly generated.
+Suppose that is dualizable.
+Then the colimit functor has a left adjoint Hence, the functor commutes with infinite products.
+By Proposition 1.52, this exactly means that (AB6) holds in Conversely, suppose that holds in Arguing as above, we see that the functor commutes with infinite products.
+Now, the category is not presentable, so we need the following additional observation to deduce the existence of the left adjoint to Let be a regular cardinal such that the category is -compactly generated.
 Denote by the colimit functor.
@@
 If is a complete lattice, then this statement is well known, see [GHKLMS80, Theorem I.2.3].
+Below we will see that the category satisfies (AB6) (Proposition 1.106) but the strong (AB5) does not hold in (Corollary 1.86).
 1.10. Automatic dualizability and images of functors
@@
 We first reduce to the case when consists of a single element.
-Consider the category Then is dualizable: it is sufficient to prove this for compactly generated categories, and in this case we have By Proposition 1.7, the induced functor is strongly continuous and its image generates
-So assume that and put We give two proofs.
+Consider the category Then is dualizable: it is sufficient to prove this for compactly generated categories, and in this case we have By Proposition 1.7, the induced functor is strongly continuous and its image generates So assume that and put We give two proofs.
 First proof.
@@
 Since the axiom (AB6) holds in we deduce that it also holds in By Proposition 1.53, is dualizable.
-Second proof. Let be a generating object, and Then for an ind-object we have an equivalence
+Second proof.
+Let be a generating object, and Then for an ind-object we have an equivalence
 Hence, the (a priori partially defined) functor is defined on Since the objects generates by assumption, we conclude that the functor exists, hence is dualizable.∎
@@
 Remark 1.58.
-In the assumptions of Corollary 1.58, if is dualizable, then is also the fiber of the cofiber of in
+Under the assumptions of Corollary 1.57, if is dualizable, then is also the fiber of the cofiber of in
 Remark 1.59.
@@
 1.11. Calkin categories
-Fix an uncountable regular cardinal We will define the Calkin and -Calkin categories as follows.
-Given a small stable Karoubi complete category we put
+Fix an uncountable regular cardinal We will define the Calkin and -Calkin categories as follows. Given a small stable Karoubi complete category we put
 It turns out that this construction has a direct analogue for dualizable categories.
@@
 For a dualizable presentable stable category we define its continuous -Calkin category as the category of compact objects in the quotient We have a short exact sequence
-in
-We have fully faithful functors for We denote by the colimit of
+in We have fully faithful functors for We denote by the colimit of
 We observe that this Calkin construction is functorial: we have a well-defined functor Indeed, if is a strongly continuous functor between dualizable categories, then we have a commutative square
@@
 Remark 1.62.
-Note that if is compactly generated, then the essential image of the functor is a stable subcategory: it is equivalent to the Verdier quotient However, in general this essential image does not have to be a stable subcategory, see Proposition B.1 below.
+Note that if is compactly generated, then the essential image of the functor is a stable subcategory: it is equivalent to the Verdier quotient However, in general this essential image does not have to be a stable subcategory even for see Proposition B.1 below.
 Proposition 1.63.
@@
 Proof.
-The second statement follows from the first one by Proposition 1.7.
 We need the following general facts about the directed colimits in and
@@
 Proof.
-We use once again the natural equivalence where the transition functors in the inverse system are Moreover, the composition is identified with the functor
-To prove 1), we simply observe that for we have isomorphisms
+We use once again the natural equivalence where the transition functors in the inverse system are Moreover, the composition is identified with the functor To prove 1), we simply observe that for we have isomorphisms
+We prove 2). By Proposition 1.7 the functors are continuous. Using again the continuity of we obtain a well-defined functor given by
+Then (1.8) is equivalent to the statement that To check this, take some objects We obtain the following isomorphisms:
+Hence, as required.∎
 Proposition 1.67.
@@
 Proof of Proposition 1.67.
-By Proposition 1.66, it is sufficient to prove the statement for
-Consider a filtered system of strongly continuous fully faithful functors where the transition functors and are strongly continuous.
+By Proposition 1.65, it is sufficient to prove the statement for Consider a filtered system of strongly continuous fully faithful functors where the transition functors and are strongly continuous.
 Put and denote by the colimit of Denote by the structural functors.
@@
 Corollary 1.69.
-Let be a presentable stable category, and suppose we have a directed family of presentable stable full subcategories where the inclusion functors are strongly continuous.
+Let be a presentable stable category, and suppose that we have a directed family of presentable stable full subcategories where the inclusion functors are strongly continuous.
 If is generated by as a localizing subcategory, then the natural functor is an equivalence.
@@
 Remark 1.70.
-The following example shows that the analogues of Proposition 1.68 and 1.70 fail in Denote by the -th prime number, i.e.
+The following example shows that the analogues of Proposition 1.67 and 1.69 fail in Denote by the -th prime number, i.e.
 Put and where We consider as a full subcategory of via restriction of scalars, and denote by the inclusion functor.
 Hence, we have
-Note that the subcategories generate as a localizing subcategory. Indeed, we have and However, we claim that the functor is not an equivalence.
+Note that the subcategories generate as a localizing subcategory.
+Indeed, we have and However, we claim that the functor is not an equivalence.
 Indeed, otherwise by Proposition 1.66 1) we would have But this clearly fails:
@@
 Proposition 1.71.
-The functor commutes with filtered colimits.
+The functor commutes with filtered colimits. Equivalently, the object is compact.
 Proof.
@@
 Here we used the fact that in filtered colimits commute with finite limits.∎
-The following example shows that Proposition 1.72 does not hold for presentable categories.
+We now give examples showing that the object is not small (i.e.
+not -compact for any regular cardinal ).
+A different argument for this was given previously by German Stefanich, using the generalizations of the derived category of quasi-coherent sheaves on We use the following notation.
+For a pair of ordinals we denote by the version of a long closed interval: we put
+and consider this set as a topological space with the interval topology with respect to the following total order.
+The order on is lexicographic, and the element is the largest one.
+Then is a connected compact Hausdorff space.
 Proposition 1.72.
 Let be a field.
-Denote by the rectangle (considered as a topological space).
+Take some regular cardinal which we also consider as an ordinal.
+For an ordinal denote by the rectangle (considered as a topological space).
 Denote by the full subcategory of sheaves whose pullback to vanishes.
-Then the pullback functors are strongly continuous, the colimit is compactly generated, but we have for
-In particular, the functor does not commute with sequential colimits.
+Then the pullback functors are strongly continuous for the colimit is non-zero and compactly generated, but we have for Therefore, the functor does not commute with -filtered colimits, i.e.
+the object is not -compact.
+In particular, the category is not presentable.
 Proof.
+The description of the category implies that it is -presentable: it is a kernel of a continuous functor between dualizable categories with an -continuous right adjoint.
 It is clear that each functor is strongly continuous.
-Identifying the colimit with the limit of the inverse sequence, we see that is equivalent to the full subcategory of consisting of sheaves whose support is contained in for all (and whose pullback to vanishes).
-This is simply the category of sheaves supported at the point which is equivalent to
-It remains to show that Since all the categories are equivalent, we may assume Suppose that is a non-zero compact object.
-For put The pullback functor is strongly continuous, hence is a compact sheaf on Since is contractible and compact, it follows from Proposition 6.14 that is a constant sheaf.
-Since this holds for arbitrarily small and since we deduce that is a non-zero constant sheaf on .
+Identifying the colimit with the limit of the inverse system, we see that is equivalent to the full subcategory of consisting of sheaves whose support is contained in for all (and whose pullback to vanishes).
+This is simply the category of sheaves supported at the point which is equivalent to It remains to show that Since all the categories are equivalent, we may assume Suppose that is a non-zero compact object.
+For put The pullback functor is strongly continuous, hence is a compact sheaf on Since is connected and compact, it follows from Proposition 6.15 that is a nowhere zero locally constant sheaf.
+Since this holds for arbitrary we deduce that is a nowhere zero locally constant sheaf on .
 The pullback functor is strongly continuous, hence is also a compact sheaf.
@@
 Proposition 1.73.
-The functor commutes with -filtered colimits. In particular, the functor commutes with -filtered colimits.
+Let be an uncountable regular cardinal. The functor commutes with -filtered colimits. In particular, the functor commutes with -filtered colimits.
 Proof.
 Recall that the functor commutes with -filtered colimits.
-Hence, so is the composition It follows that the functor commutes with -filtered colimits.
+Hence, so does the composition It follows that the functor commutes with -filtered colimits.
 Passing to compact objects, we deduce that the functor commutes with -filtered colimits.∎
-We conclude this subsection with the following elementary statement about pushouts.
+We conclude this subsection with the following statement about pushouts.
 Proposition 1.74.
@@
 Definition 1.75.
-A (finite) semi-orthogonal decomposition in (resp. ) of a presentable stable (resp. dualizable) category is a collection of presentable stable (resp. dualizable) subcategories such that
-i) the inclusion functors are strongly continuous;
-ii) For we have
-iii) The categories generate as a stable subcategory.
+A (finite) semi-orthogonal decomposition in (resp.
+) of a presentable stable (resp.
+dualizable) category is a collection of presentable stable (resp.
+dualizable) subcategories such that i) the inclusion functors are strongly continuous; ii) For we have iii) The categories generate as a stable subcategory.
 In this case we write
@@
 Proof.
-If is dualizable, then each is a retract of in hence is dualizable. Conversely, if each is dualizable, then is dualizable by Proposition 1.56.
+If is dualizable, then each is a retract of in hence is dualizable.
+Conversely, if each is dualizable, then is dualizable by Proposition 1.55.
 For the second statement we may assume that Denote by the inclusion functors.
@@
 Proof.
-Note that the functor is automatically strongly continuous.
-. It suffices to recall that where is the right adjoint to
-. Suppose that is continuous. Note that the functor is given by
+Note that the functor is automatically strongly continuous. . It suffices to recall that where is the right adjoint to . Suppose that is continuous. Note that the functor is given by
 Hence, is strongly continuous. The dualizability of follows from Proposition 1.77.∎
@@
 Let be a presentable stable category with a semi-orthogonal decomposition in Denote by the left adjoint to and the right adjoint to Let be another presentable stable category.
-1) A continuous functor is strongly continuous if and only if the compositions and are strongly continuous. We have a semi-orthogonal decomposition
+1) A continuous functor is strongly continuous if and only if the compositions and are strongly continuous.
+We have a semi-orthogonal decomposition
 2) A continuous functor is strongly continuous if and only if the compositions and are strongly continuous. We have a semi-orthogonal decomposition
@@
 The criterion of strong continuity follows from Proposition 1.8 applied to the functors The functors and are given respectively by precomposition with and The left adjoint to is given by precomposition with and the right adjoint to is given by precomposition with It follows formally that and are fully faithful and the composition is zero.
-It remains to observe that the cofiber sequence in implies that we have a cofiber sequence in
-We now prove part 1).
+It remains to observe that the cofiber sequence in implies that we have a cofiber sequence in We now prove part 1).
 The proof of a semi-orthogonal decomposition is similar; here the inclusion functors are given by the compositions with and The “only if” direction of the criterion is clear.
@@
 ∎
-For a future reference, we define the notion of a poset-indexed (possibly infinite) semi-orthogonal decomposition.
+We also define the notion of a poset-indexed (possibly infinite) semi-orthogonal decomposition.
 Definition 1.80.
@@
 in ) is a collection of presentable stable (resp.
-dualizable) subcategories such that
-i) the inclusion functors are strongly continuous;
-ii) For we have
-iii) The categories generate as a localizing subcategory.
+dualizable) subcategories such that i) the inclusion functors are strongly continuous; ii) For we have iii) The categories generate as a localizing subcategory.
 In this case we write
@@
 The first statement is proved in the same way as Proposition 1.77.
-To prove the second statement, we introduce for each finite subset the subcategory generated by By Propositions 1.56 and 1.9 the category is dualizable and the inclusion is strongly continuous.
+To prove the second statement, we introduce for each finite subset the subcategory generated by By Propositions 1.55 and 1.8 the category is dualizable and the inclusion is strongly continuous.
 Moreover, we have a semi-orthogonal decomposition By Proposition 1.77, the category is generated as a stable category by Combining Corollary 1.69 and Proposition 1.71 we deduce that the category is generated as a stable subcategory by Hence, we have a semi-orthogonal decomposition as required.∎
 1.14. Finite limits of dualizable categories
-The limits in the category of dualizable categories are much more complicated then colimits. We start their study with the case of kernels (i.e. fibers).
+The limits in the category of dualizable categories are much more complicated then colimits.
+We start their study with the case of kernels (i.e.
+fibers).
+We denote by the poset of rational numbers with the usual order.
 Proposition 1.82.
@@
 Proof.
-It is convenient to replace the poset with the isomorphic poset
-Choose a sequence in such that and each map is compact.
+It is convenient to replace the poset with the isomorphic poset Choose a sequence in such that and each map is compact.
 Define the functor by the formula Since any compact morphism in is a composition of two compact morphisms, we can define inductively the functors such that and for any such that the map is compact.
-The sequence defines a functor
-Finally, we define a functor by the formula Then satisfies the required properties (since is a dense linearly ordered set).∎
+The sequence defines a functor Finally, we define a functor by the formula Then satisfies the required properties (since is a dense linearly ordered set).∎
 The following lemma will be useful.
@@
 Proof.
-Denote by the full subcategory generated (as a localizing subcategory) by the objects described in the second assertion. We first prove that is dualizable and the inclusion is strongly continuous.
-Consider a functor as above. We may and will assume that for each rational the map is an equivalence. Indeed, as in the proof of Proposition 1.83 we may replace by the functor given by
-Now, we observe that our assumptions on imply that each object is contained in (because the linearly ordered set is isomorphic to ).
+Denote by the full subcategory generated (as a localizing subcategory) by the objects described in the second assertion.
+We first prove that is dualizable and the inclusion is strongly continuous.
+Consider a functor as above.
+We may and will assume that for each rational the map is an equivalence.
+Indeed, as in the proof of Proposition 1.82 we may replace by the functor given by Now, we observe that our assumptions on imply that each object is contained in (because the linearly ordered set is isomorphic to ).
 It follows that the functor exists and we have Moreover, the inclusion commutes with hence it is strongly continuous.
@@
 We claim that is contained in To see this, take any countably presented object Choose a functor such that each map is compact in (hence in ) and We see that as required.∎
-We can give a description of kernels in
+We can now give a description of kernels in
 Proposition 1.84.
@@
 Example 1.85.
-Let be a discrete valuation ring with the maximal ideal and with the field of fractions Then Indeed, the usual kernel is given by but the inclusion functor is strongly continuous.
+Let be a discrete valuation ring with the maximal ideal and with the field of fractions Then Indeed, the usual kernel is given by but the inclusion functor is not strongly continuous.
 The only localizing subcategory of which is not equal to is zero.
@@
 Proposition 1.87.
-Consider a pair of strongly continuous functors between dualizable categories
-1) The fiber product exists in and the natural functor is fully faithful and strongly continuous.
+Consider a pair of strongly continuous functors between dualizable categories 1) The fiber product exists in and the natural functor is fully faithful and strongly continuous.
 More precisely, the category is the largest dualizable subcategory such that the functors and are strongly continuous.
@@
 Proof.
-1) By Proposition 1.79 the lax fiber product is dualizable. It follows from Proposition 1.80 that where The usual fiber product (taken in ) is simply The rest follows from Proposition 1.85.
+1) By Proposition 1.78 the lax fiber product is dualizable.
+It follows from Proposition 1.79 that where The usual fiber product (taken in ) is simply The rest follows from Proposition 1.84.
 2) It is well-known that in a pullback of a localization is a localization.
@@
 By Proposition 1.87 we only need to show that the category is compactly generated.
-It suffices to show that the objects of the form are generated by colimits by the essential image of the diagonal functor
-By Proposition 1.20, we may assume that is of the form where and is a zero map in Then we have
+It suffices to show that the objects of the form are generated by colimits by the essential image of the diagonal functor By Proposition 1.19, we may assume that is of the form where and is a zero map in Then we have
 where the maps are chosen (non-uniquely) in such a way that they induce the pairs of maps This proves the proposition.∎
@@
 We construct the adjunction unit and counit as follows.
-Let be a dualizable stable category. Then the adjunction unit on is given by the functor (we know that this functor is strongly continuous).
-Let be an -presentable category. Then the adjunction counit is given by the colimit functor Note that this functor is -strongly continuous.
-If is dualizable, then the composition is isomorphic to the identity functor. If is -presentable, then the composition
+Let be a dualizable stable category.
+Then the adjunction unit on is given by the functor (we know that this functor is strongly continuous).
+Let be an -presentable category.
+Then the adjunction counit is given by the colimit functor Note that this functor is -strongly continuous.
+If is dualizable, then the composition is isomorphic to the identity functor.
+If is -presentable, then the composition
 is also isomorphic to the identity. This proves the adjunction. ∎
 Remark 1.90.
-Given a dualizable category and an -presentable category an -strongly continuous functor corresponds via the above adjunction to the composition
+Given a dualizable category and a -presentable category (where is uncountable), a -strongly continuous functor corresponds via the above adjunction to the composition
 Theorem 1.91.
-Let be an -category, and be a functor, Then the limit of exists, and is given by
+Let be an -category, and be a functor, Then the limit of exists, and for any uncountable regular cardinal we have an equivalence
 Proof.
@@
 It follows formally that exists and it is the largest dualizable subcategory
-such that the inclusion functor is strongly continuous.
-Note that for any functor we have a strongly continuous fully faithful embedding To conclude, we simply observe that is equivalent to
+such that the inclusion functor is strongly continuous. Note that for any functor we have a strongly continuous fully faithful embedding To conclude, we simply observe that is equivalent to
 and we have a factorization
-where the latter functor is fully faithful.
-This proves the theorem. ∎
+where the latter functor is fully faithful. This proves the theorem. ∎
 1.16. Example: fpqc descent
-To give a feeling of dualizable inverse limits, we prove the following statement. Here we denote by the category of commutatie rings.
-Proposition 1.93.
-1) The functor is a sheaf in the descendable topology.
-2) The functor is a sheaf in the fpqc topology.
+To give a feeling of dualizable inverse limits, we prove two statements about descent for functors with values in We denote by the -category of -ring spectra (without connectivity assumptions).
 Recall that the descendable topology on is given by the pretopology in which the covers are the maps such that is a descendable -algebra, see [Mat16, Definition 3.18].
 Namely, this means that the unit object is contained in the stable idempotent-complete ideal of generated by
-We need the following lemma.
+Proposition 1.92.
+The functor is a sheaf in the descendable topology.
+This can be deduced from [Ram24b, Corollary 4.85], but we would like to give an explicit proof using the description of the limit from Theorem 1.91. We need the following lemma.
 Lemma 1.93.
-Let be a strongly functor between presentable stable categories, and denote by its right adjoint.
+Let be a strongly continuous functor between presentable stable categories, and denote by its right adjoint.
 Suppose that there exists such that each object of is a direct summand of an object with a finite filtration with at most non-zero subquotients, which are contained in Then for any regular cardinal the functor preserves and reflects -compact objects.
 Proof.
-Since is continuous, takes to
-If is an object such that is -compact, then for any collection of objects we have
+Since is continuous, takes to If is an object such that is -compact, then for any collection of objects we have
 The assumption on can be reformulated as follows: for any infinite set the category is generated by the full subcategory as a stable idempotent-complete subcategory. We conclude that is -compact. ∎
 Proof of Proposition 1.92.
-1) Let be a map such that is descendable over Below all the tensor products over are always derived, including the tensor powers of i.e.
-they are taken in By [Mat16, Proposition 3.22] we have an equivalence By Lemma 1.94, it induces an equivalence By Theorem 1.92, we obtain
-It suffices to prove the following statement:
-the functor is fully faithful.
-To prove recall that by Proposition 1.65 for the complex of morphisms between and in the Calkin category is given by
+Let be a map such that is descendable over By [Mat16, Proposition 3.22] we have an equivalence By Lemma 1.93, it induces an equivalence By Theorem 1.91, we obtain
+It suffices to prove the following statement: the functor is fully faithful. To prove recall that by Proposition 1.64 for the spectrum of morphisms between and in the Calkin category is given by
 Since we know the descent for it suffices to show that the map
-is an isomorphism. This is equivalent to proving the following
-the map is an isomorphism.
-This follows easily from the descendability of over Indeed, by [Mat16, Proposition 3.20] the inverse system is pro-constant, hence we have equivalences
-2) Descent for finite disjoint unions is obvious.
-We need to prove that for a faithfully flat map we have As in the proof of 1), we only need to show that and to prove the fully faithfulness statement above.
+is an isomorphism.
+This is equivalent to proving the following: the map is an isomorphism.
+This follows easily from the descendability of over Indeed, by [Mat16, Proposition 3.20] the inverse sequence is pro-constant, hence we have equivalences
+This proves the proposition. ∎
+Remark 1.94.
+It seems plausible that for any descendable map and for any uncountable regular cardinal we have an equivalence
+However, we only checked the fully faithfulness statement (the proof for is the same). It is not clear how to prove the essential surjectivity even for
+Denote by the category of (discrete) commutatie rings.
+Proposition 1.95.
+The functor is a sheaf in the fpqc topology.
+Proof.
+Descent for finite disjoint unions is obvious.
+We need to prove that for a faithfully flat map we have As in the proof of Proposition 1.92, we only need to show that and to prove the fully faithfulness statement above.
 By Proposition 1.96 below, is isomorphic to an -directed colimit where are descendable faithfully flat -algebras (and is an -directed poset).
 We know that the functors and commute with -filtered colimits (by Proposition 1.73).
-Also, the functor commutes with -filtered colimits, since the category is countable.
-We deduce from part 1) that and the functor is fully faithful.
-As in the proof of 1), this implies that ∎
+Also, the functor
+commutes with -filtered colimits, since the category is countable and the category is compactly generated.
+We deduce from the proof of Proposition 1.92 that and the functor is fully faithful.
+This implies that as required.
+∎
 The following result is due to Peter Scholze (private communication). We give a short proof based on Positselski’s results on accessibility [Pos23].
@@
 Proof.
-The second assertion follows from the fact that a faithfully flat countably presented -algebra is descendable by [Mat16, Corollary 3.33].
+The last assertion follows from the fact that a faithfully flat countably presented -algebra is descendable by [Mat16, Corollary 3.33].
 By Lazard’s theorem, the category of flat -modules is equivalent to the ind-completion where is the category of finite projective -modules.
@@
 It remains to observe that if is a map of flat -algebras such that is faithfully flat over then is also faithfully flat over ∎
-Remark 1.96.
-It seems plausible that for any descendable map and for any uncountable regular we have an equivalence
-However, we only checked the fully faithfulness statement (the proof for is the same). It is not clear how to prove the essential surjectivity even for
 Recall the following -categorical notion.
@@
 By [Pos23, Theorem 4.1] the category is -accessible, and preserves and reflects -compactness.
-Consider as a natural transformation Then the category is equivalent to the joint equifier of the pairs and Applying [Pos23, Theorem 3.1], we conclude that the category is -accessible, and the forgetful functor to preserves and reflects -compactness.
+Consider as a natural transformation Then the full subcategory is exactly the joint equifier of the pairs and Applying [Pos23, Theorem 3.1], we conclude that the category is -accessible, and the forgetful functor to preserves and reflects -compactness.
 ∎
@@
 Given ind-objects we can consider each as an object of We denote by the product of these objects in We denote by the resulting functor.
-This is simply the right adjoint to the natural functor We give two more equivalent descriptions of the functor
+This is simply the right adjoint to the natural functor We give another two equivalent descriptions of the functor
 Proposition 1.99.
@@
 Let be a collection of small stable categories, and take some ind-objects Then the natural map
-is an equivalence.
+is an isomorphism.
 Proof.
@@
 Proof.
-Put
-1) Take some families of objects Since are homological epimorphisms, using Proposition 1.101, we obtain equivalences
+Put 1) Take some families of objects Since are homological epimorphisms, using Proposition 1.101, we obtain equivalences
 Hence, is a homological epimorphism.
@@
 Proposition 1.103.
-Let be a collection of dualizable categories.
-1) We have a short exact sequence
+Let be a collection of dualizable categories. 1) We have a short exact sequence
 2) If the categories are compactly generated, then we have an equivalence
@@
 Proof.
-By Theorem 1.92, we have an equivalence
-Now part 1) follows from Proposition 1.102, applied to the homological epimorphisms
-Part 2) follows from the fact that an infinite product of short exact sequences in is a short exact sequence, applied to the sequences
 Now we want to prove that an infinite product of short exact sequences in is again a short exact sequence.
@@
 The image of contains all objects of the form where and is the evaluation functor for Moreover, the image of is generated by the objects of this form.
-2) Let be a family of strongly continuous functors between dualizable categories. Then the following square commutes:
-Here the functors and are as in 1). The functor (resp. ) is (resp. ).
+2) Let be a family of strongly continuous functors between dualizable categories.
+Then the following square commutes:
+Here the functors and are as in 1). The functor (resp. ) is (resp. the right adjoint to ).
 Proof.
 1) The functor is obtained by applying to the inclusion in the short exact sequence (1.9).
-Note that the (fully faithful) functor takes to the functor The assertion now follows from Proposition 1.102, part 2), applied to the homologicl epimorphisms
-2) We need to show that the image of is contained in the image of By part 1), it suffices to show that for any the object is contained in the image of We have
+Note that the (fully faithful) functor takes to the functor The assertion now follows from Proposition 1.102, part 2), applied to the homological epimorphisms 2) We need to show that the image of is contained in the image of By part 1), it suffices to show that for any the object is contained in the image of We have
 By part 1), the latter object is in the image of This proves the proposition.∎
@@
 The only statement that requires a proof is that the functor is a quotient functor.
-Equivalently, it suffices to show that the functor is a quotient functor, i.e.
-its right adjoint is fully faithful.
+Equivalently, we need to show that the dual functor is fully faithful.
 By Lemma 1.104, the latter functor is the upper horizontal arrow in a commutative square
@@
 Since the functor commutes both with products and with filtered colimits, we deduce that (1.10) is an equivalence when all are compactly generated.
-By Propositions 1.68 and 1.105, filtered colimits and products of short exact sequences in are short exact. Consider the map of short exact sequences
+By Propositions 1.67 and 1.105, filtered colimits and products of short exact sequences in are short exact.
+Consider the map of short exact sequences
 By the above discussion, the middle vertical and the right vertical arrows are equivalences. Hence, so is the left vertical arrow. This proves (AB6) for ∎
 2. Dualizability is equivalent to flatness
-In this section we give a non-trivial criterion of dualizability of a presentable stable category via a “flatness” condition. It can be skipped on the first reading.
+In this section we give a non-trivial criterion of dualizability of a presentable stable category via a “flatness” condition.
+It can be skipped on the first reading.
 Let be a presentable stable category.
@@
 Definition 2.1.
-A category is called flat if the answer to the above question is affirmative, i.e. the tensor product preserves fully faithful functors (=monomorphisms).
-We have the following surprising result.
+1) A category is called flat in if the answer to the above question is affirmative, i.e.
+the functor preserves fully faithful functors (=monomorphisms).
+2) Let be a regular cardinal.
+A -presentable stable category is called flat in if the functor preserves fully faithful functors.
+Note that the flatness in is a non-trivial condition only if is uncountable. We have the following surprising result.
 Theorem 2.2.
-Let be a presentable stable category. Then is flat if and only if it is dualizable.
-The “if” direction is easy: dualizability of implies an equivalence To prove the “only if” direction, we need a certain “extra -functoriality” statement.
-Denote by the -category of presentable stable categories and exact accessible functors. We have the following result.
+Let be an uncountable regular cardinal.
+Then a -presentable stable category is flat in if and only if satisfies (AB6) for -small products.
+In particular, a presentable stable category is flat in if and only if it is dualizable.
+The “if” direction in the second statement is easy: dualizability of implies an equivalence The “if” direction in the first statement is non-trivial, see below for the proof.
+Our argument for the “only if” direction will use a certain “extra -functoriality” statement.
+After the previous version of the paper was published, German Stefanich found a different argument, which in fact works in a much more general relative context (and for not necessarily stable categories).
+We will use the model for -categories given by complete Segal space objects of the -category of (possibly very large) -categories.
+We refer to [Lur09b, Section 1] for a detailed discussion.
+More precisely, we consider -categories as cartesian fibrations of -categories satisfying the following conditions.
+Denote by the fiber of over Then for the natural functor
+(the -fold fiber product) is an equivalence of -categories (the Segal condition).
+The -category is a discrete groupoid.
+Given -categories an oplax -functor is a functor between fibered categories over If moreover is cartesian (i.e.
+preserves cartesian maps), then is an actual -functor.
+From now on, we fix an uncountable regular cardinal Denote by the -category of -presentable stable categories and exact -accessible functors.
+It is convenient to identify with the following (very large) cartesian fibration over Given -presentable stable categories we denote by the category of cartesian fibrations together with equivalences such that the transition functors are exact -accessible functors.
+Then the fiber over is given by the disjoint union of categories over all -tuples Given objects in and in the fiber of the mapping space over an order-preserving map is non-empty only if for and in this case it is the space of functors over which are compatible with the equivalences We denote by the locally full sub -category which is given by the full subcategories formed by cartesian fibrations with continuous transition functors.
+We have the following result.
 Proposition 2.3.
-Let be a presentable stable category.
-Then we have a natural oplax -functor which induces the usual Lurie tensor product when we restrict to continuous functors.
-Moreover, if is flat, then we get an actual -functor
+Let be a -presentable stable category. Then we have a natural oplax -functor which induces the usual Lurie tensor product Moreover, if is flat in then we have an actual -functor
+To prove this result, we need the following statement about semi-orthogonal decompositions in
+Lemma 2.4.
+Let be -presentable stable categories, and suppose that we have fully faithful -strongly continuous functors which give a semi-orthogonal decomposition
+Then the functors are also fully faithful, and they give a semi-orthogonal decomposition
 Proof.
-Givem presentable stable categories and let us denote by the -category of triples where and are fully faithful continuous functors, such that we have a semi-orthogonal decomposition
-Then is in fact an -category, and it is naturally equivalent to the category of exact accessible functors Namely, given a triple the corresponding functor is given by where is the right adjoint to Conversely, to an exact accessible functor we associate the category as in Definition 1.1.
-Recall that is the category of triples where and The inclusions and are given respectively by and
-Given correspondences and we can define their composition as follows.
-We form the pushout (in ), and take the full subcategory generated by the images of the compositions and Then is naturally an object of Moreover, this composition operation is compatible with the composition of the corresponding accessible functors.
-Now, we observe that for any presentable and we have a well-defined functor Indeed, given we have Here the fully faithfulness of (resp.
-) follows from the existence of a continuous right adjoint to (resp.
-a left adjoint to ).
-Now, given and as above, we observe that we have a natural equivalence Thus, we have a natural transformation
-in Moreover, if is flat, then (2.1) is an equivalence.
-Identifying the correspondences with accessible functors, we obtain an oplax -functor If is flat, we get an actual -functor.
-Finally, if the triple corresponds to a continuous functor then the functor is continuous.
-Hence, the functor is the right adjoint to and the object corresponds to the functor – the usual Lurie tensor product.∎
-Proof of Theorem 2.2..
-We already observed that dualizability implies flatness by identifying with
-Now, suppose that is flat. Using Proposition 2.3 (extra -functoriality), we will deduce that (AB6) holds in By Proposition 1.54 this will imply that is dualizable.
-Consider a family of filtered categories We define the following presentable (in fact, compactly generated) stable categories:
-We have the following natural exact accessible functors:
-- the product of the colimit functors
-- the product functor;
-- taking to the functor
-- the colimit functor.
-Since the axiom (AB6) holds in the category the natural transformation is an equivalence. So we have a commutative square
-Applying the extra 2-functoriality of we obtain a commutative square
-But it is straightforward to check that the latter square is of the form
-Moreover, the functors in this commutative square are analogous to the functors with “” replaced by “”. Therefore, commutativity of (2.2) exactly means that (AB6) holds in ∎
+It suffices to prove this for Note that the functor has a left adjoint and the functor has a continuous right adjoint Hence, the functors and are fully faithful.
+Their images generate by colimits.
+Finally, to see the the semi-orthogonality, note that is the left adjoint to and we have
+This is equivalent to the required semi-orthogonality. ∎
+Proof of Proposition 2.3..
+Given -presentable stable categories and let us denote by the -category of triples where is a -presentable stable category and are fully faithful -strongly continuous functors, such that we have a semi-orthogonal decomposition
+Then the category is naturally equivalent to the category of exact -accessible functors Namely, given a triple the corresponding functor is given by where is the right adjoint to Conversely, to an exact accessible functor we associate the category as in Definition 1.1.
+Recall that is the category of triples where and The inclusions and are given respectively by and We now formally define the -category as a (very large) cartesian fibration over satisfying the Segal condition, such that the fiber over is the (very large) set of -presentable stable categories.
+Given -presentable stable categories where we define the -category to be the category of -tuples where is a -presentable stable category and are fully faithful -strongly continuous functors such that the following conditions hold:
+the functors give a semi-orthogonal decomposition
+if we denote by the right adjoint to then we have isomorphisms
+If we define to be a single point, considered as an -category.
+We think of this point as a pair The condition (2.1) (which is non-trivial only for ) can be reformulated as follows (assuming that we have the required semi-orthogonal decomposition).
+Denote by the full stable subcategory generated by and Then (2.1) holds if and only if we have an equivalence
+The fiber over is given by the disjoint union of categories over all -tuples The fiber of the mapping space
+over an order-preserving map map is the following.
+It is non-empty only if for and in this case it is the space of functors together with isomorphisms It is easy to see that is a cartesian fibration, and it follows from (2.2) that this fibration satisfies the Segal condition.
+Now, we claim that we have a natural equivalence of -categories which induces the identity map on objects.
+The functor sends to the following cartesian fibration The fiber over is given by and for the mapping space is given by
+The inverse functor sends a cartesian fibration to the category of sections of the dual cocartesian fibration (with the same fibers and the same transition functors).
+Here for the functor is the composition of with the left adjoint to the evaluation at It is easy to see that the functors can be naturally extended to a functor between the cartesian fibrations over which preserves cartesian maps.
+Thus, we get an equivalence of -categories, as stated.
+We claim that for a -presentable stable category we have a natural functor between the cartesian fibrations over which does not necessarily preserve cartesian maps.
+On it is given by Given a sequence we have a functor
+given by
+Note that has the required semi-orthogonal decomposition by Lemma 2.4.
+Also, by loc.
+cit.
+the functors are fully faithful.
+Since satisfies the condition (2.2) and since the functor commutes with pushouts, we see that is indeed an object of We obtain an oplax -functor Using the above equivalence we obtain an oplax -functor We claim that it induces the usual Lurie tensor product Indeed, the cartesian fibration (corresponding to ) is the full subfibration of which consists of such that each functor is strongly continuous.
+Thus, the oplax -functor induces an actual -functor which corresponds to the usual Lurie tensor product.
+Finally, we need to show that if is flat in then we get an actual -functor We need an explicit description of the composition of -morphisms in Given correspondences and their composition can be obtained as follows.
+We form the pushout
+and take the full subcategory generated by the images of the compositions and Then is naturally an object of and it is the composition Now, the flatness of implies that the functor
+is fully faithful. Hence, the map
+in is an isomorphism. Therefore, is an actual -functor, as required. ∎
+Note that the oplax -functor can be described on -morphisms more explicitly.
+For a -presentable stable category we identify with the category of -left exact functors (these are functors which commutes with -small limits).
+Then for an exact -accessible functor the functor is identified with the composition
+where is the left adjoint to the inclusion. However, with this description it is not clear how to show that flatness of in implies that we get an actual -functor.
+Proof of Theorem 2.2.
+By Proposition 1.53, the second assertion follows from the first one.
+As above we fix an uncountable regular cardinal and consider a -presentable stable category For brevity we say that satisfies (AB6)κ if (AB6) for -small products holds in First suppose that satisfies (AB6) Take some fully faithful -strongly continuous functor between -presentable stable categories.
+The right adjoint to is identified with the precomposition functor
+We need to show that this is a quotient functor.
+By the proof of Proposition 1.53, our assumption on means that the colimit functor commutes with -small limits.
+Its right adjoint commutes with all limits, hence the functor (2.3) is a retract of the functor
+The latter is a quotient functor since the category is dualizable, hence flat.
+The class of accessible quotient functors is closed under retracts, hence (2.3) is a quotient functor, as required.
+Now, suppose that is flat in Using Proposition 2.3 (extra -functoriality), we will deduce that (AB6)κ holds in Consider a -small family of filtered categories We have the following natural exact -accessible functors between -presentable stable categories: - the product of the colimit functors - the product functor; - taking to the functor - the colimit functor.
+The functors and are continuous, and the functors and have left adjoints and respectively.
+Here is the diagonal functor , and the components of are the left Kan extension functors Tautologically, we have a commutative square
+Now, (AB6)κ for means that the square (2.4) satisfies the dual Beck-Chevalley condition: the induced map is an isomorphism. Note that the square (2.4) is equivalent to the following tensor product:
+Since the axiom (AB6) holds in the category the map is an isomorphism.
+Applying the -functoriality of given by Proposition 2.3 (since is flat in ), we conclude that the map is an isomorphism.
+Therefore, (AB6)κ holds in as required.
+∎
 3. Extensions of compactly generated categories
@@
 Is it true that the category is automatically dualizable?
-It turns out that in general the answer is “no”. To give a precise criterion of dualizability of we introduce some notation.
-Denote by the right adjoint to and by the right adjoint to Note that we have a semi-orthogonal decomposition Putting we obtain an equivalence
-Note that we have equivalences of categories of functors:
-We denote by the functor corresponding to
-Recall the following full subcategories of formed by (elementary) Tate objects.
+It turns out that in general the answer is “no”.
+To give a precise criterion of dualizability of we introduce some notation.
+Denote by the right adjoint to and by the right adjoint to Note that we have a semi-orthogonal decomposition Putting we obtain an equivalence Note that we have equivalences of categories of functors:
+We denote by the functor corresponding to Recall the following full subcategories of formed by (elementary) Tate objects.
 Definition 3.2.
@@
 the image of is contained in
-Assuming Proposition 3.3, we deduce the following result.
+Assuming Proposition 3.3, we deduce the following result, which first appeared in [CDH+20, Appendix A].
 Theorem 3.4.
@@
 Remark 3.5.
-The special case of Theorem 3.4 for split extensions is well-known.
+The special case of Theorem 3.4 for split extensions is well known.
 Namely, short exact sequences of the form (3.2) where both functors have right (resp.
@@
 Clearly, .
-If is dualizable, then for any object there exists an object such that . Such objects together with the objects of form a generating collection of compact objects of hence is compactly generated.
-. We use the above observation: is compactly generated if and only if for any object there exists a compact object such that Via the identification the object must be of the form
-We need to describe the necessary and sufficient conditions for to be a compact object of Note that commutes with colimits since is a compact object of We have
+.
+If is dualizable, then by Proposition 1.18 for any object there exists an object such that .
+Such objects together with the objects of form a generating collection of compact objects of hence is compactly generated.
+.
+We use the above observation: is compactly generated if and only if for any object there exists a compact object such that Via the identification the object must be of the form We need to describe the necessary and sufficient conditions for to be a compact object of Note that commutes with colimits since is a compact object of We have
 We deduce that the object is compact if and only if the right-hand side in (3.3) commutes with coproducts.
-Note that the functor is exactly the object where corresponds to as above. We conclude that the existence of such that is equivalent to the existence of objects and and an exact triangle
+Note that the functor is exactly the object where corresponds to as above.
+We conclude that the existence of such that is equivalent to the existence of objects and and an exact triangle
 By [Hen17, Theorem 3], this is equivalent to the inclusion By Thomason’s theorem the latter condition is equivalent to the inclusion This proves the equivalence of (ii) and (iii). ∎
@@
 The associated -linear functor sends to where is considered as a Tate object in the usual way.
-We give some natural examples of non-dualizable categories which are extensions of compactly generated categories.
+We give a natural “linear-algebraic” example of a non-dualizable category which is an extension of compactly generated categories.
 We use the notation for the direct sum of copies of an object indexed by the elements of a set
@@
 Proof.
-Consider the functor corresponding to By Proposition 3.3, we need to show that
-Identifying with the opposite of the category of -linear accessible functors we see that corresponds to the functor Using the fact that (AB6) holds in we obtain isomorphisms
-Putting we obtain an isomorphism
-To show that we need to show that the image of in is not a pro-constant object.
+Consider the functor corresponding to By Proposition 3.3, we need to show that Identifying with the opposite of the category of -linear accessible functors we see that corresponds to the functor Using the fact that (AB6) holds in we obtain isomorphisms
+Putting we obtain an isomorphism To show that we need to show that the image of in is not a pro-constant object.
 Assume the contrary.
@@
 Let is a field. Consider the following full subcategory of the derived category of quasi-coherent sheaves on the affine plane:
-Then Theorem 3.10 implies that is not dualizable. Indeed, the corresponding subset contains the ideals and but does not contain the ideal
+Then Theorem 3.10 implies that is not dualizable. Indeed, the corresponding subset of is given by We see that contains the ideals and but does not contain the ideal hence is not convex.
 We need the following auxiliary statement.
@@
 Since is injective, we obtain the map This map does not factor through a finite direct sum, which contradicts the commutation of with coproducts.
-Next, we show that each -module is finitely generated. We may and will assume that Suppose that is not finitely generated. Then we can construct a strictly increasing sequence of submodules
+Next, we show that each -module is finitely generated.
+We may and will assume that Suppose that is not finitely generated.
+Then we can construct a strictly increasing sequence of submodules
 such that each is finitely generated.
@@
 This contradicts the commutation of with coproducts.
-So we showed that has bounded finitely generated cohomology. To prove that is perfect, it suffices to observe that is perfect over by adjunction, the functor commutes with coproducts. ∎
+So we showed that has bounded finitely generated cohomology.
+To prove that is perfect, it suffices to observe that is perfect over by adjunction, the functor commutes with coproducts.
+∎
 Lemma 3.13.
@@
 where and are strongly continuous, and the categories on the left and on the right are compactly generated.
-Suppose that is a dualizable category. The proof of Proposition 3.3 shows that there exists an object and a map such that the functor
+Suppose that is a dualizable category.
+The proof of Proposition 3.3 shows that there exists an object and a map such that the functor
 commutes with coproducts. Since denoting by the inclusion functor, we see that the functor commutes with coproducts. By Lemma 3.12, we deduce that But choosing some we have
@@
 Suppose that the category is dualizable.
-Choose some such that and Since is noetherian, we may and will assume that is a minimal element of the set Denote by and respectively the prime ideals corresponding to the points and Put We identify the set with its image in and put
-Since the category is dualizable, so is the category Moreover, since is dualizable, the functor
+Choose some such that and Since is noetherian, we may and will assume that is a minimal element of the set Denote by and respectively the prime ideals corresponding to the points and Put We identify the set with its image in and put Since the category is dualizable, so is the category Moreover, since is dualizable, the functor
 is fully fiathful.
@@
 The “only if” part of Theorem 3.10 fails for non-noetherian rings. Namely, let be a field, and consider the rank valuation ring
-Then where is the maximal ideal, and the field of fractions of is
-Consider the non-convex subset Then we have
-– a compactly generated category.
-To see this, it suffices to observe that if is an -torsion -module, then acts by zero on Indeed, if is an element such that then It follows that for we have
+Then where is the maximal ideal, and the field of fractions of is Consider the non-convex subset Then we have
+– a compactly generated category. To see this, it suffices to observe that if is an -torsion -module, then acts by zero on Indeed, if is an element such that then It follows that for we have
 Hence, the full subcategories and in are mutually orthogonal.
@@
 To avoid set-theoretic issues, we will consider only accessible localizing invariants with values in accessible stable categories.
-Let be an accessible stable category with (not necessarily cocomplete). Recall that a functor is a localizing invariant if the following conditions hold:
+Let be an accessible stable category (not necessarily cocomplete).
+Recall that a functor is a localizing invariant if the following conditions hold:
 for any fiber-cofiber sequence of the form
@@
 If is a regular cardinal such that has -filtered colimits, we denote by the category of localizing invariants which commute with -filtered colimits.
-It follows from [BGT13] that for any regular cardinal there exists a universal localizing invariant such that the category is accessible and has -filtered colimits, and for any other such category we have
+It follows from [BGT13] that for any regular cardinal there exists a universal localizing invariant such that the category is accessible and has -filtered colimits, the functor is -continuous and for any other such category we have
 Here the right-hand side is the category of exact functors which commute with -filtered colimits.
@@
 Lemma 4.2.
-Any short exact sequence in is an -filtered colimit of short exact sequenes of -compact categories.
+Let be an uncountable regular cardinal. Any short exact sequence in is a -filtered colimit of short exact sequences of -compact categories.
 Proof.
 Consider a short exact sequence
-in Choose an -directed system such that each is in and Consider a poset of pairs where and is a full stable idempotent-complete subcategory generated by at most countable collection of objects.
-By 1.6, such a category is automatically -compact.
+in Choose a -directed system such that each is in and Consider a poset of pairs where and is a full stable idempotent-complete subcategory generated by a -small collection of objects.
+By Proposition 1.6, such a category is automatically -compact.
 The poset is -directed, and we conclude that (4.1) is a -indexed colimit of short exact sequences
@@
 Localizing invariants which commute with filtered colimits are called finitary. In this case we write
-Recall that for each regular the category has a natural (closed) symmetric monoidal structure, and the functor is symmetric monoidal.
+Recall that for each regular cardinal the category has a natural symmetric monoidal structure, and the functor is symmetric monoidal.
 Below we will often formulate results in terms of the universal localizing invariants.
@@
 In the forthcoming papers [E1, E2] we will study the categories and in great detail.
-In particular, we will prove that the category and its relative versions are rigid as a symmetric monoidal categories in the sense of Gaitsgory and Rozenblyum [GaRo17, Definition 9.1.2].
+In particular, we will prove that the category and its relative versions are rigid as a symmetric monoidal categories in the sense of Gaitsgory and Rozenblyum [GaiRoz17, Definition 9.1.2].
 This in particular will imply that the category is dualizable, which is quite non-obvious from the definition.
@@
 4.2. Localizing invariants of dualizable categories
-The notion of a localizing invariant is defined in the same way as for By Theorem C.6 the category is -presentable, and again we will consider only accessible localizing invariants.
-It turns out that any localizing invariant defined on small categories has a unique extension to a localizing invariant defined on dualizable categories, so that
-Roughly the idea is the following.
+The notion of a localizing invariant is defined in the same way as for By Ramzi’s theorem [Ram24a] (see also Theorem C.6 below), the category is -presentable, and again we will consider only accessible localizing invariants.
+It turns out that any localizing invariant defined on small categories has a unique extension to a localizing invariant defined on dualizable categories, so that Roughly the idea is the following.
 Proposition 1.61 shows that for a compactly generated presentable category the Calkin category is, up to direct summands, the quotient of by the full subcategory of compact objects For a general dualizable category one can think of as a “virtual quotient of by compact objects”.
@@
 Proof.
-By Propositions 4.6 and 4.8, the assignment is indeed a functor By Proposition 4.7, it is a right inverse to the precomposition
-On the other hand, for any and any dualizable category we have a short exact sequence
+By Propositions 4.6 and 4.8, the assignment is indeed a functor By Proposition 4.7, it is a right inverse to the precomposition On the other hand, for any and any dualizable category we have a short exact sequence
 which is functorial in This gives an isomorphism of functors This proves the theorem.∎
-Further, we have two straightforward results about commutation of with sufficiently nice pullbacks and pushouts.
+Further, we have two straightforward results about commutation of localizing invariants with sufficiently nice pullbacks and pushouts in
 Proposition 4.11.
@@
 By Proposition 1.74, the functor is fully faithful. Clearly, the functor between the quotients is an equivalence. Applying we obtain the commutation of with the pushout.∎
-We observe that there is an alternative equivalent definition of as follows.
+We observe that there is an alternative equivalent way to extend localizing invariants to dualizable categories.
 Proposition 4.13.
@@
 Theorem 4.16.
-Let be an accessible localizing invariant, where is a presentable stable category. Then the functor is the right Kan extension of via
-We will again use the fact that the category is -presentable (Theorem C.6).
-It is convenient to define the following topologies on the opposite categories and In both cases we define a pretopology for which a family of functors is a cover if consists of one element and the functor is fully faithful.
+Let be an accessible localizing invariant, where is an accessible stable category. Then the functor is the right Kan extension of via
+We will again use the fact that the category is -presentable (Ramzi’s theorem), and also one of the characterizations of -compact objects in (for ), given by Theorem C.6.
+It is convenient to define the following topologies on the opposite categories and In both cases we define a pretopology for which a family of functors is a cover of if consists of one element and the functor is fully faithful.
 To deal with set-theoretic issues we also consider the induced pretopologies on the small subcategories and (note that these subcategories are closed under finite limits), where is an uncountable regular cardinal.
 In each case we denote by the corresponding Grothendieck topology.
+Note that in each case is a subcanonical topology.
+Although the category is not necessarily complete (it is only assumed to be accessible and stable), we still have a notion of an -valued sheaf.
+Namely, the sheaf condition only requires the existence of the relevant limits, not all limits.
 We make the following observation.
@@
 Consider the Čech conerve of i.e.
-– the cofiber product of copies of It follows from Proposition 4.12 that is the Čech conerve of Since is stable, we conclude that ∎
-Next, we observe the following equivalence of categories of sheaves.
-Proposition 4.18.
-Let be an uncountable regular cardinal, and let be a presentable category. Let be a -sheaf. Then the composition is also a -sheaf. The induced functor
-is an equivalence. Its inverse is given by the right Kan extension.
-Proof.
-The first statement is obvious.
-The functor sends covers to covers, and it commutes with finite limits.
-It remains to apply Theorem C.6, part 3), condition (iv): it implies that for any fully faithful strongly continuous functor between -compact dualizable categories we can find a fully faithful strongly continuous functor where is -compact.
-In particular, the covering of is a refinement of the covering This implies the equivalence of categories of sheaves, as stated.∎
+– the cofiber product of copies of It follows from Proposition 4.12 that is the Čech conerve of Since is stable, we conclude that More precisely, for any morphism in its Čech conerve has the following property: for the map is an isomorphism.
+∎
 Proof of Theorem 4.16..
-Choose an uncountable regular cardinal such that and commute with -filtered colimits.
-It suffices to show that the restriction of to is the right Kan extension of the restriction of to This follows immediately from Propositions 4.17 and 4.18: and are -sheaves respectively on and and is the composition ∎
+Choose an uncountable regular cardinal such that is -accessible, and and commute with -filtered colimits.
+It suffices to show that the restriction of to is the right Kan extension of the restriction of to Let be a -compact dualizable category.
+By Theorem C.6, we can find a fully faithful strongly continuous functor such that is compactly generated and is -compact in Denote by the Čech conerve of in Then each is a compactly generated category, and is -compact.
+Now, the functor is a sheaf of spaces, covered by the representable sheaf with the Čech nerve By Proposition 4.17, the functors and are sheaves with values in It follows that the right Kan extension of from to exists, and its value on is given by
 4.4. Example: sheaves on the real line
-We consider the special case of sheaves on the real line.
+As an example, we will compute the finitary localizing invariants of the category of sheaves on the real line.
 Let be a dualizable category.
 Recall that the category is dualizable [Lur18].
-In this subsection we give a detailed proof of the following special case of Theorem 0.2.
+We will give a detailed proof of the following special case of Theorem 0.2.
 Proposition 4.18.
@@
 The object represents the class
-Recall that if is compactly generated, then we have a notion of a singular support for -valued sheaves on -manifolds [RS18, Definition 4.5]. We will be interested in the following special case.
-Denote by the cone of non-positive integers.
+Recall that if is compactly generated, then we have a notion of a singular support for -valued sheaves on -manifolds [RS18, Definition 4.5].
+We will be interested in the following special case.
+Denote by the cone of non-positive real numbers.
 The -topology on is defined as follows: a subset is -open if is open and In other words, is -open if either or or for some We denote by the set equipped with -topology.
@@
 This is a straightforward generalization of the special case of [KS18, Theorem 2.5].∎
-In fact there is a reasonable notion of singular support for sheaves with values in arbitrary presentable stable categories, see Remarks 4.24 and 4.25.
-To avoid the technicalities, we will use the following definition in the special case
-Notation. For a presentable stable category we denote by resp. the category resp. of -valued sheaves on resp.
+In fact there is a reasonable notion of singular support for sheaves with values in arbitrary presentable stable categories, see Remarks 4.23 and 4.24.
+To avoid the technicalities, we will use the following definition in the special case Notation.
+For a presentable stable category we denote by resp.
+the category resp.
+of -valued sheaves on resp.
 Note that for we have a tautological equivalence where the latter category was defined in Subsection 1.6.
@@
 Proposition 4.21.
-Let be a dualizable category, and choose some dense subset Denote by the set considered as a linearly ordered set with the usual order.
-1) We have a short exact sequence in
-Here the functor is given by
-2) We have
+Let be a dualizable category, and choose some dense subset Denote by the set considered as a linearly ordered set with the usual order. 1) We have a short exact sequence in
+Here the functor is given by 2) We have
 Proof of Proposition 4.21.
@@
 Now, we observe that the category is naturally equivalent to the right orthogonal This equivalence sends a sheaf to the functor Note that for we have hence
-Using (4.6), we see that the category consists of functors such that for any the map is an isomorphism. This corresponds exactly to the fact that is a sheaf on
-Finally, since the functor has a right adjoint, we have equivalences
+Using (4.6), we see that the category consists of functors such that for any the map is an isomorphism.
+This corresponds exactly to the fact that is a sheaf on Finally, since the functor has a right adjoint, we have equivalences
 This proves that we have a short exact sequence (4.4) in This proves part 1).
-2) By (4.3), it is sufficient to prove the vanishing for Take and consider the short exact sequence (4.4). We need to show that the map is an equivalence in
-Note that the categories and are compactly generated.
+2) By (4.3), it is sufficient to prove the vanishing for Take and consider the short exact sequence (4.4).
+We need to show that the map is an equivalence in Note that the categories and are compactly generated.
 We have an -indexed semi-orthogonal decomposition where the (fully faithful, strongly continuous) functors are defined in (4.5).
@@
 Again, it suffices to prove the result for We have a short exact sequence in
-where the is the inclusion. Hence, it suffices to prove the result for sheaves on
-We claim that there is a natural short exact sequence in
+where the is the inclusion. Hence, it suffices to prove the result for sheaves on We claim that there is a natural short exact sequence in
 Assuming this, the result follows from Proposition 4.21.
-The functor in (4.7) is the left adjoint to The functor is given by where is the map (extending ) to the sobrification of Note that as a set is naturally identified with (where is the generic point), and clearly
-An explicit description of the functors and is the following. For a convex open subset and for we have
+The functor in (4.7) is the left adjoint to The functor is given by where is the map (extending ) to the sobrification of Note that as a set is naturally identified with (where is the generic point), and clearly An explicit description of the functors and is the following.
+For a convex open subset and for we have
 With these descriptions it is straightforward to check that (4.7) is indeed a short exact sequence. ∎
@@
 Remark 4.23.
-If is a -manifold and is a presentable compactly generated stable category, then the singular support of a -valued sheaf on is defined in [RS18, Definition 4.5], which is a straightforward generalization of [KS90, Definition 4.4.1].
+If is a -manifold and is a compactly generated presentable stable category, then the singular support of a -valued sheaf on is defined in [RS18, Definition 4.5], which is a straightforward generalization of [KS90, Definition 4.4.1].
 However, if is a dualizable category which is not compactly generated, then already in the -dimensional case the non-characteristic deformation lemma [RS18, Theorem 4.1] fails and the conditions of [KS90, Proposition 5.1.1] are no longer equivalent.
 To illustrate this, let Consider a -valued sheaf on given by for Then according to [RS18, Definition 4.5] the singular support would be the zero section But clearly the sheaf is not locally constant, hence the definition of the singular support has to be modified.
-We mention several equivalent ways to define the singular support for a sheaf on a -manifold with values in a presentable stable category
-1) A point is not contained in if and only if for some (any) choice of local coordinates it satisfies the equivalent conditions and of [KS90, Proposition 5.1.1].
+We mention several equivalent ways to define the singular support for a sheaf on a -manifold with values in a presentable stable category 1) A point is not contained in if and only if for some (any) choice of local coordinates it satisfies the equivalent conditions and of [KS90, Proposition 5.1.1].
 2) The singular support is the smallest conical closed subset such that is contained in the essential image of the functor
@@
 3) The singular support of is the support of the conical sheaf (of spectra) on (microlocal Hom), where the bifunctor is defined as in [KS90, Definition 4.4.1].
-4) Choose a regular cardinal such that takes values in Then the composition defines a sheaf with values in the compactly generated category (since the Yoneda embedding commutes with limits).
+4) Choose a regular cardinal such that takes values in Then the composition defines a sheaf with values in the compactly generated category (since the Yoneda embedding commutes with existing limits).
 We put where the latter singular support is defined as in [RS18, Definition 4.5].
@@
 Remark 4.25.
-We first recall a simple proof of the commutation of with infinite products.
-Proposition 4.27 (Heller’s criterion).
-[KW19, Lemma 2.12] Let be a small triangulated category, and take two objects The following are equivalent:
+We first recall the argument showing the commutation of of small triangulated categories with infinite products.
+Proposition 4.26.
+[KW19, Lemma 2.12](Heller’s criterion) Let be a small triangulated category, and take two objects The following are equivalent:
 we have in
@@
 Theorem 4.28.
-1) Let be an accessible stable category with a non-degenerate -structure. Let be accessible localizing invariants and a morphism such that is an isomorphism. Then is an isomorphism.
+1) Let be an accessible stable category with a non-degenerate -structure.
+Let be accessible localizing invariants and a morphism such that is an isomorphism.
+Then is an isomorphism.
 2) In particular, both -theory and continuous -theory commute with infinite products.
@@
 Consider a collection of dualizable categories.
-We need to show that We may assume that the categories are all equivalent: indeed, each is a retract in of the category
-In other words, we need to show that for a set and a dualizable category the map is an isomorphism.
+We need to show that We may assume that the categories are all equivalent: indeed, each is a retract in of the category In other words, we need to show that for a set and a dualizable category the map is an isomorphism.
 But this is the special case of 1) for the standard -structure on Indeed, by Proposition 1.105 the functor is a localizing invariant, and so is the functor When is compactly generated, the map is an isomorphism by Corollary 4.27.∎
@@
 Proof.
-Put The functors are compatible with the functors hence we obtain a functor By construction, we have We need to show that in
-For each the inclusions define a fully faithful functor Its left adjoint is simply the structural functor Denote by the composition functors, Since the compositions vanish for we see that the direct sum of the functors is well-defined.
-Moreover, we have
-Now, denote by the composition, and put (where ). Equivalently, is the semi-orthogonal projection onto the left orthogonal to (where ). Denote by the natural transformations, Then
-Since the composition vanishes for we see that the direct sum is also a well-defined endofunctor of By the above, we obtain an exact triangle
+Put The functors are compatible with the functors hence we obtain a functor By construction, we have We need to show that in For each the inclusions define a fully faithful functor Its left adjoint is simply the structural functor Denote by the composition functors, Since the compositions vanish for we see that the direct sum of the functors is well-defined.
+Moreover, we have Now, denote by the composition, and put (where ).
+Equivalently, is the semi-orthogonal projection onto the left orthogonal to (where ).
+Denote by the natural transformations, Then Since the composition vanishes for we see that the direct sum is also a well-defined endofunctor of By the above, we obtain an exact triangle
 We conclude that
@@
 Proof.
-Since we may assume that Applying 4.22 to we obtain a short exact sequence
+Since we may assume that Applying Proposition 4.21 to we obtain a short exact sequence
 in Moreover, the source and the target of are compactly generated categories.
@@
 It suffices to show that is a -equivalence.
-As in the proof of proposition 4.22, we put and
-Note that the functor has an obvious section which corresponds to the collection of “representable” functors where is defined in (4.5). We need to show that in
-Choose any bijection Denote by the full stable subcategory generated by the objects Then is equivalent to the category where as usual denotes the poset Clearly, each inclusion has both left and right adjoints.
-Moreover, we see that the elements form a basis of
-It follows from Corollary 4.31 that Since we deduce that
+As in the proof of proposition 4.21, we put and Note that the functor has an obvious section which corresponds to the collection of “representable” functors where is defined in (4.5).
+We need to show that in Choose any bijection Denote by the full stable subcategory generated by the objects Then is equivalent to the category where as usual denotes the poset Clearly, each inclusion has both left and right adjoints.
+Moreover, we see that the elements form a basis of It follows from Corollary 4.30 that Since we deduce that
 It remains to observe that the functor induces the identity map on it sends each object to itself, ∎
@@
 Assuming that is an isomorphism for some we see that for any the composition
-is also an isomorphism. By induction, we see that is an isomorphism for
-Denote by the induced map between the localizing invariants of dualizable categories.
+is also an isomorphism.
+By induction, we see that is an isomorphism for Denote by the induced map between the localizing invariants of dualizable categories.
 Arguing as above, we see that is an isomorphism for It remains to show that if is an isomorphism for some then is also an isomorphism.
@@
 5.2. Sheaves of spectra on continuous posets
-Let be a continuous poset. We denote by the category of sheaves on (where is considered as a category) for the above topology. This is in fact the category from Subsection 1.6.
+Let be a continuous poset.
+We denote by the category of sheaves on (where is considered as a category) for the above topology.
+This is in fact the category from Subsection 1.6.
 Namely, since every object of has the smallest covering sieve, a presheaf is a sheaf if and only if
@@
 For denote by the “representable” presheaf of spectra,
-Its sheafification is given by
 Proposition 5.3.
@@
 Consider the category of -valued presheaves on (i.e.
-the category of sections of the associated biCartesian fibration over ).
+the category of sections of the associated cocartesian fibration over which is also cartesian).
 For we denote by the functor given by
-The right adjoint to is given by
-Denote by the category of -valued sheaves on For we have a functor The right adjoint to is given by
-Our goal in this section is to prove the following result.
+The right adjoint to is given by Denote by the category of -valued sheaves on For we have a functor The right adjoint to is given by Our goal in this section is to prove the following result.
 Theorem 5.4.
-1) The category is dualizable.
-2) For each compact element the functor is strongly continuous and fully faithful.
-3) The functors induce an isomorphism
+1) The category is dualizable. 2) For each compact element the functor is strongly continuous and fully faithful. 3) The functors induce an isomorphism
 Proof.
 For a presheaf we denote by its sheafification. Since any object has the smallest covering sieve we see that
-In particular, we have for This shows that for the functor commutes with coproducts. This proves 2).
-To prove 1), we first observe that the category is dualizable. Indeed, we have a -indexed semi-othogonal decomposition of the category in
+In particular, we have for This shows that for the functor commutes with coproducts.
+This proves 2).
+To prove 1), we first observe that the category is dualizable.
+Indeed, we have a -indexed semi-othogonal decomposition of the category in
 Since are dualizable, so is the category by Proposition 1.81.
@@
 To prove 3), consider the full subcategory By the above we have a short exact sequence in
-where the functor is the left adjoint to the inclusion. We first observe the following.
-Claim 1. Let be a presheaf such that for any the map is zero. is Then is contained in
+where the functor is the left adjoint to the inclusion. We first observe the following. Claim 1. Let be a presheaf such that for any the map is zero. Then is contained in
 Proof of Claim 1..
@@
 Next, we observe that the category has the following semi-orthogonal decomposition.
-Claim 2. For the functor is zero (resp. fully faithful) for (resp. for ). Moreover, we have a semi-orthogonal decomposition
+Claim 2.
+For the functor is zero (resp.
+fully faithful) for (resp.
+for ).
+Moreover, we have a semi-orthogonal decomposition
 Proof of Claim 2..
@@
 In particular, we see that is zero for If then
-Hence, the adjunction unit is an isomorphism. We conclude that the functor is fully faithful for
-Similarly, if and then for This implies the semi-orthogonality.
+Hence, the adjunction unit is an isomorphism.
+We conclude that the functor is fully faithful for Similarly, if and then for This implies the semi-orthogonality.
 Finally, taking into account the vanishing of for we deduce that the categories generate This proves the claim.∎
 5.4. Localizing invariants of categories of cosheaves on continuous posets
-Let now be a copresheaf of dualizable categories on for By a -valued copresheaf on we mean a section of the associated biCartesian fibration over We denote by this category of copresheaves.
+Let now be a copresheaf of dualizable categories on for By a -valued copresheaf on we mean a section of the associated cocartesian fibration over (which is also cartesian).
+We denote by this category of copresheaves.
 Further, we denote by the full subcategory of cosheaves: it consists of copresheaves such that We denote by the cosheafification functor, i.e.
@@
 Theorem 5.5.
-1) The category is dualizable.
-2) For each compact element the functor takes values in cosheaves, hence we have a fully faithful strongly continuous functor
-3) The functors induce an isomorphism
+1) The category is dualizable. 2) For each compact element the functor takes values in cosheaves, hence we have a fully faithful strongly continuous functor 3) The functors induce an isomorphism
 Proof.
 As in the proof of Theorem 5.4, the category is dualizable and we have a -indexed semi-orthogonal decomposition The cosheafification functor is continuous, since we have
-It follows that the category is dualizable. This proves 1).
-Part 2) is clear from the definitions.
-We prove 3). Consider the short exact sequence
+It follows that the category is dualizable. This proves 1). Part 2) is clear from the definitions. We prove 3). Consider the short exact sequence
 where is the kernel of the cosheafification, and is the left adjoint to the inclusion. We claim that the functors are fully faithful for and we have a semi-orthogonal decomposition
@@
 Since the functors take values in cosheaves, we obtain the semi-orthogonal decomposition (5.8).
+As in the proof of Theorem 5.4, we see that after applying to the sequence (5.7) we get a split exact triangle:
+This proves 3). ∎
 6. Sheaves on locally compact Hausdorff spaces
@@
 Here we consider finite CW complexes both as topological spaces and as -groupoids (i.e.
-objects of -category of finite spaces).
+as objects of the -category of finite spaces).
 Theorem 6.1.
@@
 It follows that for any finite CW complex and for any dualizable category we have
-Therefore we have a well defined functor such that for a finite CW complex we have Since it suffices to prove that commutes with pullbacks.
-This is straightforward. Namely, consider cellular closed embeddings of finite CW complexes Then both pullback functors and are strongly continuous localizations and we have a pullback square
+Therefore we have a well defined functor such that for a finite CW complex we have Since and it suffices to prove that commutes with pullbacks.
+This is straightforward.
+Namely, consider cellular closed embeddings of finite CW complexes Then both pullback functors and are strongly continuous localizations and we have a pullback square
 By Proposition 4.11 we obtain an isomorphism
@@
 We will consider -valued (pre)sheaves on Recall that is a sheaf if and only if it satisfies the following conditions:
-for any open we have a pullback square
-for any open the map is an isomorphism.
+for any open we have a pullback square in
+for any open the map is an isomorphism in
 Recall from Example 5.2 that the poset of open subsets of is continuous.
@@
 We denote by the sheafification functor. It has a factorization:
+We call a presheaf reduced if i.e. if the condition (i) holds for Denote by the full subcategory of reduced presheaves. The inclusion functor has a left adjoint where
+Note that if is continuous, then so is
+Remark 6.2.
+In what follows, we crucially use the following property of the continuous lattice if and then In other words, the way below relation in is multiplicative in the terminology of [GHKLMS80].
 We make the following observation.
 Proposition 6.3.
-Suppose that is a presheaf which satisfies descent for finite covers, i.e. the conditions (i) and (ii). Then is a sheaf. Equivalently, the natural map is an isomorphism.
-In particular, the inclusion functor is continuous.
 Proof.
 Assuming that satisfies finite descent, it suffices to show that satisfies finite descent.
-Clearly, so (i) holds. To see that (ii) holds for we first introduce the following notation: for an open we denote by the poset Now take some open subsets and consider the maps
+Clearly, so (i) holds.
+To see that (ii) holds for we first introduce the following notation: for an open we denote by the poset Now take some open subsets and consider the maps
 The standard arguments show that both maps (6.2) and (6.3) are cofinal (hence, the maps between the same posets with the opposite order are final).
 This directly implies that (ii) holds for Thus, is indeed a sheaf.
-For the final assertion of the proposition, it suffices to notice that the class of presheaves satisfying finite descent is closed under colimits. ∎
+For the final assertion of the proposition, it suffices to notice that the class of presheaves satisfying finite descent is closed under colimits.
+∎
 We consider the category as the first approximation of the category Note that if is a presheaf with values in then we know the finitary localizing invariants of the category by Theorem 5.4.
-To construct the further approximations, it is convenient to use the language of -sheaves, which we know recall.
+To construct the further approximations, it is convenient to use the language of -sheaves, which we now recall.
 We first extend the presheaf to locally closed subsets by left Kan extension: for a locally closed we put
-where runs over open subsets such that is a closed subset of In particular, we obtain a presheaf of presentable stable categories on
-Denote by the poset of compact subsets of
+where runs over open subsets such that is a closed subset of In particular, we obtain a presheaf of presentable stable categories on Denote by the poset of compact subsets of
 Definition 6.4.
-1) Denote by the category of -presheaves, i.e. the category of -valued presheaves on
-2) We denote by the full subcategory of continuous -presheaves, i.e. -presheaves such that for any compact we have an isomorphism
-Here if contains an open neighborhood of
-3) We denote by the full subcategory formed by continuous -presheaves satisfying additional conditions:
-i)
-ii) For compact subsets the following is a pullback square:
-Note that the poset is continuous, and iff Hence, the category is the category of -valued cosheaves for the topology from Subsection 5.1.
-The following is an (almost) straightforward generalization of Lurie’s theorem for the case of a constant presheaf
+1) We denote by the category of -presheaves, i.e.
+the category of -valued presheaves on 2) We denote by the full subcategory of continuous -presheaves, i.e.
+-presheaves such that for any compact we have an isomorphism
+Here if contains an open neighborhood of 3) We denote by the full subcategory formed by continuous -presheaves satisfying additional conditions: i) ii) For compact subsets the following is a pullback square:
+Note that the poset is continuous, and iff Hence, the category is the category of -valued cosheaves on for the topology from Subsection 5.1.
+Again, it is convenient to call a -presheaf reduced if The functor is defined as in (6.1).
+If is a continuous -presheaf, then so is The following is an (almost) straightforward generalization of Lurie’s theorem for the case of a constant presheaf
 Proposition 6.5.
-[Lur09, Theorem 7.3.4.9, Corollary 7.3.4.10] 1) Consider the following functors
-They are quasi-inverse equivalences.
 Proof.
-Part 1) is proved by the same argument as [Lur09, Theorem 7.3.4.9], and in fact we do not need the strong continuity of the restriction functors
-We prove 2). Let be a -valued sheaf, and denote by the associated continuous -presheaf. Clearly, Consider compact subsets and put It suffices to show that the square
+Part 1) is proved by the same argument as in [Lur09a, Theorem 7.3.4.9], and in fact we do not need the strong continuity of the restriction functors We prove 2).
+Let be a -valued sheaf, and denote by the associated continuous -presheaf.
+We need to prove that is a -sheaf.
+Clearly, Consider compact subsets and put It suffices to show that the square
 is the directed colimit of pullback squares
-over open neighborhoods
-Clearly, we have
+over open neighborhoods Clearly, we have
 Applying Proposition 1.66 and Lemma 6.6 below, we obtain the isomorphisms
-(here we used that the restriction functors are strongly continuous). The proof for the remaining two objects of the square (6.5) is similar.
-The proof in the other direction is as in [Lur09, Theorem 7.3.4.9] (and here we do not need the strong continuity of the restriction functors). ∎
+(here we used that the restriction functors are strongly continuous).
+The proof for the remaining two objects of the square (6.6) is similar.
+Therefore, is a -sheaf.
+The proof in the other direction is as in [Lur09a, Theorem 7.3.4.9] (and here we do not need the strong continuity of the restriction functors).
+∎
 Lemma 6.6.
-Let be a directed poset, and consider a functor Consider a section of the associated biCartesian fibration over Denote by the transition functors, and denote by the functors to the colimit.
+Let be a directed poset, and consider a functor Consider a section of the associated cocartesian fibration over Denote by the transition functors, and denote by the functors to the colimit.
 Then for any we have
@@
 We consider as a poset with the reverse inclusion order.
-Consider the set of all collections of compact (=closed) subsets of of the form as in (6.7). Define the partial order on by setting if Then is a directed poset.
+Consider the set of all collections of compact (=closed) subsets of of the form as in (6.8).
+Define the partial order on by setting if Then is a directed poset.
 Remark 6.7.
-1) If is locally compact, then the poset has a natural Grothendieck pretopology, where the family is a cover if is finite and Then for each compact the poset is a cofinal family of covering sieves for
-2) If is compact and is a sieve as above, then we have a final functor
+1) If is locally compact, then the poset has a natural Grothendieck pretopology, where the family is a cover if is finite and Then for each compact the poset is a cofinal family of covering sieves for 2) If is compact and is a sieve as above, then we have a final functor Here for a set we denote by the poset of non-empty subsets of
 Proposition 6.8.
@@
 Proof.
-Consider the full subcategory of presheaves satisfying only the finite descent, i.e.
+Consider the full subcategory of -presheaves satisfying only the finite descent, i.e.
 sheaves for the finitary topology on described in Remark 6.7 1).
@@
 Proposition 6.10.
-1) For each the functor is a strongly continuous localization. The essential image of the right adjoint consists of continuous presheaves such that for each compact we have
+Suppose that the presheaf is reduced, i.e.
+1) For each the functor is a strongly continuous localization.
+The essential image of the right adjoint consists of continuous presheaves such that for each compact we have
 2) Put Then the functor is a localization, and the essential image of its right adjoint is exactly the category of -sheaves In particular, we have an equivalence
 Proof.
-Observe that the analogue of part 1) for categories holds trivially. To prove the statement 1), we only need to show the following.
-Claim. Let be a -presheaf satisfying (6.9). If each of the pullbacks is in then also
+Observe that the analogue of part 1) for categories holds trivially.
+To prove the statement 1), we only need to show the following.
+Claim.
+Suppose that Let be a -presheaf satisfying (6.10).
+If each of the pullbacks is in then also
 Proof of Claim..
-We may assume that i.e. Take a compact subset It suffices to show that the map
-is an isomorphism, and similarly for and
 2) It follows from 1) that the functor is a strongly continuous localization, and the essential image of its right adjoint consists of continuous -presheaves such that for any compact and for any we have
@@
 Proposition 6.13.
-Suppose that is compact. For a sieve the natural functor
+Suppose that is compact and For a sieve the natural functor
 is an equivalence. Moreover, we have
 Proof.
-This is basically an iterated application of Propositions 4.11 and 6.9. We prove the statement by induction on where the case is clear.
-Let where We put
+This is basically an iterated application of Propositions 4.11 and 6.10. We prove the statement for a sieve by induction on where the case is clear. Let We put
 Using the induction hypothesis, we see that the functors
-are equivalences.
+are equivalences. Applying Propositions 6.9 and 6.10, we see that the composition
+is a localization (both functors are localizations). It follows from Proposition 4.11 that the pullback square
+in is also a pullback square in and applied to this square is a pullback square in This proves the proposition. ∎
+For a topological space we denote by the set of open-closed subsets of
 Proposition 6.14.
-Let be locally compact Hausdorff, and a presentable stable category. Given a -presheaf we define another -presheaf by the formula
-where runs over open-closed subsets of We have a natural map
-1) If is continuous, then so is
-2) Assuming that is continuous, the map of -sheaves is an isomorphism.
+Let be locally compact Hausdorff space, and a presentable stable category. Given a reduced -presheaf we define another -presheaf by the formula
+We have a natural map 1) If is continuous, then so is 2) Assuming that is continuous, the map of -sheaves is an isomorphism.
 Proof.
@@
 It suffices to show that the map is an isomorphism.
-Let us observe that the weak sheafification functor is given by
+Let us observe that the weak sheafification functor on reduced -presheaves is given by
 where the colimit is over the decompositions of into a finite disjoint union of compact subsets. Now it is clear that the map is an isomorphism. ∎
@@
 We observe the following: for any compact inclusion of open subsets, for any objects and for any compact map in the induced map of sheaves is compact in Indeed, this follows from the factorization
-and the fact that the functor is strongly continuous.
-Now observe that we have a generating collection of sheaves
+and the fact that the functor is continuous. Now observe that we have a generating collection of sheaves
 For any such object since is a dualizable category, we can choose a sequence in such that and all the maps are compact.
 Choose a sequence of open subsets and compact inclusions such that and each is countable at Then
-and by the above discussion each map is compact. This shows that the category is dualizable.
+and by the above discussion each map is compact.
+This shows that the category is dualizable.
 To compute we first reduce to the case when is compact.
@@
 Let be the presheaf of dualizable categories on obtained by extension by zero of Then hence we may replace the pair with the pair Thus, we may and will assume that is compact.
+Also, we may and will assume that Consider the -presheaf given by Then is a continuous -presheaf.
+By Proposition 6.12, for a compact we have a natural isomorphism
+Applying Propositions 6.10, 6.13 and 6.14, we obtain the following chain of isomorphisms
+This proves the theorem. ∎
 6.4. Compact objects
@@
 The implication is clear.
-. Suppose that is compact.
+.
+Suppose that is compact.
 We first show the following: if for some then there is an open neighborhood such that Indeed, since and is compact, we deduce that is a direct summand of some sheaf of the form where Hence the restriction of to vanishes.
-Now let be a point such that Since the functor is strongly continuous, the object is compact in Hence, there is an open neighborhood and a compact object such that Shrinking if necessary, we can lift the identity map of to a map in the category Denote by the associated map of sheaves.
-Choosing an open such that we see that is a compact object of whose stalk at vanishes.
+Now let be a point such that Since the functor is strongly continuous, the object is compact in By Proposition 1.71 there is an open neighborhood and a compact object such that Shrinking if necessary, we can lift the identity map of to a map in the category Denote by the associated map of -valued sheaves on Choosing an open such that we see that is a compact object of whose stalk at vanishes.
 By the above, we deduce that vanishes in a neighborhood of Hence, shrinking further, we may assume that as required.
-Finally, to see that has a compact support, note that Since is compact, it is a direct summand of some sheaf which has compact support. ∎
+Finally, to see that has compact support, note that Since is compact, it is a direct summand of some sheaf which has compact support.
+∎
 Corollary 6.16.
@@
 The category is compactly generated.
-The category is compactly generated, and is locally profinite.
+The category is compactly generated and is locally profinite.
 Proof.
 Consider the collection of sheaves
-Then each sheaf in is compact and the right orthogonal to consists of sheaves such that for every open-compact Since is locally profinite, this means that Hence, generates the category
-Choose a point The pullback to is a strongly continuous quotient functor Since the category is compactly generated, so is
-Next, we reduce to the case when is compact. If not, consider the short exact sequence
-Since the categories on the left and on the right are compactly generated, so is the category in the middle. Hence, we may and will assume that is compact.
+Then each sheaf in is compact and the right orthogonal to consists of sheaves such that for every open-compact Since is locally profinite, this means that Hence, generates the category Choose a point The pullback to is a strongly continuous quotient functor Since the category is compactly generated, so is Next, we reduce to the case when is compact.
+If not, consider the short exact sequence
+in Since the categories on the left and on the right are compactly generated, so is the category in the middle (by Proposition 3.3).
+Hence, we may and will assume that is compact.
 Suppose that is not profinite.
@@
 Choose a point Choose a non-zero compact object and consider the skyscraper sheaf on Then we can find a directed system of compact -valued sheaves such hat By Proposition 6.15 each is a locally constant sheaf with compact stalks.
-Since is compact in we can find and a splitting Using compactness of again, we can find an open neighborhood and a map which lifts Choose a compact such that is a neighborhood of Let be the connected component of in Then hence we may and will assume that
-So we have the map Take some Since the sheaf is locally constant, the set
+Since is compact in we can find and a splitting Using compactness of again, we can find an open neighborhood and a map which lifts Choose a compact such that is a neighborhood of Let be the connected component of in Then hence we may and will assume that So we have the map Take some Since the sheaf is locally constant, the set
 is open-closed in We have Since is connected, we have Choosing a point we see that the map is non-zero. But a contradiction. ∎
 Appendix A Monomorphisms and epimorphisms of presentable and dualizable categories
-Here we show that the monomorphisms and epimorphism in and are as expected: these are respectively the fully faithful functors and the quotient functors.
+Here we show that the monomorphisms and epimorphisms in and are as expected: these are respectively the fully faithful functors and the quotient functors.
 Proposition A.1.
-Let be a continuous functor between presentable stable categories.
-1) is a monomorphism in if and only if is fully faithful.
-2) is an epimorphism in if and only if is a quotient functor.
+Let be a continuous functor between presentable stable categories. 1) is a monomorphism in if and only if is fully faithful. 2) is an epimorphism in if and only if is a quotient functor.
 Proof.
@@
 is a monomorphism, hence is fully faithful (since and are stable).
-2) Again, the “if” part is clear. For the “only if” part, note that where both functors are Hence, the functor is an equivalence, in particular, it is fully faithful. Hence, for any the square
+2) Again, the “if” part is clear.
+For the “only if” part, note that where both functors are Hence, the functor is an equivalence, in particular, it is fully faithful.
+Hence, for any the square
 is bicartesian, thus is fully faithful and is a quotient functor.∎
 Proposition A.2.
-Let be a strongly continuous functor between presentable stable categories.
-1) is a monomorphism in if and only if is fully faithful.
-2) is an epimorphism in if and only if is a quotient functor.
+Let be a strongly continuous functor between presentable stable categories. 1) is a monomorphism in if and only if is fully faithful. 2) is an epimorphism in if and only if is a quotient functor.
 Proof.
-1) The “if” part is clear. We prove the “only if” part.
-Suppose that is a monomorphism. Then for a presentable stable category and strongly continuous functors we have By adjunction, we have
+1) The “if” part is clear. We prove the “only if” part. Suppose that is a monomorphism. Then for a presentable stable category and strongly continuous functors we have By adjunction, we have
 Put and consider the semi-orthogonal gluing Since is continuous, the inclusions of the two copies of are strongly continuous.
@@
 Applying (A.1) to we obtain
-hence which exactly means that is fully faithful. ‘
-2) follows from Proposition A.1 since the functor is conservative and commutes with colimits. ∎
+hence which exactly means that is fully faithful. ‘ 2) follows from Proposition A.1 since the functor is conservative and commutes with colimits. ∎
 Proposition A.3.
-Let be a strongly continuous between dualizable categories.
-1) is a monomorphism in if and only if is fully faithful.
-2) is an epimorphism in if and only if is a quotient functor.
+Let be a strongly continuous between dualizable categories. 1) is a monomorphism in if and only if is fully faithful. 2) is an epimorphism in if and only if is a quotient functor.
 Proof.
-1) The “if” direction is clear. We prove the “only if” direction.
+1) The “if” direction is clear.
+We prove the “only if” direction.
 Suppose that a functor is a monomorphism in By Proposition 1.87 the functor is fully faithful.
@@
 Arguing as in the proof of Proposition A.1, we conclude that is fully faithful.
-2) follows from Proposition A.1 since the functor is conservative and commutes with colimits. ∎
+2) follows from Proposition A.1 since the functor is conservative and commutes with colimits.
+∎
 Remark A.4.
@@
 Consider a homological epimorphism where and are small stable idempotent-complete categories.
-We know that if the category is compactly generated, then In particular, the essential image of is a stable subcategory of In this case by Thomason’s theorem every object of of the form can be lifted to
-It turns out that if is not compactly generated then the image of might not be a stable subcategory of More precisely, the following result shows that in general for a dualizable category there may exist an object such that the object cannot be lifted to
+We know that if the category is compactly generated, then In particular, the essential image of is a stable subcategory of In this case by Thomason’s theorem every object of of the form can be lifted to It turns out that if is not compactly generated then the image of might not be a stable subcategory of More precisely, the following result shows that in general for a dualizable category there may exist an object such that the object cannot be lifted to
 Proposition B.1.
-Let be a field and consider the dualizable category Then there exists a non-zero idempotent-complete stable subcategory such that and the intersection of with the image of is zero.
-In particular, the image of the homological epimorphism is not a stable subcategory.
+Let be a field and consider the dualizable category Then there exists a non-zero idempotent-complete full stable subcategory such that and the intersection of with the essential image of is zero.
+In particular, the essential image of the homological epimorphism is not a stable subcategory.
 Proof.
@@
 second copy of corresponds to the neighborhood of resp.
-We denote by the corresponding idempotents in the Calkin category, and denote their images by We obtain fully faithful functors which send to resp.
-We claim that any non-zero object of the image of is not contained in the image of
-Assume the contrary: suppose that for some sheaf the image of in the Calkin category is isomorphic to a non-zero object of Take some real number Since the constant sheaf is a compact object of it follows that the image of in is zero.
+We denote by the corresponding idempotent endomorphisms in the Calkin category, and denote their images by We obtain fully faithful functors which send to resp.
+We claim that any non-zero object of the image of is not contained in the essential image of Assume the contrary: suppose that for some sheaf the image of in the Calkin category is isomorphic to a non-zero object of Take some real number Since the constant sheaf is a compact object of it follows that the image of in is zero.
 Hence, the pullback of to is a compact object, i.e.
-for some perfect complex This holds for any hence for some Then and the image of in is isomorphic to the direct sum which is not contained in the image of This gives a contradiction.∎
+for some perfect complex This holds for any hence for some Then and the image of in is isomorphic to the direct sum which is not contained in the essential image of This gives a contradiction.∎
 Appendix C Presentability of the category of dualizable categories
-In this section we give an alternative, “explicit” proof of Ramzi’s theorem [Ram] which states that the category is -presentable.
-We also describe explicitly the -compact objects in for any uncountable regular We start with the following observation.
+In this section we give an alternative, “explicit” proof of the absolute version of Ramzi’s theorem [Ram24a] which states that the category is -presentable.
+We also describe explicitly the -compact objects in for any uncountable regular cardinal We start with the following observation.
 Proposition C.1.
-Let be a dualizable stable category. Consider the full subcategory which consists of pairs such that the image of generates by colimits. Then the functor is an equivalence of categories.
+Let be a dualizable stable category.
+Consider the full subcategory which consists of pairs such that the image of generates by colimits.
+Then the functor is an equivalence of categories.
 The inverse functor is given by
@@
 Indeed, for any the functor is conservative and commutes with all colimits, in particular, with geometric realizations.
-Hence, the functor is monadic by Lurie-Barr-Beck theorem, and we have an equivalence
-It remains to check that for any the category is dualizable and the functor is strongly continuous. The latter assertion is obvious, and the former one follows from Proposition 1.56.∎
+Hence, the functor is monadic by Lurie-Barr-Beck theorem, and we have an equivalence It remains to check that for any the category is dualizable and the functor is strongly continuous.
+The latter assertion is obvious, and the former one follows from Proposition 1.55.∎
 We will repeatedly use the following.
@@
 Theorem C.6.
-Put
-1) The category of dualizable categories is -presentable.
-2) Let be an uncountable cardinal and a dualizable category. The following are equivalent.
+Put 1) The category of dualizable categories is -presentable. 2) Let be an uncountable regular cardinal and a dualizable category. The following are equivalent.
 is a -compact object of
@@
 Proof.
-This is essentially a standard statement, we include the proof for completeness.
-. Since the functors and are -strongly continuous, so is the composition
-This composition is exactly the functor
-. Since the functors and are -strongly continuous, so is the composition
+This is essentially a standard statement, we include the proof for completeness. . Since the functors and are -strongly continuous, so is the composition
+This composition is exactly the functor . Since the functors and are -strongly continuous, so is the composition
 This composition sends to ∎
@@
 Consider a short exact sequence in
-1) If is -strongly continuous, then so is
-2) If is -strongly continuous, then so is
-3) If and are -strongly continuous, then so is
-4) If and are -strongly continuous, then so is
+1) If is -strongly continuous, then so is 2) If is -strongly continuous, then so is 3) If and are -strongly continuous, then so is 4) If and are -strongly continuous, then so is
 Proof.
 1) The functor is the composition of -strongly continuous functors
-hence it is -strongly continuous.
-2) Similarly, the functor is the composition of -strongly continuous functors
-hence it is -strongly continuous.
-3) It suffices to prove that is in We have an exact triangle
+hence it is -strongly continuous. 2) Similarly, the functor is the composition of -strongly continuous functors
+hence it is -strongly continuous. 3) It suffices to prove that is in We have an exact triangle
 Since is -compact, it suffices to show that is -compact. The latter object corresponds under the self-duality of to the composition
-The latter functor is -strongly continuous by our assumptions. Applying Lemma C.7 (to the categories and ) we conclude that is -compact.
-4) This is proved by reversing the proof of 3).∎
+The latter functor is -strongly continuous by our assumptions. Applying Lemma C.7 (to the categories and ) we conclude that is -compact. 4) This is proved by reversing the proof of 3).∎
 Lemma C.9.
@@
 the object is -compact.
-Denote by the continuous monad on associated with Then is a monad on and we have Since is a -compact monad, the -compactness of is equivalent to the -compactness of in The latter object is simply under the identification Buy our assumptions, is -compact.
+Denote by the continuous monad on associated with Then is a monad on and we have Since is a -compact monad, the -compactness of is equivalent to the -compactness of in The latter object is isomorphic to under the identification By our assumptions, is -compact.
 ∎
 Proof of Theorem C.6..
-We first show the following.
-Claim. For any dualizable category there exists a set and a monad such that
+We first show the following. Claim. For any dualizable category there exists a set and a monad such that
 Proof of Claim.
-Next, we prove some implications from part 3).
-. Recall that is a -compact object in Hence, any dualizable object in is also -compact. This proves the implication.
+Next, we prove some implications from part 2).
+.
+Recall that is a -compact object in Hence, any dualizable object in is also -compact.
+This proves the implication.
 .
@@
 is zero. By Proposition 1.73, we have We deduce that for some the composition
-is zero. This exactly means that the composition is strongly continuous.
-. Note that it is sufficient to prove that if is -compact in then the functors and for are -strongly continuous. Indeed, assuming this, we can apply Lemma C.8.
+is zero.
+This exactly means that the composition is strongly continuous.
+.
+Note that it is sufficient to prove that if is -compact in then the functors and for are -strongly continuous.
+Indeed, assuming this, we can apply Lemma C.8.
 So let be in It follows directly from Proposition 1.6 that the functor is -strongly continuous: this exactly means that the spectra are -compact.
-To show that (the diagonal -bimodule) is -compact, choose a -small ind-system of compact (finitely presented) objects of such that By a straightforward generalization of [TV07, Proposition 2.14], for each the object is compact in Hence, the object is -compact: it is a -small filtered colimit of compact objects, namely
-This proves the implication.
-. First, we have a short exact sequence
-Hence, satisfies and by the above it also satisfies If is a -small set, then also satisfies If is -compact, then by Lemma C.9 we see that the category satisfies
+To show that (the diagonal -bimodule) is -compact, choose a -small ind-system of compact (finitely presented) objects of such that and denote by the functors to the colimit.
+By a straightforward generalization of [TV07, Proposition 2.14], for each the object is compact in Hence, the object is -compact: it is a -small filtered colimit of compact objects, namely
+This proves the implication. . First, we have a short exact sequence
+from Proposition 4.21.
+Hence, satisfies (iv) and by the above it also satisfies (iii).
+If is a -small set, then also satisfies (iii).
+If is -compact, then by Lemma C.9 we see that the category satisfies (iii).
 We now prove 1).
@@
 Suppose that is -compact in We argue as in the proof of Claim.
-We see that is generated by a -small collection of -compact objects We obtain a monad such that Since is -compact in we deduce that is -compact in the category of dualizable categories under By Proposition C.1, this implies that is -compact.
+By the above, we see that is generated by a -small collection of -compact objects We obtain a monad such that Since is -compact in we deduce that is -compact in the category of dualizable categories under By Proposition C.1, this implies that is -compact.
 We already proved that the conditions (i), (ii), (iii) and (v) are equivalent, and they are implied by (iv).
 It remains to prove the implication .
-So let be -compact in By we have a strongly continuous functor such that the image generates.
+So let be -compact in By (ii), we have a strongly continuous functor such that the image generates.
 Put Define the category by the pushout square in
-Since is compactly generated, so is Since the categories and are -compact in so is It follows that the category is -compact in
-Since the functor is fully faithful, so is Consider the quotient Then is compactly generated and is -compact in Therefore, is -compact in Hence, the short exact sequence
+Since is compactly generated, so is Since the categories and are -compact in so is It follows that the category is -compact in Since the functor is fully faithful, so is Consider the quotient Then is compactly generated and is -compact in Therefore, is -compact in Hence, the short exact sequence
 has the required properties.∎
@@
 The category is generated by colimits by a single -compact object, namely by the category
-Note that the classical Urysohn’s lemma for compact Hausdorff spaces essentially says that the category (the opposite category of compact Hausdorff spaces) is generated by colimits by the unit interval The latter is easily seen to be an -compact object.
-We will again use the equivalence
-given by Proposition 5.3. Slightly abusing the terminology, we will call a functor strongly continuous if its left Kan extension to is strongly continuous. We denote the category of such functors by
-The following proposition essentially implies Theorem D.1.
+Note that the classical Urysohn’s lemma for compact Hausdorff spaces essentially says that the category (the opposite category of compact Hausdorff spaces) is generated by colimits by the unit interval The latter object is easily seen to be -compact in see Proposition F.5 below.
+Proof of Theorem D.1.
+We will consider the set as a linearly ordered set with the usual order. For we consider the subsets also as linearly ordered sets. We will again use the equivalence
+given by Proposition 5.3, where is a dualizable category.
+By Theorem C.6 we already know that the category is presentable and the object is -compact.
+Thus, we only need to prove the following: if a strongly continuous functor between dualizable categories induces an equivalence
+then is an equivalence. We first recall that the functor
+is essentially surjective by Corollary C.5, hence is also essentially surjective.
+To prove the fully faithfulness of we need the following observation: for any pair of strongly continuous functors we have pullback squares of spaces:
+and
+Since the posets and are isomorphic to the poset we conclude that the map
+is an equivalence of spaces.
+Now take any objects and choose strongly continuous functors such that Then we obtain the following chain of equivalences of spaces, where all the limits and colimits are over real numbers:
+Here the second equivalence follows from the isomorphism The third equivalence follows from the observation that for any we have an ind-equivalence The fourth map is an equivalence since (D.2) is an equivalence and we have obvious equivalences of posets Now, a similar chain of equivalences identifies the target of (D.3) with This proves the fully faithfulness of and the theorem.
+∎
+Remark D.2.
+Suppose that we know the cocompleteness of the category and the -compactness of the poset in (both are not difficult to check).
+Then the same argument as in the proof of Theorem D.1 shows that the category is generated via colimits by the poset A similar statement holds for the category of compactly assembled presentable categories and strongly continuous right exact functors.
+In this case the -compact generating object is the category
+For completeness we explain a direct way to recover the dualizable category from the category of strongly continuous functors Slightly abusing the terminology, we will call a functor strongly continuous if its left Kan extension to is strongly continuous.
+We denote the category of such functors by
 Proposition D.3.
-1) Let be a dualizable category. The functor is a Verdier quotient. Its kernel is the full subcategory of ind-zero functors, i.e. of functors such that for any there exists such that the map is zero.
+1) Let be a dualizable category. The functor
+is a Verdier quotient.
+Its kernel is the full subcategory of ind-zero functors, i.e.
+of functors such that for any there exists such that the map is zero.
 2) Denote by the right adjoint to Then for a strongly continuous functor we have
@@
 The notation simply means that we look only at In other words, the last condition on means that the endofunctor commutes with existing colimits.
-We will need the following basic fact about the sequential limits of directed colimits.
+We will need the following basic fact about sequential limits of directed colimits.
 Lemma D.4.
@@
 is final. It follows that for any functor we have
-where the two maps have components resp.
-Note also that the poset of order-preserving maps is cofinal in the poset of all maps Using our assumptions on we obtain the isomorphisms
+where the two maps have components resp. Note also that the poset of order-preserving maps is cofinal in the poset of all maps Using our assumptions on we obtain the isomorphisms
 Proof of Proposition D.3.
-Since we already know that the functor is essentially surjective, it suffices to prove part 2).
+We already know that the functor is essentially surjective by Corollary C.5.
+Hence, it suffices to prove part 2).
 Indeed, (D.4) implies that i.e.
 is a homological epimorphism.
-If is a function with required properties such that for all then the cofinality of implies that is an ind-zero functor Hence, the kernel of is generated by colimits by ind-zero functors, as required.
+Let be a function with required properties such that for all Then for any strongly continuous functor the functor is an ind-zero functor (since is cofinal).
+Hence, the kernel of is generated by colimits by ind-zero functors, as required.
 We now prove 2).
 Denote by the poset of functions with required properties.
-Then is a cofinal subposet of the poset of all monotone functions such that Since is directed, so is Consider another subposet which consists of monotone functions such that takes values in and is constant on and on each semi-open interval for Then is also cofinal in
-Proof of Theorem D.1.
-It suffices to prove the following: if a strongly continuous functor between dualizable categories induces an equivalence then is an equivalence.
-It is clear that is essentially surjective. The fully faithfulness follows directly from Proposition D.2. Namely, for strongly continuous we have
-Remark D.4.
-Suppose that we know the cocompleteness of the category and the -compactness of in (both are not difficult to check).
-It is easy to deduce from the proof of Proposition D.2 that is generated by colimits by
-The same applies to the category of compactly assembled presentable categories and strongly continuous right exact functors if we replace the poset with the poset with the usual order (the former poset is not a presentable category since it does not have an initial object).
+Then is a cofinal subposet of the poset of all (not necessarily strictly) monotone functions such that Since is directed, so is Consider another subposet which consists of monotone functions such that takes values in and is constant on and on each semi-open interval for Then is also cofinal in Consider two strongly continuous functors Note that for the functor is right Kan-extended from Applying Lemma D.4 and using the isomorphism we obtain the isomorphisms
+This proves (D.4). ∎
 Appendix E Phantom maps, pure injectivity and Adams representability
-Let be a dualizable category, and let be its homotopy category. As in Subsection 1.8, we denote by the ideal of compact maps in Recall that
+In this section we prove two closely related analogues of Adams representability theorem for -compact dualizable categories (Theorems E.2 and E.4).
+Let be a dualizable category, and denote by the triangulated homotopy categories of and We say that a morphism in is compact if it is an image (i.e.
+homotopy class) of a compact morphism in The following are straightforward generalizations of the usual notions for compactly generated categories [Kr00, GP04].
+Definition E.1.
+Let be a morphism in 1) is phantom if for any compact morphism with we have 2) is a pure monomorphism if the map is phantom. 3) is a pure epimorphism if the map is phantom.
+Recall that for a small triangulated category an additive functor is called homological if for any exact triangle in the sequence is exact.
+Clearly, homological functors are closed under extensions in the abelian category of left -modules (i.e.
+the category of additive functors ).
+We will also refer to homological functors as cohomological.
+Theorem E.2.
+Let be a dualizable category which is is -compact in We denote by resp.
+the triangulated homotopy category of resp.
+Consider the functor given by Then the functor is full and conservative, and its essential image consists of homological functors which commute with countable coproducts.
+Moreover, for we have a functorial short exact sequence of abelian groups
+The subgroup of phantom maps is identified with
+We will deduce Theorem E.2 from the closely related Theorem E.4 below. To formulate it, we need some preparations. We denote by resp. the ideal of compact maps in resp. Recall from Subsection 1.8 that
 Definition E.3.
 1) We say that a right -module is almost zero if it is annihilated by In other words, for any compact morphism in the map is zero.
-2) We denote by the quotient of by the full (Serre) subcategory of almost zero modules.
-The subcategory of almost zero modules is indeed a Serre subcategory since We make the following observation.
+The definition of an almost zero left -module is similar.
+2) We denote by the quotient of by the full (Serre) subcategory of almost zero right modules.
+Similarly, we denote by the quotient of by the full subcategory of almost zero left modules.
+The subcategories of almost zero modules are indeed Serre subcategories since and If is compactly generated, then both categories and are identified with (Proposition E.9 below).
+We say that an almost module is almost cohomological if it is an image of a right -module corresponding to a cohomological functor
+Theorem E.4.
+Let be a dualizable category which is -compact in Denote by the homotopy category of Then the functor is full and conservative, and its essential image is exactly the full subcategory of almost cohomological almost modules.
+More precisely, for we have a functorial short exact sequence of abelian groups:
+The abelian group of phantom maps is identified with The composition of two phantom maps is zero.
+Note that if is compactly generated, then is an -compact object of if and only if the triangulated category is countable.
+Thus, Theorem E.4 is a generalization of Neeman’s theorem [Nee97, Theorem 5.1, Proposition 4.11] if we restrict to triangulated categories with an -enhancement.
+Before proving Theorem E.4 we explain some basic properties of the abelian category and show that it is naturally equivalent to First we recall the following characterization of (co)homological functors.
+Recall that if is a small additive category with weak kernels, then a right -module is called fp-injective if for any finitely presented right -module we have
 Proposition E.5.
-The quotient functor has an exact left adjoint given by
+Let be a small triangulated category and an additive functor. The following are equivalent.
+is cohomological.
+is flat as a right -module.
+is fp-injective as a right -module.
 Proof.
+For an object we denote by the corresponding representable right -module.
+Flatness of means that for any finitely presented module any morphism factors through a representable module.
+Let for some morphism in Then the map is given by an element Note that is a weak cokernel of in Hence, the map factors through a representable module iff is in the image of the map We conclude that is flat if and only if is cohomological.
+Let be a finitely presented module, for some morphism in Let Then we have an exact sequence of -modules
+It follows that We conclude that is fp-injective if and only if is cohomological. ∎
+Next, recall from Proposition 1.48 that the tensor product is the quasi-ideal in given by We denote by the similar quasi-ideal in given by We make the following observation.
+Proposition E.6.
+The quotient functor resp. has an exact left adjoint resp. given by
+resp.
+Proof.
+Proposition E.7.
+Consider the colimit-preserving functor which is given on representable modules by Then is exact and it vanishes on almost zero right -modules. The induced functor
+is an equivalence.
+Proof.
+Exactness of means that for an object the left -module is flat.
+This follows from Proposition E.5.
+By Proposition E.6, to show the vanishing of on almost zero modules, we need to check that for any the map is an isomorphism.
+By exactness of and we may assume for some Let (in ).
+Then and
+This proves the vanishing of on almost zero modules.
+Now recall the functor from Remark 1.12 (with the roles of and interchanged).
+To avoid confusion, we denote the corepresentable left -modules by Consider the colimit-preserving functor given on corepresentable objects by We claim that Indeed, let with Choose factorizations in with Then we have and
+The isomorphism is functorial in hence we get an isomorphism as stated.
+Arguing as above, we see that the functor is exact.
+Now to prove that the functor (E.2) is an equivalence, it suffices to construct the isomorphisms
+We only need to construct these isomorphisms functorially on corepresentable left -modules. So let and let in Choose a pro-system in such that Then and we obtain
+Similarly, we obtain
+This proves the proposition. ∎
 We observe the following interpretation of local coherence of abelian categories.
@@
 Proof.
-. If is locally coherent, then the category is abelian and the inclusion functor is exact. Passing to ind-completions, we get the functor which is therefore exact.
+.
+If is locally coherent, then the category is abelian and the inclusion functor is exact.
+Passing to ind-completions, we get the functor which is therefore exact.
 .
@@
 Proposition E.9.
-1) If is compactly generated, then we have
-2) In general, the category is a Grothendieck abelian category which satisfies (AB4*) and (AB6). In particular, it is compactly assembled. The functor is exact.
+1) If is compactly generated, then we have 2) In general, the category is a Grothendieck abelian category which satisfies (AB4*) and (AB6).
+In particular, it is compactly assembled.
+The functor is exact.
 Proof.
@@
 Here the middle functor is exact by Proposition E.8 since the category has weak kernels. ∎
-Consider the functor which sends to considered as an almost -module.
+Consider the functor from Theorem E.4, i.e. considered as an almost -module.
 Lemma E.10.
@@
 It is clear that is a cohomological functor. Let be a collection of objects of By the definition of compact morphisms, we see that the -module
-is almost zero. This shows that commutes with coproducts.
-To show that is conservative, it suffices to check that if then But the vanishing of means that any compact morphism is zero. This implies that is zero.∎
-Definition E.6.
-Let be a morphism in
-1) is phantom if for any compact morphism we have
-2) is a pure monomorphism if the map is phantom.
-3) is a pure epimorphism if the map is phantom.
+is almost zero.
+This shows that commutes with coproducts.
+To show that is conservative, it suffices to check that if then But the vanishing of means that any compact morphism is zero.
+This implies that is zero.∎
 Note that is a phantom morphism resp. a pure monomorphism resp. a pure epimorphism in if and only if the morphism is zero resp. a monomorphism resp. an epimorphism in
@@
 We say that an object is pure injective if for any phantom morphism and for any morphism we have Equivalently, for any pure monomorphism and for any map there exists a map such that
-We refer to [Kr00, GP04] for the usual notions of phantom maps, pure monomorphisms and epimorphisms, and pure injective objects for compactly generated triangulated categories.
+We refer to [Kr00, GP04] for the usual notion of pure injective objects for compactly generated triangulated categories.
 We obtain the following description of the category of pure injective objects, which is a straightforward generalization of the corresponding result for compactly generated categories [GP04, Theorem 4.1].
@@
 Hence, it is represented by some object Moreover, is a pure injective object of since annihilates phantom morphisms.
-We obtain a functor from the category of injective almost -modules to the category of pure injective objects of Note that we have a functorial isomorphism
-Let now be a pure injective object, and choose a monomorphism where is injective.
+We obtain a functor from the category of injective almost -modules to the category of pure injective objects of Note that we have a functorial isomorphism Let now be a pure injective object, and choose a monomorphism where is injective.
 Then the map is a pure monomorphism.
-Since is pure injective, is a direct summand of Hence, is a direct summand of We conclude that is injective.
+Since is pure injective, it follows that is a direct summand of Hence, is a direct summand of We conclude that is injective.
 It remains to recall that is conservative, hence the map is an isomorphism for any pure injective ∎
-Recall that if is an additive category with weak kernels, then a right -module is called fp-injective if for any finitely presented right -module we have
-Proposition E.9.
-Let be a small triangulated category and an additive functor. The following are equivalent.
-is cohomological.
-is flat as an -module.
-is fp-injective as an -module.
-Proof.
-For an object we denote by the corresponding representable right -module.
-Flatness of means that for any finitely presented module any morphism factors through a representable module.
-Let for some morphism in Then the map is given by an element of Note that is a weak cokernel of in Hence, the map factors through a representable module iff is in the image of the map We conclude that is flat if and only if is cohomological.
-Let be a finitely presented module, for some morphism in Let Then we have an exact sequence of -modules
-It follows that We conclude that is fp-injective if and only if is cohomological. ∎
-We say that an almost module is almost cohomological if it is isomorphic to an image of a cohomological -module. Equivalently, this means that the -module is flat.
 Proposition E.13.
@@
 Proof.
-Indeed, for any object the -module is flat. ∎
-We can now prove a generalization of Adams representability theorem for -compact dualizable categories.
-Theorem E.11.
-Let be a dualizable category which is -compact in Denote by the homotopy category of Then the functor is full and conservative, and its essential image is exactly the full subcategory of almost cohomological almost modules.
-More precisely, for we have a functorial short exact sequence of abelian groups:
-In particular, the abelian group of phantom maps is identified with The composition of any two phantom maps is zero.
-Note that if is compactly generated, then is an -compact object of if and only if the triangulated category is countable.
-Thus, Theorem E.11 is a generalization of Neeman’s theorem [Nee97, Theorem 5.1, Proposition 4.11] if we restrict to triangulated categories with -enhancement.
-Proof.
+Indeed, for any object the -module is cohomological. ∎
+Proof of Theorem E.4.
 Essentially the theorem follows from the following statement.
-Claim. For any the object has injective dimension at most
-Assume that Claim holds. Take some objects and choose an injective resolution of of length
-Consider the pure injective objects and the map (where the functor is defined in the proof of Proposition E.8).
+Claim.
+Any almost cohomological almost -module has injective dimension at most Assume that Claim holds.
+Let be an almost cohomological almost -module.
+Choose an injective resolution of of length
+Consider the pure injective objects and the map where the functor is defined in the proof of Proposition E.12.
+Then the map is a pure epimorphism.
+Putting and using the fact that is cohomological, we obtain
+Hence, the essential image of is exactly the full subcategory of almost cohomological almost modules. Now take some objects and choose an injective resolution of as above:
 We obtain a (non-unique) map which becomes an isomorphism after applying Since is conservative, we obtain an exact triangle
 It follows that we have a short exact sequence
-By construction of the functor we see that the sequence E.2 is exactly of the form (E.1).
-In particular, we see that any phantom map factors through (and vice versa). If is another phantom map, then since is pure injective.
+By construction of the functor we see that the sequence (E.3) is exactly of the form (E.1).
+In particular, we see that any phantom map factors through (and vice versa).
+If is another phantom map, then since is pure injective.
 Proof of Claim..
@@
 Moreover, takes almost cohomological almost modules to flat modules, because the extension of scalars functor preserves flatness.
-We also observe that is exact: composing with the conservative exact functor we obtain the functor which is isomorphic to the composition of exact functors
-Now, applying Proposition E.9, we see that is fp-injective.
+We also observe that is exact: composing with the conservative exact functor we obtain a functor which is isomorphic to the composition of exact functors
+where as above Now, applying Proposition E.5, we see that for any almost cohomological almost module the -module is fp-injective.
 Since is countable, it follows that has injective dimension at most Since is exact, the functor preserves injective objects.
@@
 Theorem is proved. ∎
+Proof of Theorem E.2.
+The proof is obtained from Theorem E.4 using the relation between left -modules and right -modules.
+We say that a left almost module is almost homological if it is an image of a left -module corresponding to a homological functor Recall from Proposition E.6 that the (exact, fully faithful) left adjoint to the quotient functor is given by where is the quasi-ideal in given by The essential image of the functor consists of functors such that the composition commutes with sequential colimits.
+It follows that the functor gives an equivalence
+Indeed, this follows from the observation that for a homological functor the composition commutes with sequential colimits if and only if commutes with countable coproducts.
+Therefore, in the formulation of the theorem we can replace the functor with the composition
+and prove that this functor has the same properties as the functor from Theorem E.4. It remains to observe that the following square commutes
+where is the equivalence from Proposition E.7.
+It is clear that sends almost homological left almost -modules to almost cohomological right almost -modules.
+Therefore, the theorem follows from Theorem E.4.
+∎
 Remark E.14.
@@
 Namely, an object is fp-injective if for any and for any compact morphism the map
-is zero. If moreover is a Serre subcategory of then one can show that any fp-injective object of has injective dimension at most
-Now, for any dualizable category an almost -module is almost cohomological if and only if it is fp-injective in If is -compact, it can be shown that is a Serre subcategory in This gives an alternative (intrinsic) way to show that has injective dimension at most for any
+is zero.
+If moreover is a Serre subcategory of then one can show that any fp-injective object of has injective dimension at most Now, for any dualizable category an almost -module is almost cohomological if and only if it is fp-injective in If is -compact, it can be shown that is a Serre subcategory in This gives an alternative (intrinsic) way to show that has injective dimension at most for any
 Remark E.15.
@@
 The product in of a family is given by the subspace of the naive product, formed by elements such that with the norm given by
-Epimorphisms in are maps with a dense image. It follows that satisfies weak (AB4*), i.e. the class of epimorphisms is closed under products.
-Filtered colimits in are obtained by taking the completion (of the separation of) of the naive filtered colimit. Here the norm on the naive (filtered) colimit is defined by
-To show that (AB6) holds in recall the presentable compactly assembled category (of prenormed vector spaces and contractive maps) from Example 1.34.
+Epimorphisms in are maps with a dense image.
+It follows that satisfies weak (AB4*), i.e.
+the class of epimorphisms is closed under products.
+Filtered colimits in are obtained by taking the completion (of the separation of) of the naive filtered colimit.
+Here the norm on the naive (filtered) colimit is defined by
+To show that (AB6) holds in recall the presentable compactly assembled category (of seminormed vector spaces and contractive maps) from Example 1.33.
 Since is compactly assembled, it satisfies (AB6).
@@
 Proof.
-The functor is conservative by Urysohn’s lemma. By Tietze extension theorem, the functor detects epimorphisms (and also detects regular epimorphisms).
+The functor is conservative by Urysohn’s lemma.
+By Tietze extension theorem, the functor detects epimorphisms (and also detects regular epimorphisms).
 Now, the product of in is given by – Stone-Čech compactification of the disjoint union.
@@
 Consider a codirected system and put for Passing to Boolean algebras, we need to show that the map is an isomorphism.
-Given an open-closed consider its characteristic function Then for some we have a continuous function such that Then is an open-closed subset of and its preimage in equals
-Suppose that is an open-closed subset such that Then hence for some we have i.e. ∎
+Given an open-closed consider its characteristic function Then for some we have a continuous function such that Then is an open-closed subset of and its preimage in equals Suppose that is an open-closed subset such that Then hence for some we have i.e.
+∎
 Proposition F.7.
@@
 Proof.
-Consider the surjective map Put and Then
-Put Then the diagram is the limit of but is a one-element set for Hence, the sequential limits do not commute with coequalizers in ∎
+Consider the surjective map Put and Then Put Then the diagram is the limit of but is a one-element set for Hence, the sequential limits do not commute with coequalizers in ∎
 Appendix G K-theory of products of exact infinity-categories
-In this section we prove that -theory of exact -categories commutes with infinite products.
+In this section we prove that -theory of exact -categories commutes with infinite products (Corollary G.9).
 G.1. Products of additive infinity-categories
 Given a small additive -category we denote by its stabilization.
-By [Bon, Theorem 5.3.1, Proposition 6.2.1], if is Karoubi complete, then so is Given we denote by the full subcategory generated by We put and similarly
-The following result is due to A. Córdova Fedeli.
+By [Bon, Theorem 5.3.1, Proposition 6.2.1], if is Karoubi complete, then so is Given we denote by the full subcategory generated by We put and similarly The following result is due to A.
+Córdova Fedeli.
 Theorem G.1.
 [Cor23, Proposition 2.10] Let be a collection of Karoubi complete additive small -categories. We have an isomorphism in
-Since we will need an analogous statement for -structures, we recall the proof from [Cor23] with some simplifications.
-We introduce some notation. Put Given functions we put
+Since we will need an analogous statement for -structures, we recall the proof from [Cor23] with some simplifications. We introduce some notation. Put Given functions we put
 For any function we denote by the product of the functors
 Lemma G.2.
-2) Analogous statement holds for the same functors
+[Cor23, Lemma 2.9] 1) For each the functors
+have the same class in 2) Analogous statement holds for the same functors
 Proof.
@@
 Then and the map becomes an isomorphism in This proves that is essentially surjective.
-Let now Let be objects of and denote by and their images in
-Fullness of Consider a map represented by a roof where and Take such that and Choose and the maps such that Then we can choose factorizations and automatically
-Putting we have and Therefore, the roof represents a morphism such that
-Faithfulness of Consider a map such that We may assume that can be lifted to a morphism Then there exists a map such that and Arguing as above, we can construct a factorization of such that and Hence ∎
+Let now be objects of and denote by and their images in Fullness of Consider a map represented by a roof where and Take such that and Choose and the maps such that Then we can choose factorizations and automatically Putting we have and Therefore, the roof represents a morphism such that Faithfulness of Consider a map such that We may assume that can be lifted to a morphism Then there exists a map such that and Arguing as above, we can construct a factorization of such that and Hence ∎
 G.2. Products of stable categories with bounded -structures
 We will need the following analogue of Theorem G.1.
-Let be a family of stable categories with bounded -structures.
+Let be a family of small stable categories with bounded -structures.
 Denote by the full subcategory formed by objects such that the objects have uniformly bounded homological amplitude.
@@
 The proof is completely analogous to the proof of Theorem G.1 with minor changes, which we indicate.
-For a stable category with a bounded -structure and for integers we denote by the full subcategory formed by objects whose homology is concentrated in degrees between and and similarly for
-Analogously to Lemma G.3, one shows that we have an isomorphism
-Analogously to Lemma G.4, one proves that for any function the functor is an equivalence.
+For a stable category with a bounded -structure and for integers we denote by the full subcategory formed by objects whose homology is concentrated in degrees between and and similarly for Put Given functions we define the subcategories analogously to (G.1)-(G.3).
+Analogously to Lemma G.3, one shows that we have an isomorphism Analogously to Lemma G.4, one proves that for any function the functor is an equivalence.
 Then one concludes that proving the theorem.∎
@@
 [Bar15, Definition 3.1] An exact -category is a triple where is an additive -category and are subcategories such that the following conditions hold.
-1) Every morphism (resp. ) is in (resp. ).
-2) Pushouts of morphisms in exist, and the class of morphisms in is closed under pushouts. Pullbacks of morphisms in exist and the class of morphisms in is closed under pullbacks.
+1) Every morphism (resp.
+) is in (resp.
+).
+2) Pushouts of morphisms in exist, and the class of morphisms in is closed under pushouts.
+Pullbacks of morphisms in exist and the class of morphisms in is closed under pullbacks.
 3) For a commutative square
-the following are equivalent:
-a) the square is coCartesian, and
-b) the square is Cartesian, and
+the following are equivalent: a) the square is cocartesian, and b) the square is cartesian, and
 By definition, a short exact sequence is a bicartesian square of the form
-where and A total cofiber of such a square is by definition a primitive acyclic object.
+where and A total cofiber of such a square in is by definition a primitive acyclic object.
 The full subcategory of acyclic objects is the stable subcategory generated by the primitive acyclic objects.
@@
 Note that if is a triangulated category with a -structure, and is a weak Serre subcategory of the heart, then the full triangulated subcategory generated by has a bounded -structure with the heart In particular, the functor is -exact.
+Consider the category with its standard -structure.
+By [Kle22, Proposition 3.10], the primitive acyclic objects are contained in the heart Moreover, by [Kle22, Lemma 3.11 and Lemma 3.12] they form a weak Serre subcategory.
+The proposition follows.∎
 Theorem G.8.
@@
 We have a map of short exact sequences of stable categories:
+By Theorem G.1, applied to the middle vertical arrow is an isomorphism.
+By Theorem G.5 applied to the bounded -structures on from Proposition G.7, applied to the left vertical arrow is an isomorphism.
+Hence, applied to the right vertical arrow is an isomorphism.
+This proves the theorem.∎